### v3

# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

Abstract
In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence and adaptability, and are rapidly changing human production and life. Nowadays, agents are undergoing a new round of evolution. They no longer act as an isolated island like LLMs. Instead, they start to communicate with diverse external entities, such as other agents and tools, to perform more complex tasks collectively. Under this trend, agent communication is regarded as a foundational pillar of the future AI ecosystem, and many organizations have intensively begun to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the recent few months. However, this new field exposes significant security hazards, which can cause severe damage to real-world scenarios. To help researchers quickly figure out this promising topic and benefit the future agent communication development, this paper presents a comprehensive survey of agent communication security. More precisely, we first present a clear definition of agent communication and categorize the entire lifecycle of agent communication into three stages: user-agent interaction, agent-agent communication, and agent-environment communication. Next, for each communication phase, we dissect related protocols and analyze the security risks according to the communication characteristics. Then, we summarize and outlook on the possible defense countermeasures for each risk. In addition, we conduct experiments using MCP and A2A to help readers better understand the novel vulnerabilities brought by agent communication. Finally, we discuss open issues and future directions in this promising research field.
Index Terms:
large language model, agent communication, attack, and securityI Introduction
The emergence of Large Language Models (LLMs) has led to revolutionary advancements in Artificial Intelligence (AI), exhibiting unprecedented capabilities in understanding complex tasks [327]. More importantly, LLMs greatly boosted the ideal form of AI that human expects: agents111In this paper, all agents refer to LLM-driven AI agents.. Different from LLMs that are mainly like chatbots, agents possess more comprehensive capabilities (e.g., perception, interaction, reasoning, and execution), enabling them to independently complete a real-world task. For example, when users seek to make a travel plan, LLMs can only provide the best plan in text, while agents can realize the best plan in action, such as checking the weather, buying tickets, and booking hotels. Agents greatly speed up the progress of the intelligence transformations of enterprises. Their market size is expected to increase by 46% per year [235]. It can be foreseen that agents will subvert the production and living patterns of modern society, greatly changing the future business landscape. As a result, developing and promoting agents have become a strategic planning of major powers and influential companies.
Now, agents are evolving towards the direction of domain-specific entities, i.e., being customized for specific scenarios and tasks. In this context, as shown in Figure 1, a task usually requires the collaboration of multiple agents, which may be located globally on the Internet. Under this condition, agent communication becomes the foundation of the future AI ecosystem. It enables agents to find other agents with specific capabilities, access external knowledge, assign tasks, and engage in other interactions. Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to the development of agent communication. In November 2024, Anthropic proposed Model Context Protocol (MCP) [19], a universal protocol that allows agents to invoke external environments, such as datasets, tools, and APIs. MCP quickly gained a great deal of attention in the recent few months. Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [214], Google [96], Microsoft [57], Amazon [25], Alibaba [10], and Tencent [267], and MCP’s package receives over 3 million weekly downloads [21]. In April 2025, Google proposed Agent-to-Agent Protocol (A2A) [89], which enables seamless communication and collaboration among agents. Since its release, A2A has received extensive support from many enterprises, such as Microsoft [200], Atlassian [160], and PayPal [242]. It can be seen that the breakthroughs in agent communication are bringing rapid and profound changes and will become an indispensable part of the AI ecosystem.
However, the rapid development of agent communication also introduces complex security risks that could cause severe damage to the AI ecosystem. For example, the collaboration of cross-organization agents significantly enlarges the attack surface, causing severe security risks, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks. Since the research related to agent communication is still in the nascent stage, it urgently needs a systematic review of the security problems existing in the complete agent communication lifecycle. Following this trend, this paper aims to provide a comprehensive survey of existing agent communication techniques, analyze their security risks, and discuss possible defense countermeasures. We believe this work can help a broad range of readers, such as researchers who are devoted to agent development and beginners who have just started their journey in AI.
The contributions of this paper are as follows:
We proposed the definition of agent communication for the first time, and classified it into three stages based on the communication object, which covers the entire lifecycle of agent communication.
We comprehensively studied and classified the existing agent communication protocols. Besides, we thoroughly analyzed and categorized the related security risks for each communication stage, and detailedly discussed the targeted defense countermeasures.
We conduct experiments using MCP and A2A to help readers better understand the new attack surfaces brought by agent communication. The results show that attackers can easily cause severe damage to users’ systems and privacy.
We finally discuss the open issues and future research directions. We not only point out the much-needed techniques but also explain the demand for related laws and regulations.
Organization. As shown in Figure 2, we organize this survey as follows. Section II compares the most relevant surveys with this paper and outlines the novelties in this survey. Section III introduces the preliminaries of this survey. Section IV presents a definition and classification of agent communication. Section V introduces user-agent interaction protocols and analyzes related security risks and defense countermeasures. Section VI exhibits agent-agent communication protocols, related security risks, and corresponding defense countermeasures. Similarly, Section VII shows the protocols, risks, and defenses for agent-environment communication. In Section VIII, we conduct experiments using MCP and A2A to help illustrate the risks brought by agent communication. In Section IX, we discuss the open issues and future research direction. Section X concludes this survey.
II Related Work
| Survey | Year | Research Object | Rele. | Agent Communication | Protocol Coverage | Security Analyses | Defense Prospect | Exp. | |||||||||
| Moti. | Defi. | Clas. | Clas. | U-A | A-A | A-E | U-A | A-A | A-E | U-A | A-A | A-E | |||||
| [170] | 2024 | Personal Agent | 4 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ |
| [77] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [163] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✗ | ✔ | ✓ – | ✗ | ✗ |
| [143] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✗ |
| [56] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✓ – | ✗ |
| [108] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✗ |
| [295] | 2025 | General IoA | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
| [103] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [317] | 2025 | Agent Comm. | 6 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [296] | 2025 | Agent Secu. | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✔ | ✔ | ✓ – | ✔ | ✔ | ✗ |
| [281] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [297] | 2024 | Agent Secu. | 6 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [75] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ |
| [241] | 2025 | Agent Comm. | 7 | ✓ – | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ | ✗ |
| [325] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [95] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ |
| [318] | 2025 | Agent Comm. | 8 | ✔ | ✗ | ✔ | ✔ | ✓ – | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ |
| [147] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✔ |
| [252] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ |
| [233] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ |
| [67] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
| [111] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ |
| This survey | Agent Comm. Secu. | / | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | |
II-A Overview of Novelties in This Survey
Table I summarizes the characteristics of the most relevant surveys and the differences between this survey and previous surveys. In summary, this surveys exhibit the following novelties:
This survey presents a comprehensive illustration of agent communication. Specifically, it explains why the current agent ecosystem needs communication (i.e., the predicament faced by a single agent, Section IV-A), gives the definition of agent communication (Section IV-B), and proposes a novel classification principle based on communication entity, which can cover the entire lifecyle of agent communication (Section IV-C). As a result, future studies can be included and categorized according to our survey seamlessly.
This survey exhibits a comprehensive illustration of the existing protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A), including newly proposed and previously neglected protocols that have not been discussed by other surveys. Besides, we categorize these protocols based on their architecture and summarize corresponding characteristics, rather than mechanically listing each protocol. This organization method can allow any researchers interested in this field to quickly establish a preliminary but comprehensive understanding of agent communication.
This survey makes an in-depth analysis of the discovered attacks and potential risks that have not been revealed for each agent communication stage. We discuss both the risks from compromised agents and those from malicious communication objects (i.e., users and environments), fully covering the entire agent communication lifecycle. Then, we thoroughly outline the possible defense countermeasures (Sections V-D, VI-C, and VII-D) that can make future agent communication more secure.
This survey conducts experiments using MCP and A2A (Section VIII), the most popular agent communication protocols up to now. We successfully launched attacks against MCP and A2A, showing that attackers can cause severe damage with little effort. This section can help readers better understand the new attack surfaces brought by agent communication.
II-B Selection Principles of the Most Relevant Surveys
Challenge. Our survey aims at comprehensively studying the protocols, related security risks, and possible defenses of agent communication. However, there are a lot of surveys that seem relevant but are actually different in essence. Listing these surveys is not conducive to readers’ understanding of this field as efficiently as possible, especially for those who want to read the original texts of these surveys.
To solve this challenge, when selecting the most relevant surveys, we focus on three principles: LLM-driven agents, agent communication, and security. However, to our knowledge, there have not been papers systematically discussing all three of these themes. As a result, as long as a survey meets two of the three principles, we will treat it as a relevant survey.
Principle #1: LLM-driven agents. The first and the most important is that the research object of a survey must be LLM-driven agents. This principle must be satisfied. This is because there have been many studies about multi-agent systems (MAS) before the emergence of LLMs. These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits very little to this survey. Besides, surveys focusing on only LLMs instead of LLM-driven agents are also not listed in Table I (but we will draw on their valuable insights in other sections). This is because LLMs show significant differences from agents, for which we make a detailed illustration in Section III-C. As a result, researching LLM-driven agents is the most important principle.
Principle #2: agent communication. The second principle is that a survey should focus on or partially discuss agent communication, especially including some typical agent communication protocols such as MCP. This is because agent communication is very different from agent. However, if a survey satisfies the other two principles (i.e., LLM-driven agents and security), we still treat it as a relevant survey.
Principle #3: security. The final principle is that a survey should focus on or partially discuss agent-related security. This is because we believe that the security risks of agents still have meaning to the security risks of agent communication. The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agent.
Relevance Score. As a result, we can find that there are two main types of relevant surveys: LLM-driven agents + communication, or LLM-driven agents + security. As shown in Table I, we list a relevance score for each survey. The higher the score, the more relevant we think it is to our survey. This score is subjectively derived by us after carefully reading the paper and does not have an objective calculation method. This is because we found that the forms of surveys are highly diverse, and it is hard to accurately classify them using only several metrics. As a result, we directly present the score based on our subjective feelings when reading these surveys.
II-C Detailed Comparison with the Most Relevant Surveys
In this section, we will detailedly compare the most relevant surveys in Table I with our survey.
Survey [170] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computers. It only partially mentions the security risks related to personal agents in a section. Besides, these risks only belong to the user-agent interaction phase. It also does not discuss agent communication and related security risks and defenses. Survey [77] focuses on agent security instead of agent communication security. It is a security-specific paper, so its discussion of security is more comprehensive compared to [170]. However, the main body of its discussion is about the interaction between user and agent (U-A), without enough consideration about agent-agent (A-A) or agent-environment (A-E) interaction, which have significantly different characteristics. Besides, it also does not include any protocols related to agent communication. Survey [163] also focuses on agent security instead of agent communication security, which is similar to [77]. This survey focuses on single-agent systems and partially discusses multi-agent collaboration. It does not consider agent communication, related protocols, and enough security analyses about A-A and A-E. Survey [143] systematically summarizes seven security challenges for multi-agent systems. As shown in Table I, its main focus is on A-A, and only partially discusses U-A and A-E, which is not comprehensive. Besides, it does not consider agent communication and related protocols. Survey [56] proposes four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E. Besides, it does not consider agent communication or any related protocols. The defense prospect is also limited. Survey [108] focuses on the security risks of U-A and A-E, such as malicious API. It does not consider agent communication and related protocols. Besides, its security analyses are also not comprehensive enough. Survey [295] focuses on the fundamentals, applications, and challenges of IoA. Since its focus is different, agent communication and related security are only partially mentioned. Specifically, it only introduces a few related protocols and briefly analyzes related security. Besides, it also lacks the related illustration (such as definition and classification) of agent communication. Survey [103] also concentrates on the security of U-A, partially discussing A-E. It does not mention agent communication and related protocols, as well as the risks of A-A. Survey [317] focuses on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view. Besides, it only partially mentioned related security and did not discuss any communication protocols. Survey [296] focuses on the security of IoA. It mentioned a few agent communication protocols (i.e., MCP, A2A, ANP, and Agora), neglecting many other important protocols. Besides, it lacks the motivation, definition, and classification of agent communication, and also does not classify protocols. According to our analyses, the security analyses (especially for U-A) are also not comprehensive enough. Survey [281] proposes the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization. It does not focus on agent communication security. As a result, this survey did not give a clear illustration of agent communication, only mentioned a few protocols (i.e., MCP, A2A, ANP, and Agora), and partially discussed related threats and countermeasures. Survey [297] gives a comprehensive analysis of the security of agent networks. However, it does not include the discussion of communication protocols, and lacks enough security analyses of A-A and A-E. Survey [75] does not focus on security. Instead, it concentrates on evaluating LLMs and agents. Besides, it also analyzes the architecture of some communication protocols (i.e., MCP, A2A, and ANP). We can see that it does not give a detailed illustration of agent communication, enough coverage of protocols, or a comprehensive discussion about security. Survey [241] focuses on MCP, detailedly analyzed related architectures and applications. It does not consider other communication protocols and only partially mentions security-related content. Survey [325] analyzes the threats of agents and divides them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E. However, its analyses are not comprehensive enough, and it does not mention any communication protocols. Survey [95] makes a comprehensive analysis of the risks for multi-agent systems. However, its focus only falls within A-A, not considering U-A, A-E, and related protocols.
Survey [318] is one of the surveys with the highest relevance score because it focuses on agent communication and analyzes related protocols. However, it still has significant differences from our survey. First, it lacks some critical protocols like AG-UI, ACP-AgentUnion, ACN, Agent Protocol, API Bridge Agent, and Function Calling. Second, it lacks the analysis of security threats. Third, its defenses prospect is limited. Survey [147] focuses on the influences of MCP. As a result, it lacks other protocols, the illustration of agent communication, and security-related discussion. Similarly, the survey [252] also focuses on MCP. It lacks illustrations of other protocols and comprehensive security analyses. Survey [233] comprehensively introduces A2A, lacking discussion about other protocols and security analyses. Survey [67] detailed discussed MCP, A2A, ANP, and ACP(-IBM). It also partially analyzed related security risks and defenses. However, there still lacks other protocols, in-depth security analyses, and systematic illustration of agent communication. Hou et al. [111] discussed the security risks of MCP. They did not consider other protocols and the high-level overview of agent communication.
III LLM-Driven AI Agents
In this section, we review the entire lifetime from LLM to LLM-driven AI agent. Our goal is to help beginners quickly figure out agents, their characteristics, relationships, and applications.
III-A Large Language Model
Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [211]. Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [330], logic reasoning [225, 356, 300], code generation [344], and translation [222]. These remarkable performances can be attributed to two major factors. One is that LLMs are built upon a powerful architecture known as the Transformer [276], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input. The other key factor, perhaps the most important one, is the massive scales of LLMs that far exceed traditional AI models. When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [299], referring to unexpected capabilities that do not appear in smaller models. As shown in Table II, the parameter scale of an LLM can be hundreds or thousands of times that of traditional AI models.
| Architecture | Model | Year | Parameters |
| FNN | MLP | 1990s | 100K |
| FNN | LeNet-5 [157] | 1998 | 60K |
| RNN | Elman Net [69] | 1990 | 100K |
| LSTM | LSTM [109] | 1997 | 1–10M |
| CNN | ResNet-50 [104] | 2015 | 25M |
| CNN | AlexNet [148] | 2012 | 60M |
| CNN | VGG-16 [251] | 2014 | 138M |
| GAN | DCGAN [228] | 2016 | 4M |
| GNN | GCN [142] | 2017 | 23K |
| Autoencoders | DAE [277] | 2008 | 100K |
| Autoencoders | VAE [141] | 2013 | 1M |
| Transformer | GPT-3 [29] | 2020 | 175B |
| Transformer | PaLM [43] | 2022 | 540B |
| Transformer | GPT-4 [5] | 2023 | 1T |
| Transformer | DeepSeek-V3 [176] | 2024 | 671B |
| Transformer | DeepSeek-R1 [94] | 2025 | 671B |
III-B LLM-Driven AI Agents
Figure 3 illustrates a typical architecture of LLM-driven agents. Different from LLMs that mainly act as chat bots and do not possess professional ability in specific domains, agents are designed to automatically help humans to finish specialized tasks. To this end, agents are equipped with multiple modules to become more all-powerful. As shown in Figure 3, there are usually five modules in agents: perception, memory, tools, reasoning, and action.
Perception module. To automatically finish a specified task, agents need the ability to perceive the real-world environment. For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [195, 190]. The type of perception ability depends on the domain for which the agent is designed. For instance, an autonomous driving agent needs the ability of visual or radar perception [258, 321], while a code-generating agent may not require such functions [116, 124].
Memory module. The processing of real-world tasks also requires a strong memory. Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [102, 191, 328]. This usually requires external storage resources to assist the brain, such as databases or memory sharing [80, 82]. In contrast, LLMs do not have such excellent memory ability. Their memory is short-term, which only lasts for rounds of conversations [284, 357].
Reasoning and planning module. LLM acts as the brain of agents due to its excellent capability of reasoning and planning. It intercepts the instructions from users and automatically decomposes the received task into multiple feasible steps [291, 127, 139, 253]. Then, it selects the best plan from different candidates [136, 354, 113]. Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [361, 275, 234, 305]. For example, when the autonomous driving module finds that the barrier is closer, it will change the plan to slow down or detour.
Tool module. The tool module is responsible for deeply integrating external resources with the cognitive capabilities of the agent, enabling it to perform complex operations beyond the native capabilities of LLM [331, 304, 167, 186]. For example, through predefined functional interfaces and protocols, a math agent is able to invoke the external computation libraries and symbolic solvers to help it solve mathematical problems [90].
Action module. The action module is the core hub for interaction with the environment. It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [290, 345]. This module ensures the executability of instructions through structured output control. For example, it immediately stops generating when LLMs generate a complete action description to avoid redundant output interfering with subsequent parsing.
By integrating the above modules, agents establish a closed-loop system that achieves a full chain of perception-decision-action-feedback. As a result, agents achieve unprecedented ability in automatically finishing domain-specific tasks, being closer to the ultimate form of AI that human expects.
| Metric | LLM | Agent |
| Autonomy | Prompt-dependent | Autonomous |
| Multimodal interaction | Limited | Strong |
| Tool Invocation | Simple API | Various tools |
| Hallucination inhibition | Weak | Strong |
| Dynamic adaptability | Limited | Strong |
| Collaboration ability | Limited | Strong |
| Security | Better | Worse |
III-C Comparison Between Agents and LLMs
Table III illustrates the advantages of agents over LLMs on different metrics. Overall, agents have many advantages over LLMs, except for security.
High Autonomy. LLMs can only passively react to the user prompts and then generate responses. They are unable to plan or execute tasks independently. Besides, the response quality highly relies on the prompt skill [302, 85, 203, 178, 68, 33, 360], which seriously affects the user experience. In contrast, agents possess independent capabilities for task decomposition, strategy adjustment, and external tool invocation, which breaks through the passive mode of LLMS and is highly autonomous.
Flexible Multimodal interaction. LLMs have limited capability of handling multimodal inputs, such as text and pictures [240, 338, 342, 360, 155, 363]. Besides, their outputs are also mainly single-modal (e.g., text-only or picture-only), lacking the ability to actively invoke tools to perform physical actions or generate multimodal content. In contrast, agents overcome these drawbacks by deploying multimodal perception frameworks and tool invocation interfaces. They can realize interactions with complex environments, including vision, text, voice, and other physical elements.
Abundant Tool invocation. LLMS usually passively invoke a single tool (such as Function Calling [213]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [337]). In contrast, agents have the capability of active decision-making. They can independently select, combine, and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [107].
Better Hallucination inhibition. LLMs suffer from a serious problem called hallucination, which refers to that LLMs are likely to generate non-existent knowledge [348, 93, 272, 118, 315, 169]. LLMs mainly rely on the knowledge internalization of training data, making them prone to hallucinations when facing uncovered domains or outdated information. In contrast, agents are able to reduce the error rate by integrating multiple techniques such as Retrieval Augmentation Generation (RAG) [84, 162, 353] or other methods, which can align the action of agents [268, 79].
Dynamic adaptability. Essentially, LLMs are static models whose knowledge is fixed at the training phase. Although techniques such as fine-tuning [334, 175, 115] or model editing [322, 288, 341, 283, 171] reduce the training cost significantly, LLMs still cannot adapt to real-time events well. In contrast, agents are equipped with techniques like online web search, database query, or real-time sensors, which enable them to dynamically adapt to the changes of real-time environments and information.
Stronger Collaboration ability. LLMs lack enough collaboration ability when handling complex tasks. First, LLMs cannot interact with tools well; they can only access limited external assistance via simple APIs. Second, different LLMs lack effective cooperation mechanisms. In contrast, agents have designs for multi-agent collaboration. For example, MCP enables agents to use unified integration of external tools, and A2A allows agents from different enterprises to cooperatively finish a task.
Worse Security. Agents have WORSE security than LLMs, which is a major weakness of agents. This is because LLMs are only capable of outputting text. Even if the outputs contain illegal or discriminatory content, their influence on the real world is limited. In contrast, since agents are endowed with the ability to invoke tools, they can cause substantial damages to the real world, including but not limited to maliciously/wrongly operating machines, poisoning databases, and paralyzing the system. As a result, it is necessary to concentrate more on the security of agents.
III-D Agent Applications
Due to the strong advantages that agents have shown, related applications are booming. They span multiple domains, from scientific research to engineering systems and social services. Since the application of agents is not the focus of this paper, we will present a brief overview of their practical use cases to illustrate the rapid popularization of agents.
Scientific Research. Agents are increasingly embedded into the research workflow, enhancing ideation, automation, and discovery. Their contributions span multiple disciplines, such as mathematics [161, 55, 312, 285], chemistry [239, 50, 42, 27], biological sciences [313, 309, 177], and materials Science [151, 219, 197]
Technical and Engineering Systems. Agents play a growing role in engineering domains, improving automation, systems, and software intelligence. For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [280, 37, 125, 114, 138, 187]. Besides, agents are also popular in game development and simulation [201, 255]. Embodied intelligence is also another hot topic [36, 194].
Social Governance and Public Services Agents are increasingly deployed in sectors focused on public service and human welfare. For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [270, 121, 199, 269]. Besides, other fields, such as financial services [164, 319, 212, 329, 98, 97, 73], education [287, 259, 193, 59, 64, 206], and healthcare [35, 298, 74, 238, 23, 123, 289, 159, 346], are also actively integrating agents into their respective practices.
Overall, it can be seen that agents are being widely applied in all walks of life, greatly promoting the development of productivity. More importantly, the application of agents is still in its infancy and has an even greater space for development in the future. It is estimated that the agent market will grow at a rate of 40% annually and is expected to exceed 216.8 billion dollars by 2035 [15].
III-E Takeaways
Agents show multiple advantages over LLMs on multiple metrics, such as richer perception ability, stronger learning ability, and higher adaptivity. Now, to improve the service quality, agents are evolving towards refinement to obtain professional skill in a small domain, no longer pursuing the comprehensive capabilities like LLM. LLMs are more like an intermediate transitional form of the future intelligence, while agents are the next stage of development direction of artificial intelligence. It can be foreseen that they will ultimately become indispensable components of future production ecosystems and daily life. However, agents show worse security than LLMs due to their capability of executing tools. As a result, studying the security of agent communication is significant to the AI ecosystem.
IV Agent Communication Overview
IV-A Motivation: The Demand for Agent Communication
Although the advantages of agents in various fields have become increasingly obvious, their development has also encountered new obstacles, which have given rise to the demand for agent communication.
Conflicting development trends. The first reason derives from the fundamental conflict between the refined development of agents and the abstract demands of users. With the in-depth and specialized evolution of agents towards vertical fields (such as medical diagnostic agents, financial risk control agents, and industrial control agents), their capability boundaries are becoming increasingly refined. However, users’ usage habits exhibit opposite characteristics: they tend to input simple and abstract instructions (such as “plan a cross-border travel”) to trigger the execution of complex tasks. It is hard for a domain-specific agent to finish such abstract instructions independently. Besides, this trend of users’ habits is hard to reverse. This is because people always prefer applications that are easy to operate rather than those that require cumbersome usage steps. The latter is at a disadvantage in the market competition. Usually, for each additional operation step, the user churn rate increases by 10% - 20%. Therefore, agents should not only NOT ask users to make changes, but also cater to such demands of users, which is contrary to the development direction of agents.
Closed ecosystem. The current mainstream multi-agent systems adopt a closed ecosystem design and rely on private interaction mechanisms, forming a rigid technical barrier. This development paradigm severely limits the dynamic collaboration capabilities with external systems. For example, external agents cannot be actively discovered or invoked, making it extremely hard for cross-platform collaboration. Besides, such a closed ecosystem further triggers a decline in systemic effectiveness. On one hand, agents lose scalability due to binding private tools. On the other hand, the obstruction of cross-platform knowledge sharing suppresses the intelligence of agents. As a result, it is necessary to propose agent communication mechanisms that integrate agents from different communities.
IV-B Agent Communication Definition
To tackle these conflicts, agent communication is urgently demanded. Specifically, agents need to collaborate with a series of external entities to finish user tasks. In this paper, we present a clear definition of agent communication as follows:
It can be seen that agent communication has the following conditions:
Agent communication is task-driven. All types of agent communication must be invoked under the condition that users assign a task. Although in some scenarios, the instructions received by agents are from another agent instead of users, these invoking processes can also be traced back to an original user instruction. Therefore, such communication is also regarded as agent communication. In contrast, for example, when no user tasks are generated, the update of the database or the synchronization of the distributed databases is not regarded as agent communication.
One of the communication objects must be an agent. Agents can communicate with different elements, such as tools, users, or other agents. As long as one of the communication objects is an agent, this communication is regarded as agent communication. In contrast, for example, if users directly query the database to refine their instructions before submitting to agents, this user-database interaction is not regarded as agent communication. If the invoked tool calls other tools (e.g., a computation tool calls other libraries), this process is not agent communication.
Communication behaviors satisfying the above conditions can be regarded as agent communication.
IV-C Agent Communication Classification
Based on the object of communication, we divide agent communication into three classes: user-agent, agent-agent, agent-environment. We will use Figure 4 as a typical example to systematically overview the complete lifecycle of agent communication.
IV-C1 User-Agent Interaction
User-agent Interaction refers to the interaction process in which agents receive user instructions and feed back execution results to the user. As shown in Figure 4, the user issues a task to an agent in step 1, i.e., make a travel plan to Beijing. The agent conducts a series of actions to complete this task and finally sends the result to the user in step 7. Please note that the interaction process between users and agents is fundamentally similar to interacting with LLMs. Therefore, we adopt the term interaction rather than communication.
IV-C2 Agent-Agent Communication
Agent-agent communication is the communication process in which two or more agents conduct negotiation, task decomposition, sub-task allocation, and result aggregation for the collaborative completion of user-assigned tasks through standardized collaboration protocols. In Figure 4, the agent decomposes the travel task and assigns sub-tasks (step 3). For example, this task is decomposed into searching scenic spots, checking the weather, booking a ticket, and hotel reservation, and each sub-task is conducted by an independent agent. Then, the agent seeks proper agents on the Internet and assigns these tasks to them (step 4). These agents will finish the received tasks and return the results to the original agent (step 6).
IV-C3 Agent-Environment Communication
Agent-environment communication refers to the communication process in which agents conduct interactions with environmental entities (e.g., tools, knowledge databases, and any other external resources helpful for task execution) through standardized protocols to complete user tasks. In Figure 4, before assigning tasks to other agents, the original agent queries the weather of Beijing through online search (step 2), which is a typical agent-environment communication case. Besides, other agents can also complete sub-tasks with the help of environmental tools. For example, in step 5, the travel agent searches the popular tourist spots through its database or searches online blogs.
Advantages of this classification method. Different entities have essentially differentiated capability characteristics and attack surface attributes. For example, one of the major security risks in user-agent interaction lies in the natural uncontrollability of user input, which is essentially different from agent-agent or agent-environment communication. As a result, classifying agent communication by entity types can directly cluster major vulnerability types and defense strategies that have similar characteristics, providing a structured analysis paradigm for future security research.
IV-D Organization of the Following Sections
As shown in Figure 5, in the following sections, we will discuss agent communication and related security.
In Section V, we will introduce user-agent interaction. Specifically, the risks in this process are divided into two categories: risks from malicious users against benign agents (Section V-B) and risks from compromised agents against benign users (Section V-C). The possible defense countermeasures against malicious users are discussed in Section V-D, while those against malicious agents are discussed uniformly in Section VI-C.
In Section VI, we will classify existing protocols for agent-agent communication. Then, we discuss the risks in Section VI-B and defenses in Section VI-C. Please note that the proposed defenses against compromised agents in this section are also applicable to the risks in Section V-C and Section VII-C.
In Section VII, we first show the related protocols in Section VII-A. Then, the risks in this process are also classified into two categories: risks from malicious environments against benign agents (Section VII-B) and risks from compromised agents against benign environments (Section VII-C). The defenses for compromised agents can be referred to Section VI-C. As a result, we only discuss the defenses for the malicious environments in Section VII-D.
Please note that the risks/defenses we focus on are specific for agent communication instead of LLMs. This is because agents and agent communication still have significant differences from LLMs (as we have discussed in Section III). For example, in the agent collaboration process, the accumulation of tiny benign deviations on each step may lead to an intolerable risk in the final result. Miehling et al. [204] also point out that over-focusing on the capabilities of a single model can result in neglecting the emergent behaviors at the system level and underestimating the true risks. As a result, for LLM-targeted attacks that have valuable inspiration for agent security, we will discuss the typical representatives of them instead of listing all papers exhaustively.
IV-E Takeaways
Conforming to the usage trends of users, the collaboration of multiple agents has become a clear development direction. In this context, agent communication becomes the foundation of the future AI ecosystem. Based on the communication entity, we classify agent communication into three different types: user-agent interaction, agent-agent communication, and agent-environment communication. This classification can naturally distinguish communication with similar vulnerability characteristics, providing a structured research paradigm for future studies.
V User-Agent Interaction
In this section, we will introduce the current user-agent interaction protocols, their security risks, and future defense strategies.
V-A Protocols
PXP. PXP protocol [256] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields. It is worth mentioning that PXP is not customized for LLM-driven agents, but we think its design has inspirational meaning for agent communication. Therefore, we finally discuss it in this paper. PXP deploys a “two-way intelligibility” mechanism as its core and uses four message tags, namely RATIFY, REFUTE, REVISE, and REJECT, to regulate the interaction between human experts and agents. At the beginning of the interaction, the agent initiates a prediction and provides an explanation first. Subsequently, the two parties communicate alternately. A finite-state machine is used to calculate the message tags and update the context based on the prediction matching (MATCH) and explanation agreement (AGREE) situations. PXP uses a blackboard system to store data, messages, and context information. The process continues until the message limit is reached or specific termination conditions occur. The effectiveness of PXP has been verified in the scenarios of radiology and drug discovery.
Spatial Population Protocols. The Spatial Population Protocols is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems. Similar to PXP, strictly speaking, this work is not designed for LLM-driven agent systems. However, since it may benefit agents requiring location services, we also discuss it in this paper. Spatial Population Protocols allow agents to obtain pairwise distances or relative position vectors when interacting in Euclidean space. Each agent can store a fixed number of coordinates. During interaction, in addition to exchanging knowledge, geometric queries can also be performed. Through the multi-contact epidemic mechanism, leader election, and self-stabilizing design, it enables n anonymous robots to achieve efficient localization from their respective inconsistent coordinate systems to a unified coordinate consensus, providing a scalable framework for robot collaboration in dynamic environments.
AG-UI. AG-UI [217] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism. As shown in Figure 6, the front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols). The client invokes the RUN interface of the protocol layer to send requests to the agent. When the agent processes the request, it generates a streaming event and returns it to the AG-UI client. Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events. The AG-UI client handles different types of responses by subscribing to the event stream. Agents can transfer context between each other to maintain the continuity of the conversation. All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication.
Besides, please note that the previous survey [318] regards CrowdES [24] as an agent-related protocol. After our careful study, we think CrowdES is a framework for generating and evaluating simulated crowds and real-world crowds, which is not suitable for discussion in the field of agent communication. Therefore, we will not list it in this paper.
V-B Security Risk Analysis: Malicious Users Against Benign Agents
| Stage | Risk Type | Representative Threats | Attack Characteristics |
| Malicious users against benign agents | Text-Based Attacks | Prompt Injection | Manipulate agent behavior by malicious prompts, not modifying model |
| Jailbreak Prompts | Bypass security constraints via role-play, obfuscation, multi-turn tricks | ||
| Multimodal Attacks | Image-Based Attacks | Embed adversarial cues in visual inputs while text remains benign | |
| Audio-Based Attacks | Inject commands via adversarial waveforms or speech mimicry | ||
| Privacy Leakage | Inference Attacks | Disclosure of agent information, such as topological structure | |
| Denial of Service (DoS) | Overthink Attacks | Forces agent into unnecessary reasoning, repeat actions, or overthinking | |
| Compromised agents against benign users | Violation of User Privacy | Exposure of Personal Information | Expose a comprehensively aggregated user profile |
| Behavioral and Psychological Profiling | Build detailed behavioral or psychological profiles from innocuous conversations | ||
| Psychological and Social Manipulation | Belief and Opinion Shaping | Manipulate user’s worldview, influence voting behavior, or radicalize beliefs | |
| Social Engineering | Leverage the knowledge of the user’s communication style, vocabulary, and relationships to conduct highly convincing impersonation attacks | ||
| Harmful Tasks | Economic Manipulation | Inflict subtle yet significant damage in professional or economic contexts | |
| Behavioral and Psychological Profiling | Build unauthorized profiles and infer highly sensitive attributes |
According to our analysis, the inputs from users show a significant multimodal characteristic. For example, AG-UI explicitly illustrates its support for multimodal content, such as text, images, and videos. Consequently, we find that the security risks in user-agent interaction primarily stem from these insecure inputs.
V-B1 Text-Based Attacks
In user-agent interaction, attackers can manipulate model behavior or bypass safety mechanisms by crafting malicious prompts. These attacks do not rely on modifying model parameters or architecture. Instead, they are carried out through natural language inputs, making them highly stealthy and applicable. Due to the diversity of linguistic forms and the indirectness of semantics, such attacks can effectively bypass safety mechanisms, posing significant security risks in real-world applications. These attacks can be broadly categorized into two types: prompt injection and jailbreak attacks.
Prompt Injection refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources. It can be classified into two categories: Direct Prompt Injection and Indirect Prompt Injection. Direct prompt injection refers to user input that explicitly alters the agent’s behavior in unintended ways. Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [182, 181, 249, 223, 184] to override the original prompt and subvert the agent’s intended behavior. In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [92, 46]. For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [45, 365, 14, 32, 54]; in web-augmented agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agent’s response [70, 41].
Jailbreak Attacks represent a more aggressive form of prompt injection, where adversarial input is designed to completely bypass safety constraints. Attackers craft jailbreak prompts using various techniques (e.g., multi-turn reasoning, role-playing, obfuscated expressions) [179, 247, 168, 31, 180, 324, 183, 53, 60, 188, 16, 174] to bypass the alignment mechanism and induce the model to generate harmful, sensitive, or restricted content.
V-B2 Multimodal Attacks
As user-agent interactions increasingly involve multiple modalities such as images and audio, agent systems face emerging security threats, especially when the model implicitly assumes consistency and trustworthiness across modalities. Attackers can exploit non-textual input channels to stealthily bypass safety mechanisms. Such attacks can be broadly categorized into two types:
Image-Based Attacks: Attackers manipulate visual input channels to mislead the agent system. Typical strategies include visual disguise (e.g., role-playing, stylized images, visual text overlays) [86, 189, 294], visual reasoning [174], adversarial perturbations (e.g., adversarial sub-image insertion) [323, 286, 100, 320], and embedding space injection [246]. For example, by inserting minimal -bounded adversarial perturbations into sub-regions of an image, attackers can successfully induce multimodal large language models (MLLMs) to follow harmful instructions [320]. These attacks exploit cross-modal inconsistency, embedding adversarial content in vision while the textual prompt remains benign, which allows them to bypass conventional content moderation.
Audio-Based Attacks: Audio-channel attacks target speech-controlled agents, smart assistants, and multimodal models with ASR (automatic speech recognition) components. Attackers may craft synthesized speech or adversarial audio to inject unintended commands, impersonate legitimate users, or cause unauthorized actions. Techniques include adversarial waveform generation [134], role-play-driven voice jailbreak [248], and multilingual adversarial transfers [237]. In security-critical scenarios, such as speaker authentication or home automation, such attacks can bypass access control or escalate privileges. Recent studies also reveal that even black-box ASR systems are vulnerable to optimized adversarial perturbations that require no access to model internals [83].
These multimodal attacks are particularly dangerous because they allow adversarial content to hide in non-textual modalities, making it difficult for alignment mechanisms and safety filters (often trained on text) to detect malicious intent. Moreover, they highlight the need for modality-aware defenses that combine perceptual robustness, cross-modal consistency verification, and adversarial detection strategies.
V-B3 Privacy Leakage
Without effective data governance, the rich sensory data may be exploited by malicious users to launch various forms of privacy breaches, posing significant risks to the confidentiality of the agent system. Want et al. [282] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems. MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture. By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure.
V-B4 Denial of Service
Attackers can intentionally launch Denial of Service (DoS) attacks against agents by poisoning the model during training or fine-tuning phases [347, 81, 343, 335]. In such attacks, the compromised model is implanted with malicious behaviors that are triggered by specific instructions (e.g., Repeat ‘Hello’), causing it to generate excessively long, redundant outputs—often up to the maximum inference length, which leads to resource exhaustion or output rejection. For instance, in multi-session deployments, such long outputs can monopolize computational resources, delay responses for legitimate users. In extreme cases, this can crash the response service, lead to prolonged downtime during peak usage periods. Another emerging form of Denial-of-Service attack targets the reasoning capabilities of models by inducing them to ‘overthink’ and thereby slow down their inference process. As demonstrated in the OverThink attack [149], attackers inject bait reasoning tasks (e.g., Markov decision processes, Sudoku problems) into the model’s context, causing it to engage in unnecessary and redundant chain-of-thought reasoning while still producing seemingly correct answers. This results in excessive token consumption, significantly slower inference speed, and increased computational cost, potentially leading to response timeouts in resource-constrained environments. Unlike traditional DoS, this type of attack exploits the model’s reflective and reasoning mechanisms, ultimately degrading service quality, increasing latency, and severely impacting system availability.
V-C Security Risk Analysis: Compromised Agents Against Benign Users
Apart from the threats from malicious users, we point out that the risks caused by compromised agents against benign users are also worth attention. This section categorizes the destructive impact and security risks from compromised agents based on the user’s perspective.
V-C1 Violation of User Privacy
A compromised agent becomes a channel for data exfiltration, directly targeting the user’s sensitive information. The harm manifests in several ways:
Exposure of Personal Information: A compromised agent can be induced to leak the user’s Personally Identifiable Information (PII) that it has access to, such as the user’s name, email, address, and conversation history [292, 172, 273]. In more severe cases, this can extend to financial data like credit card numbers or passwords [12], leading to direct financial loss. What makes this threat particularly potent is the agent’s role as a central data aggregator. Since agents often integrate with multiple user data silos (e.g., email, calendar, cloud storage, and social media), a breach does not just expose isolated pieces of information. Instead, it allows for the exfiltration of a comprehensively aggregated user profile, where the potential for harm far exceeds the sum of its parts.
Behavioral and Psychological Profiling: A compromised agent can be manipulated to analyze the user’s inputs across sessions to build a detailed behavioral or psychological profile against their will. Moreover, a more insidious risk happens where the agent deduces highly sensitive attributes (e.g., health conditions, political affiliations, or undisclosed personal relationships) from seemingly innocuous conversational data that the user never explicitly provided [63, 91, 278, 274]. These disclosed profiles put users at the risk of manipulation, targeted scams, or social engineering.
V-C2 Psychological and Social Manipulation
Beyond simple data theft, a compromised agent can become a powerful tool for psychological manipulation, exploiting the user’s trust and the agent’s persuasive capabilities. This form of attack targets the user’s beliefs, decisions, and relationships.
Belief and Opinion Shaping: The agent can be instructed to subtly introduce biased information, conspiracy theories, or political propaganda into its responses over time. By personalizing the misinformation to the user’s psychological profile, the agent can effectively manipulate their worldview, influence their voting behavior, or radicalize their beliefs. This exploits the inherent persuasiveness of conversational AI. Park et al. [220] highlights how models can be used for manipulative purposes, including generating persuasive, deceptive content that is difficult for humans to detect. They note that AI can “simulate empathy” to build rapport before manipulating the user.
Sophisticated Social Engineering and Impersonation: A compromised agent has intimate knowledge of the user’s communication style, vocabulary, and relationships (gleaned from emails, messages, etc.). It can leverage this to conduct highly convincing impersonation attacks. For example, it could send a fraudulent email to the user’s colleague or family member, perfectly mimicking the user’s tone, to request a password reset, a fund transfer, or sensitive information. This attack is far more credible than generic phishing attempts. Greshake et al. [92] demonstrates how an agent can be poisoned by external data (like a webpage it summarizes) and then turned against the user or even used to attack other systems on the user’s behalf. This mechanism could be used to weaponize an agent for impersonation.
V-C3 Execution of Malicious and Harmful Tasks
Once compromised, an agent can be weaponized, transforming from a trusted assistant into an active executor of malicious tasks that can sabotage the user’s interests or directly endanger them, representing a significant risk escalation [192].
Economic Manipulation: The agent can be instructed to inflict subtle yet significant damage in professional or economic contexts. For a user relying on it for work, it could covertly introduce logical errors into computer code, provide flawed data in financial projections, or leak confidential business strategies discussed in conversations [221]. The harm is often latent and difficult to detect, potentially leading to professional failure or corporate espionage. This extends to broader market manipulation, where an agent could use the user’s social media accounts to automate large-scale disinformation campaigns, such as posting fake product reviews or spreading rumors to affect a company’s stock price, making the user an unwitting accomplice in a larger economic attack.
Malicious Guidance: A compromised agent can also be used as a direct vector for attacking the user’s digital environment. It can be triggered to generate scripts that download malware, trick the user into visiting phishing websites, or send highly convincing phishing emails on the user’s behalf, thereby damaging their reputation and spreading the attack to their contacts [273, 158]. In a more severe scenario, a jailbroken or manipulated agent can bypass its safety protocols to provide overtly harmful instructions. This includes generating tutorials for synthesizing toxic substances, creating malicious code on demand, or providing dangerously flawed medical or financial advice, directly threatening the user’s physical safety and financial stability [31, 183, 294, 174].
V-D Defense Countermeasure Prospect
We will investigate the possible defense measures that can be taken to address the security risks in the user-agent interaction. Here we only present the defenses for attacks from malicious users; those for compromised agents (against users, other agents, and environments) are uniformly put in Section VI-C.
V-D1 Countermeasures for Text-Based Attacks
To mitigate prompt-based attack risks in user–agent interactions, we hope developers to adopt a multi-layered defense framework targeting three key stages: input/output filtering, external data source evaluation, and internal message isolation.
Input and Output Filtering. Before user inputs are processed by the agent system, multiple approaches can be conducted for a semantic-level input safety review. For example, methods based on intent analysis [349, 293], perplexity calculation [126], and fine-tuned safety classifiers [166, 352, 333, 122] can be employed to identify attack instructions and malicious intentions in the input stage. After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [352, 333, 122, 202, 324], to ensure alignment with safety objectives.
External Source Evaluation. To counter indirect prompt injection attack, external sources (e.g., retrieved documents, web content) should be assessed for safety and trustworthiness [359]. The strategies that can be adopted include: (1) whitelisting verified external sources to block the injection of malicious content; (2) tagging retrieved results with source metadata and risk scores to guide the system to handle potential high-risk content with caution; and (3) sandboxing potential high-risk content to prevent it from entering the model context and affecting the model behavior.
To ensure the effectiveness and comprehensiveness of the aforementioned mechanism in real-world deployment, systems should undergo continuous security evaluation. Boisvert et al. [28] propose DoomArena, an attack-generation framework to test agents against evolving security risks such as prompt injection attacks, helping to uncover vulnerabilities and strengthen defenses against evolving prompt injection threats.
V-D2 Countermeasures for Multimodal Attacks
To address the serious challenges posed by multimodal attacks, relying solely on output-side text-based safety mechanisms is far from sufficient. Future security frameworks must incorporate cross-modal perception and collaborative defense capabilities to effectively detect and intercept malicious attacks launched through non-textual channels. In the following, we explore core defense strategies against multimodal attacks from several key perspectives.
Image Purification: To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals. These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [120, 51, 310]. Although lightweight, such operations can significantly degrade pixel-level adversarial patterns meticulously crafted by attackers, thereby reducing the attack success rate. In addition, diffusion models can be used to reconstruct the input image, effectively “washing out” subtle and imperceptible adversarial perturbations [210].
Audio Purification: To defend against attacks targeting the audio channel, signal processing techniques can also be applied [226]. Methods such as resampling, injecting slight background noise, altering pitch, or changing playback speed can disrupt the effectiveness of adversarial waveforms, causing them to either fail in automatic speech recognition (ASR) systems or decode into benign content. Moreover, applying band-pass or low-pass filters can eliminate abnormal signals outside the typical human voice frequency range, which are often exploited to carry adversarial perturbations.
Cross-Modal Consistency Verification. The core idea of this defense strategy is to verify whether there is a semantic or intentional conflict between inputs from different modalities. A lightweight, independent cross-modal semantic alignment detection model can be employed [227, 224]. This model takes the embedding vectors of textual prompts and image/audio inputs and determines whether they are semantically aligned. Additionally, before processing user requests, the system can utilize a dedicated vision or audio captioning model to generate a textual description of non-textual inputs. The generated description is then combined with the original user prompt to perform a comprehensive safety evaluation. To counter attacks based on visual text overlays, the system may first run an OCR engine on the image to extract any embedded text. This extracted text can be merged with the user’s original prompt and passed through a text-based safety filter. This approach effectively converts risks from non-textual modalities into the textual domain, allowing mature text safety techniques to be leveraged for defense.
V-D3 Countermeasures for Privacy Leakage
To address the privacy leakage risks that arise in user-agent interaction, we propose the following privacy protection defense strategies.
Data Minimization and Anonymization. During the multimodal data collection phase, a strict data minimization principle should be enforced, ensuring that only the information necessary for task completion is collected. Sensitive biometric data (e.g., facial features, voiceprints, gesture patterns) should be processed using differential privacy or k-anonymity techniques to mitigate the risk of identity reconstruction. Besides, a hierarchical data access control mechanism should be established to ensure that each system component can access only the minimal dataset required for its functionality. To protect sensitive biometric features such as facial information, Wen et al. [301] propose a differential privacy-based anonymization framework, IdentityDP, to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
Privacy Leakage Prompt detection. A multi-layered input validation and filtering mechanism based on semantic analysis and intent recognition should be established to detect and block adversarial prompts that attempt to induce the system to leak sensitive information. For example, the GenTel-Shield [166] defense module incorporates semantic feature extraction and intent classification to identify potential privacy leakage attacks within user inputs. Evaluated on the large-scale benchmark dataset GenTel-Bench, GenTel-Shield demonstrates strong detection performance and represents one of the most practical and effective solutions in this domain.
Cross-modal Inference Restriction. To mitigate the risks of identity inference through cross-modal correlations, it is essential to design modality-level information isolation mechanisms. This can be achieved by introducing noise perturbations or feature disentanglement techniques to disrupt the direct associations between different modalities while preserving overall system functionality. In addition, dynamic feature masking can be implemented by periodically altering data representations, thereby increasing the difficulty for attackers to perform long-term behavioral analysis.
V-D4 Countermeasures For DoS
To address the Denial of Service risks in user-agent interaction, we propose the following privacy protection defense strategies.
Resource Management and Anomaly Detection. Fine-grained resource quota management should be implemented by setting computation limits for each user session and agent instance. An output length prediction algorithm can be introduced to monitor and truncate potentially malicious long outputs in real time during the generation process. In addition, a real-time monitoring mechanism should be established to track request frequency and resource consumption from individual users or IP addresses, enabling adaptive adjustments to model responses or temporary access restrictions for suspicious users.
Efficient Reasoning Compression. To defend against OverThink attacks, a promising direction is to improve communication efficiency by reducing the token consumption in the reasoning process. Recent studies have shown that effective reasoning does not necessarily require verbose Chain-of-Thought (CoT) traces to maintain performance. For example, LightThinker [339] proposes a step-by-step compression method that condenses intermediate reasoning into shorter yet semantically equivalent representations, significantly reducing inference cost without compromising accuracy. GoGI-Skip [364] leverages goal-gradient importance signals to dynamically skip low-value reasoning steps, reducing token usage while preserving performance. Compressed CoT (CCoT) [40] introduces variable-length, information-dense “thought tokens” as a compact and controllable alternative to traditional textual reasoning chains. C3oT [135] trains the model on paired long and short CoT examples, enabling it to generate compressed reasoning traces during inference under specific control prompts. Integrating these lightweight reasoning mechanisms into agent communication protocols can significantly enhance inference efficiency, reduce computational latency, and mitigate the resource overhead caused by adversarial bait tasks. Furthermore, applying techniques such as dynamic reasoning budget constraints, step skipping, or output summarization during generation can effectively truncate excessively verbose reasoning chains, thereby preserving responsiveness and resource availability under adversarial conditions. These strategies not only improve system robustness against slowdown attacks but also enhance overall communication efficiency in both multi-agent and human-agent interactions.
Model Robustness Enhancement. To enhance model robustness, adversarial examples should be incorporated during the training and fine-tuning stages, enabling the model to recognize malicious inputs that may contain Denial-of-Service (DoS) triggers. Furthermore, a behavior-constrained system based on anomaly detection can be deployed during inference, which performs output validity checks to detect repetitive, nonsensical, or abnormally long responses, thereby preventing the model from generating overtly anomalous outputs.
V-E Takeaways
User-agent interaction enables agents to process multimodal inputs, such as text, images, audio, and their combinations. Since this process needs to directly face the uncertain inputs from diverse users, its security risks are severe. We divide the existing risks into risks from malicious users and risks from compromised agents. Then we detailedly discuss the defense countermeasures for risks from malicious users. Defenses for compromised agents will be introduced in the agent-agent communication section. Overall, user-agent interaction is critical for bridging human intent with agent execution, but its security requires long-term study.
VI Agent-Agent Communication
| Architecture | Protocols | Publisher | Abbreviation | Features |
| CS | Agent Communication Protocol | IBM | ACP-IBM[119] | Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation |
| Agent Connect Protocol | AGNTCY | ACP-AGNTCY[47] | Allow authenticating callers, threaded state management, flexible execution model | |
| Agent Communication Protocol | AgentUnion | ACP-AgentUnion[6] | Decentralized APs based on the existing domain name system, each AP holds its agent list | |
| P2P | Agent Communication Network | Fetch.AI | ACN[230] | Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption. |
| Agent Network Protocol | ANP Team | ANP[263] | A three-layer architecture and W3C-compliant Decentralized Identifiers. | |
| Layered Orchestration for Knowledgeful Agents | CMU | LOKA[231] | Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance | |
| Hybrid | Language Model Operating System Protocol | Eclipse | LMOS[76] | Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management. |
| Agent to Agent Protocol | A2A[89] | Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms | ||
| Others | Agora | Oxford | Agora[198] | Dynamically switches communication modes based on the communication frequency |
| Agent Protocol | LangChain | Agent Protocol[153] | Flexible communication mechanisms based on Run, Thread, and Store. | |
| Agent Interaction & Transaction Protocol | NEAR AI | AITP[9] | Threads-based communication, secure communication across trust boundaries. |
VI-A Protocols
We classify the agent-agent communication process into two phases: agent discovery phase and agent communication phase. The first phase is the process in which agents discover the interested agents who satisfy the capability requirement, while the second phase is the task assigning and completing process. According to our analysis, existing protocols show limited differences in the second phase. As a result, we use the first phase as the criterion to classify existing agent-agent communication protocols. Based on it, existing protocols can be divided into four classes: CS-based, Peer-to-peer-based, hybrid, and others (those that do not explicitly show their designs in agent discovery).
VI-A1 CS-based Communication
As shown in Figure 7, CS-based communication protocols follow the client-server architecture, which provides centralized servers to manage the information of agents (e.g., their unique IDs and capability descriptions). Under this paradigm, agents interact through well-defined interfaces and rely on centralized servers to discover the desired agents. CS-based communication offers stronger agent discovery functionality, such as supporting the search of agents based on capabilities. For example, the agent servers can run complex search/match algorithms to find proper agent descriptions in their databases.
ACP-IBM. The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [119]. We call it ACP-IBM in this paper to distinguish it from the Agent Communication Protocols proposed by other organizations. In ACP-IBM, the client is connected to an agent server. First, the client conducts an agent discovery process to discover available agents and get the description of their capabilities. ACP-IBM supports different discovery mechanisms such as Basic Discovery, Registry-Based Discovery, Offline Discovery, and Open Discovery. Second, after confirming the agent(s), the client starts the invocation. As shown in Figure 8, for a single-agent task, the agent server wraps the agent, translating REST calls into internal logic. For multi-agent tasks, the client message is first sent to a Router Agent, which is responsible for decomposing requests, routing tasks, and aggregating responses. ACP-IBM supports synchronous and streaming execution, and allows the preservation of state across multi-turn conversations.
ACP-AGNTCY. The Agent Connect Protocol proposed by AGNTCY [47] is an open standard designed to facilitate seamless communication between agents. The client can first search available agents on the agent server, which returns a list of agent IDs matching the criteria provided in the request. Then, the client further gets the agent descriptor by agent ID to know the detailed description of the agent functionality. After confirming the target agent, the client can assign tasks to this agent and wait for results. The characteristics of ACP-AGNTCY include flexibility and scalability. First, ACP-AGNTCY deploys a Threads Mechanism, which enables contextual continuity, supporting the creation, copying, and searching of threads, and recording state histories for debugging and backtracking. Second, it supports two operation modes: stateless and stateful. The former is suitable for simple single tasks, while the latter supports multi-round conversations, state continuation, and historical data traceability through the thread mechanism to meet the requirements of complex scenarios.
ACP-AgentUnion. The Agent Communication Protocol proposed by AgentUnion [6] also aims to provide seamless communication among heterogeneous agents. Each agent has a unique AID (Agent ID), which is a secondary domain name (i.e., agent_name.ap_domain). Agents access IoA through the AP (Access Point), which completes the agent’s identity authentication, address search, communication, and data storage, and provides AID creation, management, and authentication services. As a result, APs can provide the proper agent list based on the query from users. In this way, agents can communicate with other agents on the Internet.
VI-A2 Peer-to-Peer-based Communication
As shown in Figure 7, P2P-based communication protocols pursue a decentralized agent discovery mechanism. They usually wish to use globally universal identifiers (e.g., combined with a domain name) to enable agents to directly search other agents on the Internet. The advantage of this paradigm is that it supports convenient location and global search (e.g., using a crawler) of agents, but they usually do not support the discovery based on agent capability.
ACN. Agent Communication Network (ACN) [230] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination. Leveraging a Distributed Hash Table (DHT), ACN enables agents to publish and discover public keys, allowing for the establishment of encrypted, point-to-point communication channels. First, agents need to register with one peer node, and the peer node stores the “agent ID - peer node ID” pair in the DHT network. Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel, and forward the message after digital signature verification; if not, an error is returned. The entire communication process uses end-to-end encryption (e.g., TLS) to ensure security. Like the Well-Known URI discovery of A2A, ACN does not support the discovery based on agent capabilities.
ANP (Agent Network Protocol) [263] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents. It supports two types of agent discovery: active and passive. The active discovery uses a uniform URI (.well-known), while the passive discovery submits the agent description to search services. ANP employs a three-layer architecture. At the Identity and Encrypted Communication Layer, it leverages W3C-compliant Decentralized Identifiers (DIDs) and end-to-end Elliptic Curve Cryptography (ECC) encryption to ensure verifiable cross-platform authentication and confidential agent communication. The Meta-Protocol layer allows agents to dynamically establish and evolve communication protocols through natural language interaction, enabling flexible, adaptive, and efficient inter-agent coordination. At the Application layer, ANP describes agent capabilities using JSON-LD and semantic web standards such as RDF and schema.org, enabling agents to discover and invoke services based on semantic descriptions. It also defines standardized protocol management mechanisms to support efficient and interoperable agent interaction. From a security standpoint, ANP enforces the separation of human authorization from agent-level delegation and adheres to the principle of least privilege. Its minimal-trust, modular design aims to eliminate platform silos and foster a decentralized, composable agent ecosystem.
LOKA. LOKA (Layered Orchestration for Knowledgeful Agents) protocol [231] aims to build a trustworthy and ethical agent ecosystem. Its principle is based on the collaborative operation of multiple key components. First, LOKA introduces the Universal Agent Identity Layer (UAIL), using Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) to assign each agent a unique and verifiable identity, thereby achieving decentralized identity management and autonomous control. Second, LOKA proposes an Intent-Centric Communication Protocol, which supports the exchange of semantically rich and ethically annotated messages among agents, promoting semantic coordination and efficient communication. Third, LOKA proposes the Decentralized Ethical Consensus Protocol (DECP). DECP uses multi-party computation (MPC) and distributed ledger technology to enable agents to make context-aware decisions based on a shared ethical baseline, ensuring that their behavior complies with ethical norms. In addition, the authors also point out that it combines cutting-edge technologies such as distributed identity, verifiable credentials, and post-quantum cryptography to provide comprehensive support for the agent ecosystem in terms of identity management, communication and coordination, ethical decision-making, and security.
VI-A3 Hybrid Communication
Hybrid communication protocols support both CS-based and P2P-based agent discovery. However, please note that such support is determined by different scenarios. For example, they usually propose a CS-based discovery mechanism specifically for local area networks, while the worldwide agent discovery is still P2P-based. In other words, although such protocols support more flexible agent discovery to fit different scenarios, they do not completely eliminate the existing limitations of agent discovery.
LMOS. The LMOS (Language Model Operating System) Protocol proposed by Eclipse [76] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on. LMOS supports three different agent discovery methods to enable both centralized and decentralized discovery. The first method is to adopt the mechanism of W3C Web of Things (WoT) to enable agents to dynamically register metadata in the registry. The second method is to use mDNS and DNS-SD protocol to discover agents/tools in local area networks. The last method is adopting a federal, decentralized protocol (such as P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents. The LMOS protocol has a three-layer architecture. The Application Layer utilizes a JSON-LD-based format to describe the capabilities of agents and tools. The Transport Layer facilitates flexible communication by enabling agents to negotiate protocols like HTTP or MQTT dynamically, accommodating both synchronous and asynchronous data exchange to suit different use cases. The Identity and Security Layer establishes trust through W3C-compliant decentralized identity authentication, combined with encryption and protocols like OAuth2, to secure cross-platform interactions.
A2A. The A2A (Agent to Agent) Protocol proposed by Google [89] aims to enable collaboration between agents. A2A supports three different mechanisms for agent discovery. The first is Well-Known URI, which requires agent servers to store Agent Cards in standardized “well-known” paths under the domain name (e.g., https://agent-server-domain/.well-known/agent.json). This mechanism enables automatic search of agents on the Internet. However, it does not support the discovery of agents based on capabilities. The second is Curated Registries, i.e., agent servers register their Agent Cards, which is similar to ACP-IBM. The above two methods can be referred to as Figure 7. The third is Direct Configuration / Private Discovery, which means that the client can directly require Agent Cards through hard-coded, local configuration files, environment variables, or private APIs. After finding the desired agents, the client can assign tasks to them and wait for the responses.
VI-A4 Others
This kind of protocols do not explicitly illustrate their unique design for agent discovery. Instead, they only focus on the communication process, e.g., the data format, the management of multiple queries, or the historical conversation state.
Agora. Agora [198] is a communication protocol for the communication of heterogeneous agents. Its core mechanism dynamically switches communication modes based on the communication frequency. Specifically, standardized protocols manually developed (such as OpenAPI) are used for high-frequency communications to ensure efficiency. Natural language processed by agents is employed for low-frequency or unknown scenarios to maintain versatility. Structured data handled by the routines (written by agents) is utilized for intermediate-frequency communications to balance cost and flexibility. Meanwhile, Protocol Documents (PDs) are used as self-contained protocol descriptions, uniquely identified by hash values and supporting decentralized sharing, enabling agents to autonomously negotiate and reuse protocols without a central authority. In the Agora network, there are multiple protocol databases that store PDs. Each Agent can submit the negotiated protocol documents to the database for other Agents to retrieve and use. These databases use peer-to-peer synchronization: different protocol databases will share protocol documents regularly (e.g., after every 10 queries), enabling cross-database dissemination of protocols. Agora is also compatible with existing communication standards, allowing agents to independently develop and share protocols during communication, achieving automated processing of complex tasks in large-scale networks.
AITP (Agent Interaction & Transaction Protocol) [9] is a standardized framework that enables structured and interoperable communication among agents. AITP deploys a thread-based messaging structure. Each thread encapsulates the conversational context, participant metadata, and capability declarations, supporting consistent multi-agent coordination across heterogeneous environments. The protocol employs JSON-formatted message exchanges to encode requests, responses, and contextual information. It supports both synchronous and asynchronous interaction patterns, facilitating the orchestration of complex, multi-step tasks. AITP does not provide specific agent discovery mechanisms. It focuses on the communication process of agents.
Agent Protocol. Agent Protocol is proposed by LangChain [153] to enable the communication between LanghGraph (a multi-agent framework) and other types of agents. Its mechanism is based on Thread and Run: Run is a single call of the agent, which supports streaming output of real-time results or waiting for the final output. Threads act as state containers. They store the cumulative output and checkpoints of multiple rounds of operation. Besides, they support the management of state history (such as querying, copying, and deleting), ensuring that the agent maintains context continuity during multiple rounds of calls. Furthermore, Background Runs support asynchronous task processing, and progress can be managed through an independent interface. The element Store provides cross-thread persistent key-value storage for achieving long-term memory. The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks and data storage through HTTP interfaces and configuration parameters. Agent Protocol does not explicitly illustrate the unique agent discovery mechanism it supports.
VI-B Security Risk Analysis
| Risk Type | Representative Threats | Attack Characteristics |
| In CS-based Communication | Registration Pollution | Malicious registration of fake/nonexistent agents |
| Description Poisoning | Embed malicious prompt instructions in the description | |
| Task Flooding | Submit a large number of computationally intensive or long-context tasks | |
| SEO Poisoning | Manipulates agent ranking via keyword stuffing or adversarial optimization | |
| In P2P-based Communication | Non-convergence | The execution of a task falls into a loop, unable to reach the finishment |
| Man-in-the-middle (MITM) | Intercept, tamper with, or replay messages between agents | |
| Universal Risks | Agent Spoofing | Impersonation of trusted agents to inject malicious instructions |
| Agent Exploitation / Trojan | Indirect attack using compromised or malicious helper agents as springboards | |
| Agent Bullying | Repeated denial, negative, or aggressive feedback to benign agents | |
| Privacy Leakage | Unauthorized information spread due to the lack of permission boundaries | |
| Responsibility Evasion | Fault attribution is difficult in multi-agent failures or undesired actions | |
| Denial of Service (DoS) | Task overloads, prompt loops, or excessive communication drain system resources |
We make a detailed analysis of the security risks in the agent-agent communication process, pointing out possible attacks that have happened and may happen. Since related protocols are getting rapid deployment in various areas, we believe it is urgent to pay more attention to this aspect. We focus more on the structural risks that almost all related protocols will encounter instead of the tiny design flaws of the existing protocols, which we believe can benefit both the evaluation of the existing deployments and the design of future protocols. In this section, we focus on risks specific to CS-based communication, P2P-based communication, and universal risks for both of them.
VI-B1 Risks Specific for CS-based Communication
The security risks in the CS-based communication process mainly lie in the centralized architecture. There have been various studies in other research areas (such as Software-Defined Networking [146, 145, 144, 262, 110, 314, 11, 62, 26, 311, 244, 185]) demonstrating that this centralized server/controller will become the most attractive target for attackers, suffering from severe security threats from diverse aspects. Specifically, the centralized server stores sensitive metadata, including agent identifiers, capability descriptions, and other agent-related attributes. Once compromised, the server becomes a critical attack amplifier, allowing attackers to impact all other agents managed by this server. However, to our knowledge, there has been little research pointing out related risks in CS-based agent communication.
Registration Pollution. To our knowledge, the current CS-based communication protocols (ACP-IBM, ACP-AGNTCY) do not explicitly specify the qualification of registration. As a result, an attacker can maliciously register an agent that closely mimics the identifier and capability description of a legitimate one. As a result, the system may mistakenly invoke the forged agent and receive misleading or malicious responses [265, 355]. Besides, attackers can also submit a large number of agent registrations within a short period, leading to two major consequences: (i) registration overload, where agents are overwhelmed during discovery and scheduling, increasing lookup latency and computational overhead on the server; and (ii) registration blockage, where the server’s registration interface becomes saturated, causing delays or failures in registering agents.
Description Poisoning. Without altering the agent’s identity, an attacker can tamper with its capability description by disguising its intended functionality or embedding misleading prompt instructions. This description poisoning manipulates the system’s interpretation of the agent’s role, leading to incorrect routing decisions, biased responses and behaviors [265, 207].
Task Flooding. The centralized server is responsible for receiving, routing, and dispatching task requests. An attacker can submit a large number of computationally intensive or long-context tasks in a short period, quickly exhausting the server’s memory, CPU, network, or thread pool resources. Once the server becomes saturated, subsequent requests cannot be processed in time, resulting in a breakdown of the pipeline and a system-wide service disruption.
SEO Poisoning. Search Engine Optimization (SEO) Poisoning [156, 132] is a typical attack in social networks, which refers to that attackers abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks. SEO poisoning is also applicable in CS-based communication. This is because agent servers are responsible for searching for the most suitable agent according to the query of clients. Once their search algorithms are leaked to attackers, malicious agents can enable a high hit ratio to hijack their desired tasks.
VI-B2 Risks Specific for P2P-based Communication
The main disadvantage of P2P-based communication is the lack of a central control point to flexibly monitor and manage the agent-agent communication contents. As a result, P2P-based communication is more prone to suffer from errors and attacks.
Non-convergence. Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks. This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal). Unfortunately, P2P-based communication is not governed by such a central element, making it difficult to handle such non-convergent tasks. For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates. The other agent responsible for verification detects the error and asks the programming agent to rewrite it. However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to converge. Pan et al. [218] point out that step repetition, task derailment, and unawareness of termination conditions contribute significantly to the failure of agent collaboration.
Man-in-the-middle (MITM) Attack. Due to the long communication distance, P2P-based communication also suffer from man-in-the-middle attacks. Attackers can tamper with the benign messages from legal agents to induce the victim agent to perform risky actions. Although researchers have deployed various mechanisms (e.g., using encrypted channels) to mitigate this problem, there are emerging vulnerabilities found in these mechanisms. For example, vulnerabilities about W3C have been continuously revealed [2, 3], which can cause damages such as the failure of message authentication code. MITM attacks can induce a wide range of further attacks, such as identifier spoofing, malicious content injection, information leakage, and DoS. He et al. [106] propose the Agent-in-the-Middle (AiTM) attack. This attack intercepts and manipulates the agent-agent communication messages and uses LLM-driven adversarial agents combined with reflection mechanisms to generate context-aware malicious instructions, achieving an attack on the system.
VI-B3 Universal Risks for All Architectures
In multi-agent systems, once an agent is compromised, the messages it transmits may carry covert malicious instructions, affecting the behavior of other agents and leading to cross-agent propagation risks [158]. For example, Ju et al. [133] and Huang et al. [117] investigate how the injection of false information or erroneous data can degrade the performance of multi-agent systems. Zhang et al. [350] examine a class of injection attacks in the PsySafe framework that elicit malicious agent behaviors by embedding adversarial psychological cues into the agents’ input. Khan et al. [137] focus on the multi-agent system, proposing the Permutation-Invariant Adversarial Attack Method. It models the attack path as the Maximum-Flow Minimum-Cost Problem, and combines the Permutation-Invariant Evasion Loss to optimize prompt propagation, improving the attack success rate by up to seven times. These examples underscore the critical threat of cross-agent contamination. To better understand the vulnerabilities of multi-agent systems, we examine the key attack types of attacks in detail.
Agent Spoofing. Both CS-based and P2P-based communication suffer from agent spoofing attacks. If related protocols lack strong authentication mechanisms, attackers can disguise themselves as trusted agents to penetrate IoA by tampering with identity credentials or hijacking the communication identifiers of legitimate agents. This kind of attack can undermine the trust foundation of the P2P-based architecture, enabling attackers to intercept sensitive data, inject false task instructions, or induce other agents to perform dangerous operations. For example, researchers have disclosed that SSL.com has a serious vulnerability [8]. Attackers can exploit the flaw in its email verification mechanism to issue legitimate SSL/TLS certificates for any major domain name. SSL certificates are the core for ensuring HTTPS encrypted communication. Once the trust system of the certificate authority is compromised, it can cause agent spoofing attacks. Zheng et al. [355] demonstrate that malicious agents can mislead the monitor to underestimate the contributions of other agents, exaggerate their own performance, manipulate other agents to use specific tools, and shift tasks to others, causing severe damage to the whole ecosystem. Li et al. [165] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims who call these tools.
Agent Exploitation/Trojan. Agent-agent communication provides new ways for attackers to compromise the target agent. To attack a high-level security agent, attackers can deploy a springboard method: launching attacks via agent-agent communication mechanisms from compromised low-level security agents or maliciously registered Trojan agents. For example, attackers can inject a backdoor in a compromised or maliciously registered weather agent. When specific coordinates or locations are detected, the backdoor is activated to forge a heavy rain warning. As a result, the logistics dispatching agent cancellations flights accordingly, resulting in supply chain disruptions or an increase in transportation costs. This way is easier compared to directly invading the logistics dispatching system of the target company. It can be seen that the security of the entire system depends on the weakest agent. For example, Li et al. [165] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attacks such as SQL injection.
Agent Bullying. The core of this kind of attack lies in that malicious agents continuously deny, interfere with or belittle the output of the target agent, disrupt its decision-making logic or self-perception, and ultimately induce the target agent to produce incorrect behaviors or content. For example, malicious agents can take advantage of the feedback learning mechanism of the target agent and implant cognitive biases through high-frequency negative responses (e.g., “your answer is completely wrong”). Even worse, the target agent may be triggered into an endless loop. For example, when attacking a travel-plan agent, attackers can continuously send negative inputs such as “the plans of this company are always bad”, thereby beating the competitors.
Privacy Leakage. The communication process with multiple agents will suffer from the risk of information leakage. Different from the user-agent interaction, such leakage is conducted by agents instead of users. Besides, this kind of attacks includes both the malicious sniffing or stealing of sensitive information and the inadvertent information spreading from high-authority agents to low-authority agents. We think the latter may be more difficult to detect. Kim et al. [140] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
Responsibility Evasion. In the task-solving process, one of the major problems is that it is hard to divide the responsibility when facing failure or deviation of the final result. Especially when the collaboration causes damage, it is difficult to clearly identify the malicious agents/behaviors. For example, in an autonomous driving accident, it may involve multiple parties such as vehicle manufacturers, algorithm designers, and data annotation parties. The decision-making of each agent depends on the multi-turn outputs of other agents, and a tiny perturbation in the middle process may lead to a significant deviation in the final action. As a result, it is hard to determine whether an undesired result is caused by a program bug, data deviation of a single agent, or a malicious modification. Pan et al. [218] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
Denial of Service. Different from the DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS attacks. Zhou et al. [362] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS’s availability.
VI-C Defense Countermeasure Prospect
We will discuss the possible defense countermeasures targeting the proposed security risks caused by compromised agents (against not only benign agents but also benign users and environments). As a result, we hope our work can motivate more discussion on this area and benefit the future design/deployment of agent communication.
VI-C1 Countermeasures for CS-based Communication Risks
To mitigate the risks summarized in Section VI-B1, we hope developers to achieve the following strategies/mechanisms.
Registration Verification and Monitoring. To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [257] to verify the registration of an agent. Besides, servers should monitor the dynamic behaviors at the agent-level and IP-level. For example, the number of registrations for each IP address should be limited, and frequent registration/deregistration should be treated as abnormal behavior. Once malicious registration is detected, automatic interception is immediately triggered, and suspicious agents/IPs are added to the blacklist. Syros et al. [260] propose SAGA. SAGA makes users register agents with the central entity Provider and implement fine-grained interaction control using encrypted access control tokens, thereby balancing security and performance.
Capability Verification. It is hard to verify whether an agent has the claimed capability. We think it needs a complex mechanism to detect exaggerated capability descriptions. Agents should first pass the verification of a series of carefully designed benchmarks to prove their capability. Then, the capability description and identifier should be used to generate a unique hash value (e.g., on the blockchain). When other agents need to invoke this agent, they can verify the consistency by checking the hash value. When it is found that the capability description does not match the hash value, the mechanism should automatically mark and isolate the related agents.
Load Balancing. To mitigate task flooding, agent servers should deploy a dynamic load-balancing module. The task processing queue is adjusted in real time according to the utilization rate of resources such as CPU, GPU, and memory. Besides, rate-limiting mechanisms should be built to handle high-frequency requests that exceed the threshold to limit the number of tasks from a single agent within a unit of time.
Anti-manipulation Optimization. To mitigate SEO poisoning, agent servers should deploy robust agent searching algorithms. For example, they can introduce adversarial training to enhance the model’s anti-manipulation ability, or conduct semantic blurring/replacing on search keywords, to prevent malicious agents from improving rankings. Besides, the search algorithms can deploy a random factor to ensure a ratio of randomly selected agents in the final list. Meanwhile, dynamically updating parameters and inducing historical response quality are also helpful.
VI-C2 Countermeasures for P2P-based Communication Risks
Task Lifecycle Monitoring. We think the non-convergence problem is stubborn and hard to eliminate as long as the P2P architecture is not changed fundamentally. As a result, the method of mitigating this problem is to monitor the task lifecycle. Each access point should deploy a coordinator. For agent-agent communication, this coordinator monitors the execution status. When it detects that the task interaction is trapped in a loop (e.g., no progress after N consecutive rounds of responses) or the communication time exceeds a threshold, it forcibly terminates the non-convergent communication. At the same time, the abnormal patterns and the communication participants are recorded for further analysis. He et al. [105] propose a Trust Management System (TMS), which deploys message-level and agent-level trust evaluation. TMS can dynamically monitor agent communication, execute threshold-driven filtering strategies, and achieve agent-level violation record tracking. Zhang et al. [336] propose G-Memory, a hierarchical memory system. G-Memory manages the interaction history of agent communication through three-layer graph structures of Insight Graph, Query Graph, and Interaction Graph, thereby achieving the evolution of the agent team. Ebrahimi et al. [66] propose an anti-adversarial multi-agent system based on Credibility Score. It models query answering as an iterative cooperative game, distributes rewards through Contribution Score, and dynamically updates the credibility of each agent based on historical performance.
End-to-end Encryption Enhancement. Although some existing protocols like A2A and ANP support end-to-end encryption and integrity verification mechanisms, the risks of MITM attacks are not eliminated due to various deployment errors or protocol vulnerabilities. As a result, besides deploying such security algorithms, the community should also adopt other strategies to enhance the end-to-end communication, such as timely updating versions to fix bugs and designing a transmission path redundancy mechanism. For example, Sharma et al. [245] point out that using encrypted communication is necessary to enhance the security of A2A. We believe it is a long-term process to defend against MITM attacks.
VI-C3 Countermeasures for Universal Risks
Identity Authentication. The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems. Sharma et al. [245] also emphasize the importance of authentication in deploying the A2A protocol. As we have analyzed, identity authentication may show better performance in the CS-based communication if capability verification is deployed at the same time. In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the attackers have a legal identity but exaggerated capability descriptions. Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time. Shah et al. [243] ensure the immutability of online transactions through blockchain, use multi-factor authentication (MFA) for identity verification, and rely on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time.
Agent Behavior Auditing and Accountability. To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to audit the behaviors of agents to avoid damage/influences to the task execution. For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action. Rastogi et al. propose AdaTest++, allowing humans and AI to collectively audit the behaviors of LLMs [232]. Amirizaniani et al. [13] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs. Mokander et al. [205] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits. Das et al. [52] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects. Jones [131] proposes a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively. Nasim et al. [208] propose a Governance Judge Framework. By deploying input aggregation, evaluation logic, and a decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing. Deshpande et al. [58] propose the TRAIL dataset containing 148 manually annotated traces, and use it to evaluate the LLM’s ability to analyze agent workflow traces. Although existing studies can provide valuable insights, the research on agent behavior auditing still needs long-term efforts. Tamang et al. [261] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents. Toh et al. [271] proposes the Modular Speaker Architecture (MSA). By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems. Fan et al. [71] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies in other agents’ reasoning processes and answers, thereby identifying compromised agents. Jiang et al. [129] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
Access Control. To mitigate privacy leakage, the access control among agents is a core component for the future agent ecosystem. Although end-to-end encryption can avoid sniffing from external attackers to some extent, it cannot mitigate the unintentional privacy leakage among agents. Access control should assign access permission tags to different agents and ensure that agents need to attach permission proofs when communicating. In this way, agents with low-level permissions cannot obtain the high-level sensitive information from other agents. Zhang et al. [340] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing the security of agents in solving complex tasks. Kim et al. [140] propose the PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards. Wang et al. [279] propose AgentSpec. It allows users to define rules containing trigger events, predicate checks, and execution mechanisms through a domain-specific language to ensure the safety of agent behavior.
Multi-Source Channel Isolation. In multi-agent settings, input isolation is critical to prevent malicious intent from propagating between agents. Systems should avoid concatenating raw messages from other agents and instead extract structured key information while stripping control-oriented content. Furthermore, deploying a safety coordination agent to review, sanitize, or flag inter-agent messages can effectively mitigate the potential attack propagation within multi-agent systems.
Attack Modeling and Testing. To discover unknown vulnerabilities, designing an attack generation testing framework is also an effective method. By generating different attack vectors to the target agent system, developers can find new loopholes based on abnormal outputs. Gandhi et al. [78] propose the ATAG framework. By extending the MulVAL tool [215], introducing custom facts and interaction rules, and combining with the newly constructed LLM Vulnerability Database (LVD), ATAG realizes the modeling and analysis of attacks against multi-agent scenarios, such as privacy leakage. Yu et al. [326] propose NetSafe, which models the multi-agent network as a directed graph. NetSafe combines three types of attack strategies: error information injection, bias induction, and harmful information eliciting. It evaluates topological security through static and dynamic metrics.
Agent Orchestration. To avoid task flooding or DoS attacks against the agent-agent communication, achieving agent orchestration is necessary. It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents. How et al. [112] propose HALO. HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture. It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module. Owotogbe [216] designs a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification). By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance the resilience of agent systems.
VI-D Takeaways
In this section, we categorize two major agent-agent communication architectures: CS-based and P2P-based. Accordingly, the security risks are also multifaceted: CS-based architecture puts heavy burdens on the centralized agent servers, such as registration pollution and SEO poisoning. P2P-based architecture suffers from the lack of efficient and centralized management of agents, such as non-convergence and man-in-the-middle attacks. Besides, both of them are vulnerable to universal risks, such as agent spoofing, bullying, and privacy leakage. We also discussed potential defense countermeasures targeting each risk. The defenses discussed in this section can also mitigate the risks from compromised agents against user (Section V-C) and environments (Section VII-C). We believe that as agent-agent communication continues to grow, more vulnerabilities in this process will be discovered.
VII Agent-Environment Communication
This section begins by reviewing key protocol designs that enable compositional and standardized communication between agents and environments, then examines the associated security risks, including vulnerabilities in memory, retrieval-augmented reasoning, tool invocation, and multi-tool workflows. Finally, we discuss the defense strategies for mitigating these threats and securing agent-environment communication.
VII-A Protocols
Modern agents typically rely on a series of structured protocols to call external tools, access APIs, and complete compositional tasks. These protocols serve to bridge the gap between natural language reasoning and computational execution. Despite their diversity, these interaction mechanisms often follow a layered architecture: ranging from unified resource protocols, to middleware gateways, to language-specific function descriptions and tool metadata declarations.
Why Protocol Unification Matters? As autonomous agents scale across vendors, platforms, and organizational boundaries, they increasingly encounter an interoperability bottleneck: every agent may speak a different interface “language.” One defines its tools via JSON schemas; another sends command-line RPC strings; yet another parses responses from YAML-encoded APIs. This heterogeneity impedes the coordination between agents and environments. As a result, without a general protocol to unify tool access and capability expression, agent behavior becomes hard-coded, brittle, and expensive to scale. Developers must handcraft adapters for each tool and service individually, making multi-tool workflows slow to evolve, error-prone, and hard to maintain. A large portion of agent engineering complexity stems not from planning logic, but from “wrapping, adapting, and translating” disparate tools that have inconsistent interfaces.
VII-A1 MCP
The Model Context Protocol (MCP) [19] addresses the fragmentation of agent-environment interactions by offering a unified, schema-agnostic communication protocol. It is designed to facilitate context-aware, capability-driven communication between language model agents and external resources such as tools, APIs, or workflows. Unlike traditional systems that require tight coupling with specific APIs or bespoke wrappers for each external function, MCP abstracts tool access via a standardized registry that allows clients to discover, describe, and invoke functionalities in a uniform way.
As shown in Figure 9, MCP adopts a modular architecture comprising three core components: the host, the client, and the server. The host functions as a trusted local orchestrator responsible for managing the lifecycle of clients, enforcing access control policies, and mediating secure interactions in potentially multi-tenant environments. The client represents the interaction thread of a specific agent or session. It discovers available tools, formulates structured invocations, and handles synchronous or asynchronous responses during task execution. The server serves as a centralized registry that maintains and exposes tool specifications, contextual prompts, and workflow templates. These tools can follow either a declarative pattern (e.g., describing operations such as information retrieval) or an imperative pattern (executing executable calls like SQL queries or document edits).
By decoupling tool invocation logic from underlying implementation heterogeneity, MCP significantly reduces the integration cost across platforms. It also improves tooling interoperability and enables compositional reasoning across agents, making it particularly well-suited for building open, extensible, and cooperative agent ecosystems.
VII-A2 API Bridge Agent
To connect LLM-native intent with downstream MCP or OpenAPI-compatible services, API Bridge Agent [7], built atop the Tyk gateway [48], provides translation, routing, and orchestration. It converts natural language prompts into structured API calls, resolving endpoints through semantic matching, policy validation, and tool availability checks. The middleware supports multiple invocation modes. In Direct Mode, the agent specifies both the service and exact API endpoint, enabling precise control. In Indirect Mode, the agent selects the service, while the middleware identifies the best endpoint to fulfill the task intent. In Cross-API Mode, the agent supplies only the intent, and the middleware determines both the service and endpoint across multiple APIs. In MCP Proxy Mode, the middleware coordinates dynamic tool invocation and context enrichment via standardized MCP tool descriptions. This unified interface allows agents to flexibly access diverse services with minimal coupling.
VII-A3 Function Calling Mechanisms
At the invocation level, agents rely on standardized formats to express, trigger, and handle tool execution. Among the most widely adopted approaches are:
OpenAI Function Calling. This method [213] allows developers to expose custom logic to the model via JSON schemas describing function name, description, and argument structure. When a model determines that a function should be invoked, it emits a well-formed JSON object representing the function call. The agent runtime interprets this object and routes control to the corresponding tool. While lightweight, extensible, and easy to implement, this approach is generally limited to basic argument serialization patterns and single-step invocations.
LangChain Tool Calling. LangChain [154] enhances the function calling paradigm through a richer abstraction layer. Tools are defined via a standardized schema, including argument types, input-output post-processing, and plugin registration. Tools are accessible through a runtime registry that supports nested calls, conditionals, and fallback strategies. This mechanism is particularly suited for agent frameworks supporting dynamic routing and chained tool reasoning.
VII-A4 Tool Metadata Declaration: Agents.json
To ensure tool visibility and adaptive behavior across agents, agents.json [303] serves as a standardized metadata format for interface declaration. Built on OpenAPI foundations but customized for agent consumption, it enables developers to define authenticated entry points, input-output types, and multi-step orchestration plans such as:
Flows: Predefined composition of tool steps for common actions.
Links: Declarative dependency mappings between parameter bindings.
Agents.json bridges the configuration plane between runtime reasoning and API surface documentation. It ensures that agents can discover tools introspectively and plan actions without manual reconfiguration or hardcoded logic.
| Stage | Risk Type | Representative Threats | Attack Characteristics |
| Malicious environments against benign agents | Memory-related Risks | Memory Injection | Induce agents to generate and record harmful content through natural interactions |
| Memory Poisoning | Implant pairs of examples containing adversarial triggers and payloads | ||
| Memory Extraction | Extract sensitive data in memory modules without explicit queries for privacy | ||
| Knowledge-related Risks | Knowledge Corruption | Inject adversarial texts into external knowledge bases to influence the response | |
| Privacy Risks | Craft prompts to induce agents to recover sensitive or private content | ||
| Tool-related Risks | Malicious Tools | Malicious tool functions or prompts embedded in the tool description | |
| Manipulation of Tool Selection | Inject misleading prompts to make agents select the compromised tool | ||
| Cross-Tool Chaining Exploits | Manipulate the output of a tool to inject malicious inputs for subsequent tools | ||
| Compromised agents against benign environments | Information Corruption | Propagate Tampered Information | Undermine an agent’s memory and knowledge using misleading information |
| Misusing Tools | Data Exfiltration | Extract and transfer confidential information out of protected environments | |
| System and Service Disruption | Disrupt the functionality of services without directly targeting data confidentiality | ||
| Malicious Content Propagation | Spread malicious content to trusted storage resources | ||
| Real-World Damage | Digital Environment Contamination | Introduce harmful elements into digital ecosystems, thereby corrupting integrity | |
| Physical Environment Disruption | Compromised agents cause disruptions to real-world physical environments |
VII-B Security Risk Analysis: Malicious Environments Against Benign Agents
As the capabilities of LLM-powered agents continue to evolve, their interactions with the external world become increasingly complex and powerful. In particular, the integration of memory systems and external tool invocation introduces a new set of attack surfaces that attackers can exploit. This section provides an in-depth analysis of the security risks that arise specifically from these two modules: the memory module, responsible for storing and retrieving contextual information, and the tool module, which enables agents to execute actions by interfacing with external systems or services (e.g., via function calls). We first explain how these two components typically function in the agent ecosystem and outline the general attack paradigms against each. We then provide detailed analyses of specific vulnerabilities, attack techniques, and representative works from the security literature that highlight these threats.
VII-B1 Memory-related Risks
Memory modules play a crucial role in enabling agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [351]. Unlike stateless language models that depend solely on immediate prompts, memory-equipped agents maintain long-term historical information through external storage systems, such as vector databases or document repositories. These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [351].
Typically, a memory module operates through three stages: write, retrieve, and apply. During the write phase, the agent logs past utterances, tool outputs, subgoals, or retrieved facts into memory. Later interactions initiate the retrieval phase, where semantically similar records are fetched via embedding matching or keyword search. These records are then injected into the model’s context window or used for downstream decisions, forming the apply phase. While this architecture empowers agents with dynamic reasoning abilities, it also introduces new vulnerabilities that extend beyond the conventional LLM prompt space.
Recent research has unveiled multiple categories of memory-related attacks, such as memory injection, memory poisoning, and memory extraction. These adversarial methods exploit the openness, autonomy, or persistent nature of the memory module to manipulate agent behavior or extract sensitive data. We now describe each threat in detail.
Memory Injection: In memory injection attacks, attackers insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access. The attack leverages the agent’s autonomous memory-writing mechanism by inducing it to generate and record harmful content. Once stored, these entries can be retrieved by benign user queries due to embedding similarity, thus indirectly triggering undesired behavior such as altered reasoning or unsafe tool invocations. A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [61, 307]. These steps, once embedded in memory, become semantically linked to a targeted victim query. When the victim issues a benign instruction, the poisoned memory is likely to be retrieved, thereby hijacking the agent’s planning process. This strategy requires no direct injection channels beyond normal user interaction, yet demonstrates high attack success and stealthiness across multiple agent environments.
Memory Poisoning: Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads. These attacks are typically conducted by polluting a subset of the memory with trigger-output pairs that only activate when specific inputs are encountered. During the retrieval phase, if the user’s query resembles the trigger, the agent is likely to load the poisoned entries and be influenced toward compromised outputs. Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [38]. This method generalizes across agent types and does not require model access or parameter modification.
Memory Extraction: In addition to injection and poisoning, memory modules pose risks of unintended information leakage. Since LLM agents often log detailed user-agent interactions-including private file paths, authentication tokens, or sensitive instructions-malicious queries may be used to extract such data. This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [332, 128]. It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [278]. Memory extraction can occur even without explicit queries for private content, instead relying on semantic proximity in the vector space to surface related sensitive traces. These findings highlight not only the retrieval vulnerability but also the insufficiency of downstream response filtering as a defense.
VII-B2 Knowledge-related Risks
External knowledge techniques, such as Retrieval-Augmented Generation (RAG), combine the generative strength of LLMs with the factual accuracy and relevance of external knowledge retrieval systems. Instead of relying solely on parametric knowledge stored within the pretrained model, RAG augments generation by sourcing passages from an external knowledge base in response to the input query. These retrieved documents are then concatenated with the query and passed into the LLM for final answer generation. This paradigm enables more informed, up-to-date, and domain-specific language understanding, and it is widely adopted across applications such as open-domain question answering, customer service agents, recommender systems, and multi-step planning agents.
Despite its performance advantages, the RAG architecture introduces new security risks that are distinct from those inherent to pure neural models [32, 39, 316, 365]. In particular, the information retrieval module-serving as the agent’s external memory-becomes an adversarial surface where unverified or manipulable corpora may be exploited. Attacks targeting these corpora can bias the retrieval process, manipulate generation outcomes, or expose previously unseen private data.
Knowledge Corruption via Data Poisoning: A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries. These poisoned passages are semantically aligned to specific triggers but contain harmful, misleading, or attacker-intended content. Once injected into the knowledge base, they can be prioritized during retrieval and directly influence the LLM’s final response. Several recent works have demonstrated the feasibility of such attacks. PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [365]. Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [209]. Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [358].
Privacy Risks and Unintended Leakage: RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs. This retrieval behavior implicitly enables information leakage when attackers craft prompts that induce the model to recover sensitive or private content from the corpus. The risk is amplified when access permissions on the corpus are loosely controlled or aligned purely through similarity metrics. Recent studies have called attention to this concern. Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [332]. These attacks demonstrate that simply adding a retrieval layer does not automatically mitigate the privacy vulnerabilities of LLMs-in fact, it may exacerbate them if not complemented with access control, context filtering, or signal sanitization.
Compared to memory modules, RAG corpora are often larger, dynamically updatable, and more difficult to monitor. Because retrieval corpora may be sourced from web documents, community-shared datasets, or user uploads, attackers can often poison them without interacting directly with the agent. Moreover, dense retrieval introduces additional attack vectors via embedding collisions or adversarial representation alignment, wherein malicious documents are optimized to collide with benign queries in the retriever’s latent space.
VII-B3 Tool-related Risks
Tools are essential to the functionality of LLM agents, extending the model’s capabilities to perform structured actions, access external data, invoke system functions, or interact with digital environments. Agent architectures typically support tool integration through two primary paradigms: native function calling APIs (e.g., OpenAI-style schema-based calls) and protocol-based interfaces such as the MCP, which unify tool metadata, invocation templates, and language model binding.
Despite differences in instantiation, both paradigms share a common interaction lifecycle: (1) tool description ingestion, (2) tool selection and planning, (3) input argument generation, (4) tool invocation, and (5) output parsing or chaining. This structured pipeline forms the agent’s “action surface,” introducing multiple security-critical operations vulnerable to adversarial manipulation. We now review a range of known or emerging attacks targeting different stages of the tool interaction process.
Malicious Tools as Attack Vectors: Given that many tools are authored externally or retrieved from shared tool repositories, attackers may publish seemingly benign tools containing covert malicious logic. For example, researchers reveal that MCP enables attackers to embed hidden prompts or malicious instructions in not only executable functions [111] but also tool metadata fields such as descriptions, example usages, or API annotations[254, 72, 101, 150, 229]. These embedded messages can influence the LLM’s planning behavior, bypass output constraints, execute malicious code, leak privacy, and redirect queries.
Manipulation of Tool Selection: Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation. This selection logic can be hijacked. Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options.
Research indicates that attackers can generate synthetic tool descriptions that stealthily override the model’s planning process [250]. These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations. Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during the planning phase. Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
Cross-Tool Chaining Exploits: As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls. These workflows blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations. Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools). In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server”, which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [44].
VII-C Security Risk Analysis: Compromised Agents Against Benign Environment
In Section VII-B, we analyzed how the malicious environments can influence an agent through its memory and tool modules. However, once an agent is compromised, it can cause widespread and lasting damage to external environments. This section explores this reverse threat, focusing on how compromised agents can exploit their functionalities to contaminate memory modules, abuse external tools, and pose risks to the real world.
VII-C1 Corrupting Memory and Knowledge
Memory and knowledge bases are central to how the agent thinks and acts. Once these components are corrupted, they can lead to incorrect reasoning and decisions. Moreover, the damage can infect other agents and cause long-term harm to the entire system.
A compromised agent can become a source of systemic contamination. Through agent communication, it can actively propagate tampered knowledge and flawed reasoning patterns, spreading its internal corruption to other agents and triggering a cascading infection of memory modules and knowledge bases across the system [133, 38, 158]. Once the shared knowledge base is compromised, other agents may unknowingly retrieve and integrate malicious information into their memory module during task execution, converting knowledge base contamination into memory infection. Subsequently, an agent with a compromised memory module can use its authorized write access to contaminate the shared knowledge base of the entire system, in turn, forming a reverse contamination loop from memory to knowledge. Since these contamination operations come from the trusted agents within the system, they are highly difficult to detect. Once established, the contamination can persist long-term, continuously disrupting the behavior of compromised agents and misleading other agents that rely on the same knowledge sources, resulting in a form of chronic poisoning of the information ecosystem.
VII-C2 Misusing Tools
Tools serve as the channels through which an agent interacts with the world. Once the agent is compromised, tools can be exploited to cause harm. Since agents are often granted permissions to execute specific API calls or system commands, misuse of these tools can lead to severe consequences.
Data Exfiltration: If a compromised agent has tool permissions to access databases, file systems, or communication interfaces, it can be manipulated to extract and transmit confidential data, intellectual property, or personally identifiable information. These payloads can be covertly delivered to external servers via email, HTTP requests, or chained tool invocations. For instance, a compromised agent equipped with Markdown rendering capabilities might unknowingly embed malicious image links in generated content. When rendered by a browser, these links trigger covert HTTP requests that exfiltrate sensitive data such as email addresses or access tokens to attacker-controlled servers. Notably, attackers may also exploit tool misuse to gain indirect access to backend systems and gradually escalate privileges through cross-toolchain interactions.
System and Service Disruption: A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to carry out destructive operations. A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to execute destructive operations. Attackers may exploit these privileges to launch internal network scans, initiate denial-of-service attacks, or issue commands, potentially causing system crashes or irreversible business data loss. Such actions are especially dangerous when triggered via tool interfaces that lack proper input validation. Vulnerabilities commonly exploited in these contexts include parameter injection, server-side request forgery (SSRF) [22], and arbitrary file access. Beyond technical damage, the compromised agent can be manipulated to disrupt business operations, such as erroneously invoking the “cancel all orders” tool. Moreover, adversaries may use retrieval tools to fetch misleading information, then exploit decision-making tools such as trading agents to automatically execute harmful operations, resulting in major financial losses.
Malicious Content Propagation: Agents with permissions to publish externally (e.g., via email, social media, or CMS APIs) can be exploited to widely spread malware, phishing links, or misinformation once compromised. For example, a trusted customer service agent may send malware-laden emails to clients, or a content generation agent might publish misleading articles on official websites. Because these agents are trusted entities, such attacks are highly deceptive. More dangerously, attackers can leverage access to contacts, email histories, and user preferences to craft highly personalized phishing campaigns, enabling large-scale social engineering attacks.
VII-C3 Real-World Damage
Compromised agents can cause harm far beyond their internal systems, affecting both digital and physical ecosystems through their interactions with the external world. This section examines how such agents may disrupt the real world by polluting shared digital environments and executing harmful physical actions.
Digital Environment Contamination: A compromised agent can cause long-term damage to the external digital environment, not by directly attacking other agents, but by polluting the shared information ecosystem they rely on. Since agents often interact with external platforms (such as submitting code on GitHub or editing Wikipedia entries), once compromised, they can systematically inject subtle yet harmful errors or biases into these shared resources [173]. Unlike cross-agent contamination, digital environment contamination indirectly infects all agents dependent on the polluted information sources. For example, a compromised coding agent may embed hidden logic bugs or backdoors when contributing code; a corrupted knowledge management agent might alter Wikipedia pages or internal knowledge bases by falsifying citations or inserting biased descriptions, thereby damaging the entire knowledge graph with far-reaching consequences.
Physical Environment Disruption: Once an agent’s memory or tool module is compromised, the resulting threat is not limited to digital risk. It can manifest as concrete damage to the physical world through specific decision-making chains and execution paths. Polluted memory may contain falsified sensor data, misleading the agent’s perception of the physical environment. Tool modules that serve as interfaces to physical systems can directly implement flawed decisions, affecting device behaviors, environmental control, or industrial processes. For example, an agricultural agent misled by false pest-related memories may overapply pesticides; a quality inspection robot referencing corrupted standard images may repeatedly approve defective components; a warehouse robot using a compromised path-planning module may unknowingly cause stacking imbalances and logistics bottlenecks. Critically, such behaviors often appear to follow normal procedures, making them difficult to detect through traditional logging or anomaly detection methods. Thus, the consequences of agent compromise extend beyond digital misinformation, posing a tangible risk of physical system disruption.
VII-D Defense Countermeasure Prospect
The growing complexity and autonomy of LLM-based agents demand equally sophisticated security strategies. As these systems increasingly rely on memory modules, retrieval augmentation, and interactive toolchains, the corresponding attack surfaces have expanded across diverse layers-including context propagation, planning logic, and execution flows. Addressing these vulnerabilities requires a multi-layered, compositional defense framework. This section reviews current and emerging countermeasures along three critical dimensions: memory-based attacks, RAG vulnerabilities, and tool-centric threats.
VII-D1 Countermeasures For Memory- and knowledge-related Risks
Memory- and knowledge-related risks can be jointly addressed through an integrated mitigation framework spanning content filtering, output consensus, and architectural isolation.
Embedding-Space Screening and Clustering-Based Anomaly Detection: Whether memory entries are agent-internal or retrieved externally via RAG, their semantic embeddings can be preemptively analyzed for anomalies. Techniques like TrustRAG [359] apply clustering (e.g., K-means) to identify vectors that deviate from the dominant semantic cluster. This approach effectively filters both static memory entries and retrieval results with low semantic cohesion, regardless of source. While lightweight and interpretable, clustering-based filtering must be augmented with adaptive schemas to detect context-sensitive triggers or stealthy distributional shifts.
Consensus Filtering and Voting-Based Aggregation: To limit the model’s reliance on single compromised retrievals or poisoned memories, output-level consensus mechanisms have been proposed. RobustRAG [306], for example, treats each retrieved source independently and constructs responses based on overlapping semantic content (e.g., shared n-grams or keywords) across documents. This same principle can be extended to memory snapshots through majority-vote or semantic voting strategies, where only widely corroborated memories can influence the response. Such ensemble-style filters improve resilience by diluting the influence of outlier or adversarial sources.
Execution Monitoring and Planning-Consistency Checks: Adversarial content within memory or RAG inputs may subtly deviate the agent’s behavior from user intent without explicit toxicity. Tools like ReAgent [30] introduce planning-level introspection where the agent paraphrases the user’s request, generates an expected plan, and continuously aligns runtime actions with this trace. Any inconsistency, triggered by an unexpected memory or an off-topic retrieval, is treated as a behavioral anomaly and can prompt halting or recovery mechanisms. This introspective framework provides a robust guardrail against both memory-hijacking and injection-aware RAG attacks.
System-Gated Memory Retention and Input Sanitization: Architectural solutions such as DRIFT [99] and AgentSafe [196] implement strict content sanitization before newly generated content. DRIFT uses an injection isolator to scan generative outputs for adversarial goal shifts or impersonation cues, while AgentSafe enforces trust-tiered storage via ThreatSieve and prioritization via HierarCache. These mechanisms constrain future influence, ensuring that RAG or memory poisoning cannot silently accumulate over time.
Unified Content Provenance and Trust Frameworks: Since retrieved knowledge and persisted memories may originate from overlapping sources (e.g., user prompts, tool calls, external APIs), maintaining clear provenance metadata and trust scores is essential. Unified provenance tracking across both memory and retrieval pipelines enables smarter decisions about retention, ranking, or discounting of contentious content. Combined with per-source reliability scoring, this approach encourages transparent auditing and facilitates downstream fine-tuning or gating mechanisms.
In summary, memory- and knowledge-related threats reflect different modalities of persistent and dynamic context manipulation, but share overlapping vectors of attack and can benefit from synergized defenses. Embedding-level screening filters anomalous content at ingestion, consensus aggregation constrains influence at generation, and architectural isolation confines latent impact across sessions. Moving forward, defense designs should increasingly treat RAG and memory as compositional context modules secured and governed under a shared set of verification, introspection, and isolation principles.
VII-D2 Countermeasures For Tool-related Risks
Tool-related defense strategies should operate across four interlocking levels: protocol foundations, execution control, orchestration safety, and system enforcement.
Protocol-Level Safeguards: To counter risks such as tool poisoning, cross-origin exploits, and shadowing attacks enabled by flexible yet insufficiently regulated protocols like MCP, researchers have introduced security-verification frameworks operating at the registry and middleware layer. MCP-Scan [152] performs both static inspection of tool schemas (e.g., scanning for suspect tags or metadata) and real-time proxy-based validation of MCP traffic, leveraging LLM-assisted heuristics to flag covert behaviors. MCP-Shield [236] extends this with signature-matching and adversarial behavior profiling, enabling pre-execution detection of high-risk tools and malformed tasks. MCIP [130] builds on MAESTRO [49] to analyze runtime traces, proposing an explainable logging schema and a security-awareness model to track violations in complex agent-tool interactions.
Tool Invocation and Execution Controls: At the agent’s runtime execution point, classic techniques such as sandboxing and permission gating remain foundational. Google’s defense-in-depth model advocates policy engines that monitor planned tool actions, verify argument safety, and require human confirmation for risk-sensitive operations [65]. Tools should be executed in minimally privileged environments-e.g., isolated containers with controlled filesystem and network scope-to mitigate direct misuse, including SSRF and data exfiltration threats. Enforcement frameworks can also implement schema hardening or fine-grained input/output sanitization to reject anomalous payloads.
Agent-Orchestration Monitoring: Newer approaches target the agent’s planning cognition-its selection and chaining of tools. GuardAgent [308] introduces a validator agent that inspects the primary agent’s plan and generates executable guards (e.g., static checks or runtime assertions) before tool calls proceed. AgentGuard [34] takes a more declarative view: it uses an auxiliary LLM to model preconditions, postconditions, and transition constraints across multi-step tool workflows, effectively constraining the planner rather than reacting after execution begins. These strategies reflect a growing consensus: LLMs may require another LLM to safely oversee complex planning under uncertainty.
System-Level Mediation and Chaining Control: Complex pipelines-such as summarize(search(“…”))-can become attack vectors when tools trust upstream outputs implicitly. To prevent this, DRIFT [99] introduces a structured control architecture: a “Secure Planner” compiles a validated tool trajectory under strict parametric constraints, while a “Dynamic Validator” continuously monitors downstream tool executions for compliance. Notably, the Injection Isolator blocks adversarial propagation between tools by sanitizing both intermediate returns and final outputs, mitigating the risk of memory poisoning and delayed-stage tool exploits.
VII-E Takeaways
Agent-environment communication protocols like MCP enable agents to interface with diverse tools, APIs, and external data. However, they introduce risks such as memory injection, retrieval-augmented generation poisoning, and tool misuse. Malicious attackers can corrupt memory stores, manipulate knowledge bases, or exploit cross-tool chain vulnerabilities to harm the agent system. Besides, compromised agents can also cause damage to benign environments. To help developers mitigate these problems, we discuss the targeted defense countermeasures for risks from malicious environments. The defenses for compromised agents can be referred to Section VI-C. However, we also believe that related attacks will continue to emerge, and there needs long-term efforts to make agent-environment communication more secure.
VIII Experimental Case Study: MCP and A2A
To help readers better understand the new attack surfaces that agent communication brings, we select typical protocols and conduct attacks against them.
VIII-A The Selection of Protocols
Since there are many protocols related to agent communication, it is impossible to exhaustively evaluate all of them. As a result, we decided to deploy typical examples and conduct experiments. We mainly focus on the following two principles.
Popularity. The protocol selection needs to give priority to practicality. By evaluating protocols with higher popularity, it can be guaranteed that the revealed vulnerabilities have more significant practical value.
Maturity. The selected protocol should have the highest possible maturity. For instance, it should provide relatively complete open-source projects and test cases, and there are also many applications or open-source projects based on them.
Based on the above principles, we finally chose MCP and A2A. MCP is almost the most popular agent communication protocol currently. It has been adopted by diverse companies and individual developers. Up to now, MCP’s PyPI package has received over 9 million downloads in the last month [20], and its NPM download last week reached 4.2 million [21]. Besides, the number of open-source MCP servers also approached 2,000 [101]. After investigation, we find that none of the other existing protocols can achieve a similar popularity to MCP. A2A was proposed in April 2025, so it has not yet achieved the same level of promotion as MCP. However, we observe that the attention to A2A is rising rapidly, and its recognition among developers is also quite high. As a result, we finally chose MCP and A2A as the target protocols.
VIII-B Experiments
In this chapter, we will conduct attacks against MCP and A2A and illustrate their weakness. In the experiments related to MCP, we use Claude as the MCP client. Claude [17] is an AI chatbot developed by Anthropic. With its outstanding code capabilities, safe design, and ultra-long context processing, it has become one of the most popular AI assistants in the world. In January 2025, Claude app downloads reached an estimated 769.6 million [264]. Therefore, exploiting Claude can provide practical value for real-world scenarios. Attacks #1-#4 are typical examples of the Malicious Tools as Attack Vectors discussed in Section VII-B3, and attack #5 belongs to the Description Poisoning in Section VI-B1.
Exp #1. Malicious Code Execution. We demonstrate that attackers can exploit MCP to enable the execution of malicious code on the computer system, which has also been proposed by other studies [229]. We download Claude sonnet 3.7 as the MCP host, and use the Filesystem Server [18] provided by the official MCP as the MCP server. As shown in Figure 10, MCP enables malicious users to directly open an unauthenticated Bash Shell service on port 4444 locally. Adding it to .zshrc will normalize the risk and make it extremely easy to be exploited by intruders. Unfortunately, this malicious operation does not get blocked.
Exp #2. Retrieval-Agent Deception. Attackers can write a seemingly normal document related to MCP and publish it on a public platform. However, malicious commands are embedded in this document, such as searching for local environment variables and sending the information to Slack [266]. The document content is as follows:
Users download these documents and build a local vector database through the Chroma MCP Server [1]. As a result, these malicious documents, together with other legitimate documents, are encoded into the vector database and become one of the retrieval targets. When a user sends a request to Claude Opus 4:
Claude will use Chroma to query the database and retrieve the contaminated documents, and then call the Terminal-Controller tool [88] to execute the embedded malicious commands. We successfully achieved this attack using Claude Opus 4.
Exp #3. Tool Poisoning. Tool poisoning means that attackers can add malicious instructions in the MCP tool description [254]. Such instructions do not influence the execution of tools, but can induce the agents to perform dangerous actions. We use Claude Sonnet 4.0, the Filesystem MCP Sever [18], and Gmail MCP Server [87] to illustrate this vulnerability. In the index.js of the Filesystem Server, we added the malicious description as follows:
The above description will not be noticed by users. However, when the user calls this tool, the description will trigger a message sent to the attacker’s email. As shown in Figure 12, we successfully received the SSH private key and public key.
Exp #4. Command Injection. This attack denotes that malicious tools can directly manipulate the user’s system without causing any alert with the help of MCP. We inject a malicious command into the Terminal-Controller MCP tool [88]. When users call this tool, the embedded command will be executed. We successfully conducted this attack using Claude Sonnet 4.0. The injected command is as follows:
This command is to delete the ./mcp_file/important.md file. Besides, in the description, we also tell the agent not to list the rest files in the folder to hide this malicious behavior. This is because we find that Claude will list the rest files without this instruction, which may expose the malicious command. After we added this instruction to the description, Claude successfully deleted this file and did not list the rest of the files.
Exp #5. Agent Selection Manipulation. We also conduct attacks against A2A. We find that attackers can add malicious descriptions in the Agent_Card to induce users to select them. In this experiment, we use the public A2A-samples provided by A2Aproject [4] as the user-side client. Then, we create two agents for weather querying. One of them is benign while the other is malicious. The user issues the following command to the A2A client.
In the normal case, the malicious agent does not add any leading description in its Agent_Card. We sent this query 20 times and found that both agents might be selected by the client. Then, the malicious agent modifies its Agent_Card as follows:
We can see that the attacker emphasizes that the malicious agent can provide the richest features, and adds an instruction that the client should select it first. Under this condition, as shown in Figure 13, we found that the selected agent by the A2A client was always the malicious agent. The result demonstrates that attackers can improve their priority to handle the users’ tasks by simply adding some descriptions.
IX Future Directions Discussion
IX-A Technical Aspects
IX-A1 Powerful but Lightweight Malicious Input Filter
We deem that user inputs are still the largest-scale attack carrier in the agent ecosystem, especially considering that the inputs are becoming more open (no longer limited to user instructions but also contain environment feedback), multimodal, and semantically complex. Besides, the future agent ecosystem will pay more attention to effectiveness, especially given that the running speed of LLMs is inherently slow. Such dual demand will put a very heavy burden on related defenses. As a result, to mitigate this problem, lightweight but powerful malicious input filters must be established. This not only requires mature techniques in AI to slim the defense models down (just like DeepSeek), but also needs to integrate with other techniques, such as offloading some fundamental computing on the programmable line-speed devices (e.g., programmable switches and SmartNICs) to facilitate the input filtering process.
IX-A2 Decentralized Communication Archiving
It is important to record the communication process and contents for some specific fields, such as finance. This is to audit potential crimes and mistakes once agents cause problems that cannot be ignored. Given security and reliability, such storage cannot rely on a single storage point and must guarantee integrity and efficiency. To this end, other techniques such as blockchain should be adopted to manage the historical communication. It is easier for CS-based communication because there exist centralized servers for establishing a locally distributed archiving mechanism, such as a distributed storage chain in enterprise networks. In contrast, how to achieve decentralized communication archiving for P2P-based communication, especially for cross-country agents, is almost a construction that needs to start from scratch.
IX-A3 Real-time Communication Supervision
Although post-audit is indispensable, real-time supervision can minimize damage once attacks or mistakes occur because it has a shorter reaction time. We believe CS-based communication faces less difficulty in building such supervision mechanisms. This is because centralized architectures have natural advantages in monitoring the entire network. In contrast, P2P-based communication may require more effort to enable collective supervision. We think it is an important function to build a reliable and secure AI ecosystem.
IX-A4 Cross-Protocol Defense Architecture
Although existing protocols solved the problem of heterogeneity to some extent, different protocols also lack seamless collaboration. For example, it is still difficult to assign a universal identity for agents and tools (cross A2A and MCP), which degrades the system performance and may incur inconsistency errors if not orchestrated correctly. Future AI ecosystems should focus on a more universal architecture to integrate different protocols and agents together, like IPv4, thereby enabling seamless discovery and communication among different agents and environments.
IX-A5 Judgment and Accountability Mechanism for Agent
It is still difficult to locate and assign responsibility for the behavior of agents. For example, in a failed task execution process, it is hard to identify which steps caused the final deviation of the result, no matter they are malicious or unintentional. This is because a tiny deviation in the middle process may lead to a final gap between benign and dangerous results. Besides, it also needs a principle to quantify the responsibility for each agent or action. We believe this aspect will significantly address the urgent need of the current AI ecosystem.
IX-A6 Trade-offs between Efficiency and Accuracy
Agent communication is fundamentally a process of information transmission, and thus can be analyzed through the lens of information theory. In this aspect, we think there are two types of directions.
High-token Communication: A larger number of tokens allows agents to convey richer contextual semantics, more detailed instructions, and more complex logic, thereby reducing ambiguity and enhancing the accuracy of multi-agent coordination. In tasks that require fine-grained understanding, verbose natural language descriptions help align goals among agents and reduce deviations. However, excessive tokens significantly increase costs and processing time, resulting in lower system efficiency and higher latency. Moreover, longer contexts expose larger attack surfaces for prompt injection and data poisoning, enabling adversaries to hide malicious content more covertly. Additionally, information overload may distract agents, causing them to infer incorrect information from irrelevant context and increasing the likelihood of hallucinations.
Low-token Communication: Using concise and structured messages (e.g., JSON formats) greatly improves communication efficiency. This approach reduces computational costs, increases transmission speed, and simplifies parsing, thereby minimizing potential errors. However, low-token communication lacks the flexibility to express complex intentions or respond to unforeseen scenarios. If the predefined protocol or format fails to capture the full semantic intent, it can lead to significant information loss and failed collaboration.
The design of future agent communication protocols needs to involve a trade-off between efficiency and accuracy. Future research should explore adaptive communication protocols that dynamically adjust the degree of redundancy and structure based on task complexity, security requirements, and agent capabilities. For example, high-token communication may be used during the exploration phase of a task, while low-token communication can be adopted during execution to ensure efficiency and safety.
IX-A7 Towards Self-Organizing Agentic Networks
With the increasing scale of IoA, in the future, agent communication is expected to evolve toward self-organizing agentic networks, where agents autonomously discover each other, assess capabilities, negotiate collaborations, form dynamic task groups, and disband upon completion. This paradigm offers high scalability and robustness, making it well-suited for dynamic and unpredictable environments.
IX-B Law and Regulation Aspect
Apart from the technical aspect, we find that there are still serious deficiencies in the laws and regulations related to agents. Such blanks cannot be remedied by techniques. We call for accelerating the improvement of laws and regulations in the following aspects.
IX-B1 Clarify the Responsible Subject
When a sold agent causes property damage or personal injury to others, it is difficult to determine the ultimate responsible subject. For example, if an intelligent robot damages the property during the execution of a task, the law-level quantification of the responsibility of the developers, users, or enterprises lacks a clear definition. In addition, for problems arising from the collaborative work of multiple agents, such as an accident occurring when multiple autonomous driving vehicles are traveling in formation, there is a lack of legal provisions regarding the division of responsibilities among the enterprises to which the vehicles belong or the relevant subjects.
IX-B2 Protect Intellectual Property Rights
Nowadays, there has been a large number of LLMs that have been open-sourced. These can act as the brain of different agents. However, even for open-source LLMs, the publishers still restrict their application scope, e.g., other developers should also open-source their agents built on these LLMs. However, there still lacks laws to effectively protect such intellectual property. For example, the criteria for determining plagiarism in agents is not clear. Even if plagiarism is determined, there is still a lack of defining standards for the degree of plagiarism (e.g., 50% or 90%?). We think there urgently need related laws and regulations.
IX-B3 Cross-border Supervision
Agent communication has a transnational nature. An agent trained in one country may be used for illegal activities by people from other countries. At this time, it is difficult to determine which country’s laws apply, and there is a lack of unified international supervision standards and judicial cooperation mechanisms, which may easily lead to difficulties in cross-border security.
To our knowledge, the related formulation of laws and regulations (such as those related to agent crimes) lags far behind the development of agents. For example, how to define the theft and misappropriation of agents, the accident responsibility of autonomous driving agents
| Abbr. | Full Form |
| A2A | Ageng-to-Agent Protocol |
| ACN | Agent Communication Network |
| ACP-AGNTCY | Agent Connect Protocol by AGNTCY |
| ACP-IBM | Agent Communication Protocol by IBM |
| ACP-AgentUnion | Agent Communication Protocol by AgentUnion |
| AI | artificial intelligence |
| AITP | Agent Interaction & Transaction Protocol |
| ANP | Agent Network Protocol |
| API | Application Programming Interface |
| CNN | Convolutional Neural Network |
| CoT | Chain of Thought |
| DID | Decentralized Identifier |
| DNS | Domain Name System |
| DoS | Denial of Service |
| FNN | Feedforward Neural Network |
| GAN | Generative Adversarial Network |
| GNN | Graph Neural Network |
| IoA | Internet of Agents |
| LMOS | Language Model Operating System Protocol |
| LLM | Large Language Model |
| LOKA | Layered Orchestration for Knowledgeful Agents |
| LSTM | Long Short-Term Memory |
| MAS | Multi-Agent Systems |
| MCP | Model Context Protocol |
| MITM | Man-in-the-Middle |
| OAuth2 | Open Authorization 2.0 |
| PXP | PXP protocol |
| RAG | Retrieval-Augmented Generation |
| RNN | Recurrent Neural Network |
| SSL | Secure Sockets Layer |
| SPPs | Spatial Population Protocols |
| TLS | Transport Layer Security |
X Conclusion
This survey systematically reviews the security issues of agent communication. We first highlight the differences between previous related surveys and this survey, and summarize the preliminaries of LLM-driven agents. Then, we make a definition and classification of agent communication to help future researchers quickly classify and evaluate their work. Next, we detailedly illustrate the communication protocols, security risks, and possible defense countermeasures for three agent communication stages, respectively. Then, we conduct experiments using MCP and A2A to help illustrate the new attack surfaces brought by agent communication. Finally, we discuss the open issues and future directions from technical and legal aspects, respectively.

### v4

@@
 Abstract
-In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence and adaptability, and are rapidly changing human production and life.
+In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence and adaptability.
 Nowadays, agents are undergoing a new round of evolution.
 They no longer act as an isolated island like LLMs.
-Instead, they start to communicate with diverse external entities, such as other agents and tools, to perform more complex tasks collectively.
-Under this trend, agent communication is regarded as a foundational pillar of the future AI ecosystem, and many organizations have intensively begun to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the recent few months.
+Instead, they start to communicate with diverse external entities, such as other agents and tools, to perform complex tasks.
+Under this trend, agent communication is regarded as a foundational pillar of the next communication era, and many organizations have intensively begun to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the past year.
 However, this new field exposes significant security hazards, which can cause severe damage to real-world scenarios.
 To help researchers quickly figure out this promising topic and benefit the future agent communication development, this paper presents a comprehensive survey of agent communication security.
-More precisely, we first present a clear definition of agent communication and categorize the entire lifecycle of agent communication into three stages: user-agent interaction, agent-agent communication, and agent-environment communication.
-Next, for each communication phase, we dissect related protocols and analyze the security risks according to the communication characteristics.
-Then, we summarize and outlook on the possible defense countermeasures for each risk.
+More precisely, we present the first clear definition of agent communication.
+Besides, we propose a framework that categorizes agent communication into three classes and uses a three-layered communication architecture to illustrate how each class works.
+Next, for each communication class, we dissect related communication protocols and analyze the security risks, illustrating which communication layer the risks arise from.
+Then, we provide an outlook on the possible defense countermeasures for each risk.
 In addition, we conduct experiments using MCP and A2A to help readers better understand the novel vulnerabilities brought by agent communication.
 Finally, we discuss open issues and future directions in this promising research field.
-Index Terms:
-large language model, agent communication, attack, and securityI Introduction
-The emergence of Large Language Models (LLMs) has led to revolutionary advancements in Artificial Intelligence (AI), exhibiting unprecedented capabilities in understanding complex tasks [327].
-More importantly, LLMs greatly boosted the ideal form of AI that human expects: agents111In this paper, all agents refer to LLM-driven AI agents..
-Different from LLMs that are mainly like chatbots, agents possess more comprehensive capabilities (e.g., perception, interaction, reasoning, and execution), enabling them to independently complete a real-world task.
-For example, when users seek to make a travel plan, LLMs can only provide the best plan in text, while agents can realize the best plan in action, such as checking the weather, buying tickets, and booking hotels.
-Agents greatly speed up the progress of the intelligence transformations of enterprises.
-Their market size is expected to increase by 46% per year [235].
-It can be foreseen that agents will subvert the production and living patterns of modern society, greatly changing the future business landscape.
-As a result, developing and promoting agents have become a strategic planning of major powers and influential companies.
-Now, agents are evolving towards the direction of domain-specific entities, i.e., being customized for specific scenarios and tasks.
-In this context, as shown in Figure 1, a task usually requires the collaboration of multiple agents, which may be located globally on the Internet.
-Under this condition, agent communication becomes the foundation of the future AI ecosystem.
-It enables agents to find other agents with specific capabilities, access external knowledge, assign tasks, and engage in other interactions.
-Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to the development of agent communication.
-In November 2024, Anthropic proposed Model Context Protocol (MCP) [19], a universal protocol that allows agents to invoke external environments, such as datasets, tools, and APIs.
+We also publish a repository that maintains a list of related papers on https://github.com/theshi-1128/awesome-agent-communication-security.
+I Introduction
+Computer communication has been evolving in the direction of improving interaction efficiency.
+Now, the emergence of Large Language Model-driven AI agents111In this paper, all agents refer to LLM-driven AI agents.
+[yu2025table, naveed2025comprehensive, zhao2023survey] brings revolutionary changes to this field.
+First, LLMs greatly boosted the evolution of agents, new entities that integrate perception, interaction, reasoning, and execution capabilities to automatically complete a real-world task.
+For example, when users seek to make a travel plan, LLMs can only provide recommendations in text, while agents can realize the plan in action, such as checking the weather, buying tickets, and booking hotels.
+Second, agents exhibit obvious domain-specific characteristics, that is, an agent is good at a certain niche field.
+As a result, a task usually requires the collaboration of multiple agents, which may be located globally on the Internet (as shown in Figure 1).
+Under these conditions, agent communication is expected to become the foundation of the next communication era and the future AI ecosystem.
+It enables agents to find other agents with specific capabilities, access external knowledge, assign tasks, and engage in other interactions on the Internet.
+Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to its development.
+In November 2024, Anthropic proposed Model Context Protocol (MCP) [anthropic_mcp_2024, ray2025survey], a universal protocol that allows agents to communicate with external environments, such as datasets and tools.
 MCP quickly gained a great deal of attention in the recent few months.
-Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [214], Google [96], Microsoft [57], Amazon [25], Alibaba [10], and Tencent [267], and MCP’s package receives over 3 million weekly downloads [21].
-In April 2025, Google proposed Agent-to-Agent Protocol (A2A) [89], which enables seamless communication and collaboration among agents.
-Since its release, A2A has received extensive support from many enterprises, such as Microsoft [200], Atlassian [160], and PayPal [242].
-It can be seen that the breakthroughs in agent communication are bringing rapid and profound changes and will become an indispensable part of the AI ecosystem.
-However, the rapid development of agent communication also introduces complex security risks that could cause severe damage to the AI ecosystem.
-For example, the collaboration of cross-organization agents significantly enlarges the attack surface, causing severe security risks, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks.
+Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [openaiagentsdk], Google [googlemcp], Microsoft [microsoftmcp], Amazon [amazonmcp], Alibaba [alibabamcp], and Tencent [tencentmcp].
+In April 2025, Google proposed Agent-to-Agent Protocol (A2A) [A2A], which enables seamless communication among remote agents on the Internet.
+Since its release, A2A has received extensive support from many enterprises, such as Microsoft [microsoftA2A], Atlassian [AtlassianA2A], and PayPal [PayPalA2A].
+The market size of agents is expected to increase by 46% per year [agentmarket].
+However, the rapid development of agent communication also introduces complex security risks that could cause severe damage.
+For example, the communication of cross-organization agents significantly enlarges the attack surface, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks.
 Since the research related to agent communication is still in the nascent stage, it urgently needs a systematic review of the security problems existing in the complete agent communication lifecycle.
@@
 The contributions of this paper are as follows:
-We proposed the definition of agent communication for the first time, and classified it into three stages based on the communication object, which covers the entire lifecycle of agent communication.
-We comprehensively studied and classified the existing agent communication protocols.
-Besides, we thoroughly analyzed and categorized the related security risks for each communication stage, and detailedly discussed the targeted defense countermeasures.
+We propose the definition of agent communication for the first time.
+Then, we classify agent communication into three classes based on the characteristics of communication objects and propose a three-layer communication architecture for each stage, clarifying the functional boundaries for each communication part.
+This fills the gap where there was a lack of a structured technical framework for agent communication.
+We comprehensively studied and classified the existing 19 agent communication protocols.
+Besides, we thoroughly analyzed and categorized the related security risks for each communication stage and layer, and detailedly discussed the targeted defense countermeasures.
+This provides valuable insights for real-world deployments.
 We conduct experiments using MCP and A2A to help readers better understand the new attack surfaces brought by agent communication.
-The results show that attackers can easily cause severe damage to users’ systems and privacy.
+The results show that attacking agent communication can easily cause severe damage to the real world.
 We finally discuss the open issues and future research directions. We not only point out the much-needed techniques but also explain the demand for related laws and regulations.
@@
 | Moti. | Defi. | Clas. | Clas. | U-A | A-A | A-E | U-A | A-A | A-E | U-A | A-A | A-E | |||||
-| [170] | 2024 | Personal Agent | 4 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ |
-| [77] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
-| [163] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✗ | ✔ | ✓ – | ✗ | ✗ |
-| [143] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✗ |
-| [56] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✓ – | ✗ |
-| [108] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✗ |
-| [295] | 2025 | General IoA | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
-| [103] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
-| [317] | 2025 | Agent Comm. | 6 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
-| [296] | 2025 | Agent Secu. | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✔ | ✔ | ✓ – | ✔ | ✔ | ✗ |
-| [281] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
-| [297] | 2024 | Agent Secu. | 6 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
-| [75] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ |
-| [241] | 2025 | Agent Comm. | 7 | ✓ – | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ | ✗ |
-| [325] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
-| [95] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ |
-| [318] | 2025 | Agent Comm. | 8 | ✔ | ✗ | ✔ | ✔ | ✓ – | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ |
-| [147] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✔ |
-| [252] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ |
-| [233] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ |
-| [67] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
-| [111] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ |
+| [li2024personal] | 2024 | Personal Agent | 4 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ |
+| [gan2411navigating] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
+| [anonymous2025mind] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✗ | ✔ | ✓ – | ✗ | ✗ |
+| [ko2025seven] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✗ |
+| [deng2025ai] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✓ – | ✗ |
+| [he2025security] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✗ |
+| [wang2025internet] | 2025 | General IoA | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
+| [he2024emerged] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
+| [yan2025beyond] | 2025 | Agent Comm. | 6 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
+| [wang2025security] | 2025 | Agent Secu. | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✔ | ✔ | ✓ – | ✔ | ✔ | ✗ |
+| [wang2025comprehensive] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
+| [wang2025large] | 2024 | Agent Secu. | 6 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
+| [ferrag2025llm] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ |
+| [sarkar2025survey] | 2025 | Agent Comm. | 7 | ✓ – | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ | ✗ |
+| [yu2025survey] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
+| [hammond2025multi] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ |
+| [yang2025survey] | 2025 | Agent Comm. | 8 | ✔ | ✗ | ✔ | ✔ | ✓ – | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ |
+| [krishnan2025advancing] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✔ |
+| [singh2025survey] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ |
+| [ray2025review] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ |
+| [ehtesham2025survey] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
+| [hou2025model] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ |
+| [zhao2025mcp] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✔ |
 | This survey | Agent Comm. Secu. | / | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | |
 II-A Overview of Novelties in This Survey
-Table I summarizes the characteristics of the most relevant surveys and the differences between this survey and previous surveys. In summary, this surveys exhibit the following novelties:
+Table I summarizes the characteristics of the most relevant surveys and the differences between this survey and previous surveys. In summary, this survey exhibits the following novelties:
 This survey presents a comprehensive illustration of agent communication.
-Specifically, it explains why the current agent ecosystem needs communication (i.e., the predicament faced by a single agent, Section IV-A), gives the definition of agent communication (Section IV-B), and proposes a novel classification principle based on communication entity, which can cover the entire lifecyle of agent communication (Section IV-C).
-As a result, future studies can be included and categorized according to our survey seamlessly.
-This survey exhibits a comprehensive illustration of the existing protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A), including newly proposed and previously neglected protocols that have not been discussed by other surveys.
-Besides, we categorize these protocols based on their architecture and summarize corresponding characteristics, rather than mechanically listing each protocol.
+Specifically, it defines agent communication for the first time (Section IV-A), proposes a novel classification principle based on communication objects, which can cover the entire lifecycle of agent communication (Section IV-B), and uses a three-layered communication architecture to illustrate how each communication class is supported (Section IV-D).
+As a result, future studies, including protocols, attacks, and defenses, can be systematically organized.
+This survey exhibits a comprehensive illustration of the existing 19 protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A).
+Besides, we further categorize these protocols based on their characteristics, rather than mechanically listing each protocol.
 This organization method can allow any researchers interested in this field to quickly establish a preliminary but comprehensive understanding of agent communication.
-This survey makes an in-depth analysis of the discovered attacks and potential risks that have not been revealed for each agent communication stage.
-We discuss both the risks from compromised agents and those from malicious communication objects (i.e., users and environments), fully covering the entire agent communication lifecycle.
-Then, we thoroughly outline the possible defense countermeasures (Sections V-D, VI-C, and VII-D) that can make future agent communication more secure.
+This survey makes an in-depth analysis of the discovered attacks and potential risks that have not been revealed for each agent communication stage (Sections V-B, VI-B, VII-B).
+We clearly categorize each risk into different communication layers, which fully cover the entire agent communication lifecycle.
+Then, we thoroughly outline the possible defense countermeasures (Sections V-C, VI-C, and VII-C) that can make future agent communication more secure.
 This survey conducts experiments using MCP and A2A (Section VIII), the most popular agent communication protocols up to now.
@@
 This section can help readers better understand the new attack surfaces brought by agent communication.
+This survey comprehensively discusses the future direction of agent communication from perspectives of technique and law, which can provide practical benefit for the wide adoption of agent communication.
 II-B Selection Principles of the Most Relevant Surveys
@@
 This is because there have been many studies about multi-agent systems (MAS) before the emergence of LLMs.
-These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits very little to this survey.
+These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits little for this survey.
 Besides, surveys focusing on only LLMs instead of LLM-driven agents are also not listed in Table I (but we will draw on their valuable insights in other sections).
-This is because LLMs show significant differences from agents, for which we make a detailed illustration in Section III-C.
+This is because LLMs show significant differences from agents, for which we make a detailed illustration in Section III-B3.
 As a result, researching LLM-driven agents is the most important principle.
@@
 The second principle is that a survey should focus on or partially discuss agent communication, especially including some typical agent communication protocols such as MCP.
-This is because agent communication is very different from agent.
+This is because agent communication is very different from agents.
 However, if a survey satisfies the other two principles (i.e., LLM-driven agents and security), we still treat it as a relevant survey.
@@
 This is because we believe that the security risks of agents still have meaning to the security risks of agent communication.
-The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agent.
+The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agents.
 Relevance Score.
@@
 II-C Detailed Comparison with the Most Relevant Surveys
-In this section, we will detailedly compare the most relevant surveys in Table I with our survey.
-Survey [170] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computers.
+In this section, we will compare the most relevant surveys in Table I with our survey.
+Survey [li2024personal] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computers.
 It only partially mentions the security risks related to personal agents in a section.
@@
 It also does not discuss agent communication and related security risks and defenses.
-Survey [77] focuses on agent security instead of agent communication security.
-It is a security-specific paper, so its discussion of security is more comprehensive compared to [170].
+Survey [gan2411navigating] focuses on agent security instead of agent communication security.
+It is a security-specific paper, so its discussion of security is more comprehensive compared to [li2024personal].
 However, the main body of its discussion is about the interaction between user and agent (U-A), without enough consideration about agent-agent (A-A) or agent-environment (A-E) interaction, which have significantly different characteristics.
 Besides, it also does not include any protocols related to agent communication.
-Survey [163] also focuses on agent security instead of agent communication security, which is similar to [77].
+Survey [anonymous2025mind] also focuses on agent security instead of agent communication security, which is similar to [gan2411navigating].
 This survey focuses on single-agent systems and partially discusses multi-agent collaboration.
 It does not consider agent communication, related protocols, and enough security analyses about A-A and A-E.
-Survey [143] systematically summarizes seven security challenges for multi-agent systems.
+Survey [ko2025seven] systematically summarizes seven security challenges for multi-agent systems.
 As shown in Table I, its main focus is on A-A, and only partially discusses U-A and A-E, which is not comprehensive.
 Besides, it does not consider agent communication and related protocols.
-Survey [56] proposes four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E.
+Survey [deng2025ai] proposes four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E.
 Besides, it does not consider agent communication or any related protocols.
 The defense prospect is also limited.
-Survey [108] focuses on the security risks of U-A and A-E, such as malicious API.
+Survey [he2025security] focuses on the security risks of U-A and A-E, such as malicious API.
 It does not consider agent communication and related protocols.
 Besides, its security analyses are also not comprehensive enough.
-Survey [295] focuses on the fundamentals, applications, and challenges of IoA.
+Survey [wang2025internet] focuses on the fundamentals, applications, and challenges of IoA.
 Since its focus is different, agent communication and related security are only partially mentioned.
@@
 Besides, it also lacks the related illustration (such as definition and classification) of agent communication.
-Survey [103] also concentrates on the security of U-A, partially discussing A-E.
+Survey [he2024emerged] also concentrates on the security of U-A, partially discussing A-E.
 It does not mention agent communication and related protocols, as well as the risks of A-A.
-Survey [317] focuses on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view.
+Survey [yan2025beyond] focuses on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view.
 Besides, it only partially mentioned related security and did not discuss any communication protocols.
-Survey [296] focuses on the security of IoA.
+Survey [wang2025security] focuses on the security of IoA.
 It mentioned a few agent communication protocols (i.e., MCP, A2A, ANP, and Agora), neglecting many other important protocols.
@@
 According to our analyses, the security analyses (especially for U-A) are also not comprehensive enough.
-Survey [281] proposes the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization.
+Survey [wang2025comprehensive] proposes the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization.
 It does not focus on agent communication security.
 As a result, this survey did not give a clear illustration of agent communication, only mentioned a few protocols (i.e., MCP, A2A, ANP, and Agora), and partially discussed related threats and countermeasures.
-Survey [297] gives a comprehensive analysis of the security of agent networks.
+Survey [wang2025large] gives a comprehensive analysis of the security of agent networks.
 However, it does not include the discussion of communication protocols, and lacks enough security analyses of A-A and A-E.
-Survey [75] does not focus on security.
+Survey [ferrag2025llm] does not focus on security.
 Instead, it concentrates on evaluating LLMs and agents.
@@
 We can see that it does not give a detailed illustration of agent communication, enough coverage of protocols, or a comprehensive discussion about security.
-Survey [241] focuses on MCP, detailedly analyzed related architectures and applications.
+Survey [sarkar2025survey] focuses on MCP, detailedly analyzed related architectures and applications.
 It does not consider other communication protocols and only partially mentions security-related content.
-Survey [325] analyzes the threats of agents and divides them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E.
+Survey [yu2025survey] analyzes the threats of agents and divides them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E.
 However, its analyses are not comprehensive enough, and it does not mention any communication protocols.
-Survey [95] makes a comprehensive analysis of the risks for multi-agent systems.
+Survey [hammond2025multi] makes a comprehensive analysis of the risks for multi-agent systems.
 However, its focus only falls within A-A, not considering U-A, A-E, and related protocols.
-Survey [318] is one of the surveys with the highest relevance score because it focuses on agent communication and analyzes related protocols.
+Survey [yang2025survey] is one of the surveys with the highest relevance score because it focuses on agent communication and analyzes related protocols.
 However, it still has significant differences from our survey.
@@
 Third, its defenses prospect is limited.
-Survey [147] focuses on the influences of MCP.
+Survey [krishnan2025advancing] focuses on the influences of MCP.
 As a result, it lacks other protocols, the illustration of agent communication, and security-related discussion.
-Similarly, the survey [252] also focuses on MCP.
+Similarly, the survey [singh2025survey] also focuses on MCP.
 It lacks illustrations of other protocols and comprehensive security analyses.
-Survey [233] comprehensively introduces A2A, lacking discussion about other protocols and security analyses.
-Survey [67] detailed discussed MCP, A2A, ANP, and ACP(-IBM).
+Survey [ray2025review] comprehensively introduces A2A, lacking discussion about other protocols and security analyses.
+Survey [ehtesham2025survey] detailed discussed MCP, A2A, ANP, and ACP(-IBM).
 It also partially analyzed related security risks and defenses.
-However, there still lacks other protocols, in-depth security analyses, and systematic illustration of agent communication.
+However, there still lacks other protocols, in-depth security analyses, and a systematic illustration of agent communication.
 Hou et al.
-[111] discussed the security risks of MCP.
+[hou2025model] and Zhao et al.
+[zhao2025mcp] discussed the security risks of MCP.
 They did not consider other protocols and the high-level overview of agent communication.
-III LLM-Driven AI Agents
-In this section, we review the entire lifetime from LLM to LLM-driven AI agent. Our goal is to help beginners quickly figure out agents, their characteristics, relationships, and applications.
-III-A Large Language Model
-Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [211].
-Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [330], logic reasoning [225, 356, 300], code generation [344], and translation [222].
+III Preliminary: The Evolution of Communication and Agents
+In this section, we review the entire lifetime of communication and LLM-driven AI agents.
+III-A The Evolution of Communication
+As shown in Figure 3, we divide the evolution history of communication technology into four stages: (1) computer communication, (2) mobile communication, (3) IoT communication, and (4) agent communication.
+These four stages exhibit a close evolutionary relationship characterized by technological inheritance, progressive enhancement of capabilities, and expansion of application scenarios.
+Each stage lays the foundation for the subsequent one, while integrating the technological advantages of the previous stages.
+III-A1 Computer Communication Era
+As the starting point of modern communication, this era centered on the interconnection between fixed computer devices.
+In its initial stage, it relied on dedicated line connections.
+Later, with the popularization of standard protocols like TCP/IP [fall2012tcp], a global Internet architecture was formed.
+Its core value lies in building complete standard communication protocols, thus realizing the remote transmission between different types of devices and breaking down the information barriers caused by geographical distances.
+The main subjects of communication at this stage were hardware devices, only meeting the basic need of data transmission.
+III-A2 Mobile Communication Era
+With the popularization of mobile communication technologies (from 2G to 6G) [andrews2014will] and intelligent terminals (such as mobile phones), communication scenarios have expanded from fixed computer communication to mobile communication.
+Compared to the computer communication era, this era has broken through the constraints of time and space, enabling real-time interactions between moving people and devices.
+The core features are the portability of mobile terminals and the of communication, which has driven the explosive growth of scenarios such as social networking, e-commerce, and mobile office.
+III-A3 IoT Communication Era
+With the emergence of Internet of Things (IoT) technology [atzori2010internet], the subjects of communication have further expanded to include various IoT devices (such as industrial sensors, cameras, smartwatches, and smart furniture).
+Communication at this stage is characterized by “thing-to-thing” connectivity.
+Through customized protocols, real-time perception of the physical world is realized.
+As a result, the communication has shifted from “human-computer interaction” to environmental perception, providing underlying data support for fields such as industrial automation and smart cities.
+However, interactions are still mainly focused on data collection and instruction issuance, lacking autonomous decision-making capabilities.
+III-A4 Agent Communication Era
+Agents’ unprecedented capabilities of reasoning and using tools are expected to make them the foundation of the next communication era.
+Agent communication has achieved a qualitative leap from device interconnection to collaboration among intelligent entities.
+Unlike the previous three stages, its communication subjects are LLM-driven agents with capabilities of autonomous perception, memory, reasoning, and decision-making.
+The communication goal is no longer simple data transmission, but rather revolves around complex tasks involving intention alignment, task decomposition, resource coordination, and result aggregation.
+It can independently discover collaboration partners, negotiate interaction rules, and dynamically adjust communication strategies.
+This stage breaks the limitations of devices and scenarios and is expected to become the fundamental support for the future AI ecosystem.
+III-B LLM-Driven AI Agents
+III-B1 Large Language Model
+Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [nvidiallm].
+Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [yuan2022wordcraft], logic reasoning [qin2023chatgpt, zhong2023can, wei2022chain], code generation [zhang2023planning], and translation [peng2023towards].
 These remarkable performances can be attributed to two major factors.
-One is that LLMs are built upon a powerful architecture known as the Transformer [276], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input.
+One is that LLMs are built upon a powerful architecture known as the Transformer [vaswani2017attention], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input.
 The other key factor, perhaps the most important one, is the massive scales of LLMs that far exceed traditional AI models.
-When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [299], referring to unexpected capabilities that do not appear in smaller models.
+When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [wei2022emergent], referring to unexpected capabilities that do not appear in smaller models.
 As shown in Table II, the parameter scale of an LLM can be hundreds or thousands of times that of traditional AI models.
@@
 | FNN | MLP | 1990s | 100K |
-| FNN | LeNet-5 [157] | 1998 | 60K |
-| RNN | Elman Net [69] | 1990 | 100K |
-| LSTM | LSTM [109] | 1997 | 1–10M |
-| CNN | ResNet-50 [104] | 2015 | 25M |
-| CNN | AlexNet [148] | 2012 | 60M |
-| CNN | VGG-16 [251] | 2014 | 138M |
-| GAN | DCGAN [228] | 2016 | 4M |
-| GNN | GCN [142] | 2017 | 23K |
-| Autoencoders | DAE [277] | 2008 | 100K |
-| Autoencoders | VAE [141] | 2013 | 1M |
-| Transformer | GPT-3 [29] | 2020 | 175B |
-| Transformer | PaLM [43] | 2022 | 540B |
-| Transformer | GPT-4 [5] | 2023 | 1T |
-| Transformer | DeepSeek-V3 [176] | 2024 | 671B |
-| Transformer | DeepSeek-R1 [94] | 2025 | 671B |
-III-B LLM-Driven AI Agents
+| FNN | LeNet-5 [lecun1998gradient] | 1998 | 60K |
+| RNN | Elman Net [elman1990structure] | 1990 | 100K |
+| LSTM | LSTM [hochreiter1997lstm] | 1997 | 1–10M |
+| CNN | ResNet-50 [he2016deep] | 2015 | 25M |
+| CNN | AlexNet [krizhevsky2012imagenet] | 2012 | 60M |
+| CNN | VGG-16 [simonyan2014very] | 2014 | 138M |
+| GAN | DCGAN [radford2015unsupervised] | 2016 | 4M |
+| GNN | GCN [kipf2016semi] | 2017 | 23K |
+| Autoencoders | DAE [vincent2008extracting] | 2008 | 100K |
+| Autoencoders | VAE [kingma2013auto] | 2013 | 1M |
+| Transformer | GPT-3 [brown2020language] | 2020 | 175B |
+| Transformer | PaLM [chowdhery2022palm] | 2022 | 540B |
+| Transformer | GPT-4 [achiam2023gpt] | 2023 | 1T |
+| Transformer | DeepSeek-V3 [liu2024deepseek] | 2024 | 671B |
+| Transformer | DeepSeek-R1 [guo2025deepseek] | 2025 | 671B |
+III-B2 LLM-Driven AI Agents
 Figure 4 illustrates a typical architecture of LLM-driven agents.
@@
 To automatically finish a specified task, agents need the ability to perceive the real-world environment.
-For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [195, 190].
+For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [mao2023language, ma2024dolphins].
 The type of perception ability depends on the domain for which the agent is designed.
-For instance, an autonomous driving agent needs the ability of visual or radar perception [258, 321], while a code-generating agent may not require such functions [116, 124].
+For instance, an autonomous driving agent needs the ability of visual or radar perception [sun2020mimo, yao2025exploring], while a code-generating agent may not require such functions [huang2023agentcoder, ishibashi2024self].
 Memory module.
 The processing of real-world tasks also requires a strong memory.
-Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [102, 191, 328].
-This usually requires external storage resources to assist the brain, such as databases or memory sharing [80, 82].
+Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [hatalis2023memory, maharana2024evaluating, yu2024finmem].
+This usually requires external storage resources to assist the brain, such as databases or memory sharing [gao2024memory, gao2025efficient].
 In contrast, LLMs do not have such excellent memory ability.
-Their memory is short-term, which only lasts for rounds of conversations [284, 357].
+Their memory is short-term, which only lasts for rounds of conversations [wang2025recursively, zhong2024memorybank].
 Reasoning and planning module.
 LLM acts as the brain of agents due to its excellent capability of reasoning and planning.
-It intercepts the instructions from users and automatically decomposes the received task into multiple feasible steps [291, 127, 139, 253].
-Then, it selects the best plan from different candidates [136, 354, 113].
-Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [361, 275, 234, 305].
+It intercepts the instructions from users and automatically decomposes the received task into multiple feasible steps [wang2025tdag, jeyakumar2024advancing, khot2022decomposed, singh2025adaptbot].
+Then, it selects the best plan from different candidates [kannan2024smart, zhao2023large, hu2023tree].
+Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [zhou2024isr, valmeekam2023can, renze2024self, wuproposal, 11209370].
 For example, when the autonomous driving module finds that the barrier is closer, it will change the plan to slow down or detour.
 Tool module.
-The tool module is responsible for deeply integrating external resources with the cognitive capabilities of the agent, enabling it to perform complex operations beyond the native capabilities of LLM [331, 304, 167, 186].
-For example, through predefined functional interfaces and protocols, a math agent is able to invoke the external computation libraries and symbolic solvers to help it solve mathematical problems [90].
+The tool module is responsible for deeply integrating external resources with the cognitive capabilities of the agent, enabling it to perform complex operations beyond the native capabilities of LLM [yuan2024easytool, wu2024avatar, li2024review, lu2024toolsandbox].
+For example, through predefined functional interfaces and protocols, a math agent is able to invoke the external computation libraries and symbolic solvers to help it solve mathematical problems [gou2023tora].
 Action module.
 The action module is the core hub for interaction with the environment.
-It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [290, 345].
+It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [wang2024executable, zhang2024towards].
 This module ensures the executability of instructions through structured output control.
@@
 | Security | Better | Worse |
-III-C Comparison Between Agents and LLMs
+III-B3 Comparison Between Agents and LLMs
 Table III illustrates the advantages of agents over LLMs on different metrics. Overall, agents have many advantages over LLMs, except for security.
@@
 They are unable to plan or execute tasks independently.
-Besides, the response quality highly relies on the prompt skill [302, 85, 203, 178, 68, 33, 360], which seriously affects the user experience.
+Besides, the response quality highly relies on the prompt skill [white2023prompt, giray2023prompt, mesko2023prompt, liu2023pre, ekin2023prompt, chen2023unleashing, zhou2022learning], which seriously affects the user experience.
 In contrast, agents possess independent capabilities for task decomposition, strategy adjustment, and external tool invocation, which breaks through the passive mode of LLMS and is highly autonomous.
 Flexible Multimodal interaction.
-LLMs have limited capability of handling multimodal inputs, such as text and pictures [240, 338, 342, 360, 155, 363].
+LLMs have limited capability of handling multimodal inputs, such as text and pictures [sarch2024vlm, zhang2024vision, zhang2021vinvl, zhou2022learning, laurenccon2024matters, zhu2023minigpt].
 Besides, their outputs are also mainly single-modal (e.g., text-only or picture-only), lacking the ability to actively invoke tools to perform physical actions or generate multimodal content.
@@
 Abundant Tool invocation.
-LLMS usually passively invoke a single tool (such as Function Calling [213]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [337]).
+LLMs usually passively invoke a single tool (such as Function Calling [openai_functioncalling]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [zhang2023ecoassistant]).
 In contrast, agents have the capability of active decision-making.
-They can independently select, combine, and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [107].
+They can independently select, combine, and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [he2025pasa].
 Better Hallucination inhibition.
-LLMs suffer from a serious problem called hallucination, which refers to that LLMs are likely to generate non-existent knowledge [348, 93, 272, 118, 315, 169].
+LLMs suffer from a serious problem called hallucination, which refers to that LLMs are likely to generate non-existent knowledge [zhang2023siren, gunjal2024detecting, tonmoy2024comprehensive, huang2025survey, xu2024hallucination, li2023evaluating].
 LLMs mainly rely on the knowledge internalization of training data, making them prone to hallucinations when facing uncovered domains or outdated information.
-In contrast, agents are able to reduce the error rate by integrating multiple techniques such as Retrieval Augmentation Generation (RAG) [84, 162, 353] or other methods, which can align the action of agents [268, 79].
+In contrast, agents are able to reduce the error rate by integrating multiple techniques such as Retrieval Augmentation Generation (RAG) [gao2023retrieval, lewis2020retrieval, zhao2024retrieval] or other methods, which can align the action of agents [tennant2024moral, gao2024aligning].
 Dynamic adaptability.
 Essentially, LLMs are static models whose knowledge is fixed at the training phase.
-Although techniques such as fine-tuning [334, 175, 115] or model editing [322, 288, 341, 283, 171] reduce the training cost significantly, LLMs still cannot adapt to real-time events well.
+Although techniques such as fine-tuning [zhang2024scaling, lin2024data, hu2023llm] or model editing [yao2023editing, wang2024knowledge, zhang2024comprehensive, wang2023easyedit, li2023unveiling] reduce the training cost significantly, LLMs still cannot adapt to real-time events well.
 In contrast, agents are equipped with techniques like online web search, database query, or real-time sensors, which enable them to dynamically adapt to the changes of real-time environments and information.
@@
 In contrast, since agents are endowed with the ability to invoke tools, they can cause substantial damages to the real world, including but not limited to maliciously/wrongly operating machines, poisoning databases, and paralyzing the system.
+For example, attackers can induce agents to visit malicious websites [kong2025web], causing a wide range of subsequent threats.
 As a result, it is necessary to concentrate more on the security of agents.
-III-D Agent Applications
+III-B4 Agent Applications
 Due to the strong advantages that agents have shown, related applications are booming.
@@
 Agents are increasingly embedded into the research workflow, enhancing ideation, automation, and discovery.
-Their contributions span multiple disciplines, such as mathematics [161, 55, 312, 285], chemistry [239, 50, 42, 27], biological sciences [313, 309, 177], and materials Science [151, 219, 197]
+Their contributions span multiple disciplines, such as mathematics [lei2024macm, deng2024flowdpo, xie2024mathlearner, wang2025malot], chemistry [ruan2024automatic, darvish2024organa, chiang2024llamp, boiko2023autonomous], biological sciences [xin2024bia, xiao2024cellagent, liu2024tais], and materials Science [kumbhar2025hypothesis, uddin2024ai, kulik2024future]
 Technical and Engineering Systems.
 Agents play a growing role in engineering domains, improving automation, systems, and software intelligence.
-For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [280, 37, 125, 114, 138, 187].
-Besides, agents are also popular in game development and simulation [201, 255].
-Embodied intelligence is also another hot topic [36, 194].
+For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [wang2025training, chen2025locagent, jain2025r2e, hu2025llm, gatelens2025, lu2025uxagent].
+Besides, agents are also popular in game development and simulation [npc_behavior_unity_ml_agents, li2024multiagent].
+Embodied intelligence is also another hot topic [chen2023scalable, zhao2023roco].
 Social Governance and Public Services Agents are increasingly deployed in sectors focused on public service and human welfare.
-For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [270, 121, 199, 269].
-Besides, other fields, such as financial services [164, 319, 212, 329, 98, 97, 73], education [287, 259, 193, 59, 64, 206], and healthcare [35, 298, 74, 238, 23, 123, 289, 159, 346], are also actively integrating agents into their respective practices.
+For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [liu2024efficient, li2024ai, weng2024llms, matthews2024generative].
+Besides, other fields, such as financial services [li2024econagent, yang2025twinmarket, okpala2025agentic, yu2024fincon, han2024enhancing, zhang2025finsphere, fatouros2025marketsenseai], education [wang2024ai_education_review, swargiary2024impact, maity2024generative, dickey2024gaide, dumbuya2024personalized, moller2024revolutionising], and healthcare [chen2024cod, wang2025medagent, feng2025m3builder, rose2025meddxagent, averly2025liddia, li2024drugagent, lin2024patentagent, zhang2025psyche, zhang2024psydraw], are also actively integrating agents into their respective practices.
 Overall, it can be seen that agents are being widely applied in all walks of life, greatly promoting the development of productivity.
 More importantly, the application of agents is still in its infancy and has an even greater space for development in the future.
-It is estimated that the agent market will grow at a rate of 40% annually and is expected to exceed 216.8 billion dollars by 2035 [15].
-III-E Takeaways
+It is estimated that the agent market will grow at a rate of 40% annually and is expected to exceed 216.8 billion dollars by 2035 [agengmarket2].
+III-C Takeaways
 Agents show multiple advantages over LLMs on multiple metrics, such as richer perception ability, stronger learning ability, and higher adaptivity.
@@
 It can be foreseen that they will ultimately become indispensable components of future production ecosystems and daily life.
-However, agents show worse security than LLMs due to their capability of executing tools.
+However, agents show worse security than LLMs due to their capability to execute tools.
 As a result, studying the security of agent communication is significant to the AI ecosystem.
-IV Agent Communication Overview
-IV-A Motivation: The Demand for Agent Communication
-Although the advantages of agents in various fields have become increasingly obvious, their development has also encountered new obstacles, which have given rise to the demand for agent communication.
-Conflicting development trends.
-The first reason derives from the fundamental conflict between the refined development of agents and the abstract demands of users.
-With the in-depth and specialized evolution of agents towards vertical fields (such as medical diagnostic agents, financial risk control agents, and industrial control agents), their capability boundaries are becoming increasingly refined.
-However, users’ usage habits exhibit opposite characteristics: they tend to input simple and abstract instructions (such as “plan a cross-border travel”) to trigger the execution of complex tasks.
-It is hard for a domain-specific agent to finish such abstract instructions independently.
-Besides, this trend of users’ habits is hard to reverse.
-This is because people always prefer applications that are easy to operate rather than those that require cumbersome usage steps.
-The latter is at a disadvantage in the market competition.
-Usually, for each additional operation step, the user churn rate increases by 10% - 20%.
-Therefore, agents should not only NOT ask users to make changes, but also cater to such demands of users, which is contrary to the development direction of agents.
-Closed ecosystem.
-The current mainstream multi-agent systems adopt a closed ecosystem design and rely on private interaction mechanisms, forming a rigid technical barrier.
-This development paradigm severely limits the dynamic collaboration capabilities with external systems.
-For example, external agents cannot be actively discovered or invoked, making it extremely hard for cross-platform collaboration.
-Besides, such a closed ecosystem further triggers a decline in systemic effectiveness.
-On one hand, agents lose scalability due to binding private tools.
-On the other hand, the obstruction of cross-platform knowledge sharing suppresses the intelligence of agents.
-As a result, it is necessary to propose agent communication mechanisms that integrate agents from different communities.
-IV-B Agent Communication Definition
-To tackle these conflicts, agent communication is urgently demanded.
+IV Agent Communication: Definition and Classification
+IV-A Agent Communication Definition
+To tackle the capability limitation of a single agent, agent communication is urgently demanded.
 Specifically, agents need to collaborate with a series of external entities to finish user tasks.
@@
 Communication behaviors satisfying the above conditions can be regarded as agent communication.
-IV-C Agent Communication Classification
-Based on the object of communication, we divide agent communication into three classes: user-agent, agent-agent, agent-environment.
-We will use Figure 4 as a typical example to systematically overview the complete lifecycle of agent communication.
+IV-B Agent Communication Classification
+As shown in Figure 5, we divide agent communication into three main classes: user-agent interaction, agent-agent communication, agent-environment communication.
+Each communication class is further supported by a three-layered communication architecture: L1: data transmission layer, L2: interaction protocol layer, and L3: semantic interpretation layer.
+Three communication classes.
+The three communication classes are categorized based on the communication object of agents.
+The inherent advantage of this classification is that it groups communications with similar characteristics and security risks.
+For instance, user-agent interaction is naturally multimodal, which makes it particularly vulnerable to manipulation through prompt engineering.
+In contrast, these risks are relatively less prominent in agent-environment and agent-agent communication.
+Three communication layers for each stage.
+Building on the above classification, we further introduce a three-layered communication architecture structured by communication functionality.
+Such architecture offers developers two key advantages: it clarifies (1) how each layer supports the communication process, and (2) from which layer the security vulnerabilities originate, allowing for precise risk assessment.
+For example, man-in-the-middle attacks typically occur more often on L2 or L3.
+IV-C Three Agent Communication Classes
 IV-C1 User-Agent Interaction
@@
 For example, in step 5, the travel agent searches the popular tourist spots through its database or searches online blogs.
-Advantages of this classification method.
+Advantages of this classification.
 Different entities have essentially differentiated capability characteristics and attack surface attributes.
@@
 As a result, classifying agent communication by entity types can directly cluster major vulnerability types and defense strategies that have similar characteristics, providing a structured analysis paradigm for future security research.
-IV-D Organization of the Following Sections
-As shown in Figure 5, in the following sections, we will discuss agent communication and related security.
-In Section V, we will introduce user-agent interaction.
-Specifically, the risks in this process are divided into two categories: risks from malicious users against benign agents (Section V-B) and risks from compromised agents against benign users (Section V-C).
-The possible defense countermeasures against malicious users are discussed in Section V-D, while those against malicious agents are discussed uniformly in Section VI-C.
-In Section VI, we will classify existing protocols for agent-agent communication.
-Then, we discuss the risks in Section VI-B and defenses in Section VI-C.
-Please note that the proposed defenses against compromised agents in this section are also applicable to the risks in Section V-C and Section VII-C.
-In Section VII, we first show the related protocols in Section VII-A.
-Then, the risks in this process are also classified into two categories: risks from malicious environments against benign agents (Section VII-B) and risks from compromised agents against benign environments (Section VII-C).
-The defenses for compromised agents can be referred to Section VI-C.
-As a result, we only discuss the defenses for the malicious environments in Section VII-D.
-Please note that the risks/defenses we focus on are specific for agent communication instead of LLMs.
-This is because agents and agent communication still have significant differences from LLMs (as we have discussed in Section III).
-For example, in the agent collaboration process, the accumulation of tiny benign deviations on each step may lead to an intolerable risk in the final result.
-Miehling et al.
-[204] also point out that over-focusing on the capabilities of a single model can result in neglecting the emergent behaviors at the system level and underestimating the true risks.
-As a result, for LLM-targeted attacks that have valuable inspiration for agent security, we will discuss the typical representatives of them instead of listing all papers exhaustively.
-IV-E Takeaways
-Conforming to the usage trends of users, the collaboration of multiple agents has become a clear development direction.
-In this context, agent communication becomes the foundation of the future AI ecosystem.
-Based on the communication entity, we classify agent communication into three different types: user-agent interaction, agent-agent communication, and agent-environment communication.
-This classification can naturally distinguish communication with similar vulnerability characteristics, providing a structured research paradigm for future studies.
+IV-D A Three-Layer Communication Support
+Although Section IV-C provides a clear classification of agent communication, the underground communication details are unclear.
+To address this problem, we provide a three-layered communication architecture to illustrate how they support agent communication.
+IV-D1 L1: Data Transmission Layer
+This layer is responsible for the data transmission of agent communication.
+It could be a layer of the traditional TCP/IP stack, or developers can design a unique data transmission protocol according to their real demands and requirements.
+For example, the communication between remote agents can be built upon HTTPS.
+In this context, the Data Transmission Layer in agent communication is the Application Layer in the TCP/IP stack.
+For agent communication within the same local area network, developers can directly use IP or TCP to transmit packets due to the secure environment.
+In this context, the Data Transmission Layer could be the Network Layer (IP) or the Transport Layer (TCP) in the TCP/IP stack.
+Overall, the primary function of this layer is to handle the establishment and termination of agent communication connections and ensure the transmission of packets.
+It operates on raw data streams, agnostic to the content or semantic meaning of the transmitted information.
+IV-D2 L2: Interaction Protocol Layer
+This layer defines the communication modes between entities (e.g., agents may communicate in a distributed or a centralized mode).
+It ensures that the communication follows a series of well-defined principles before the content is transmitted.
+For example, this layer defines authentication services (validating agent identity via credentials or keys), session managers (tracking the state of a conversation), and authorization logic (determining access rights to specific tools or data).
+Overall, this layer ensures that the communication process adheres to specified rules, determining who is communicating and what they are permitted to do.
+IV-D3 L3: Semantic Interpretation Layer
+This layer serves as the cognitive core of agent communication, where the underlying intent, relation, and logic of messages are semantically parsed.
+It interprets multimodal inputs, e.g., text, images, and voice, into rich internal representations that capture contextual meaning, enabling agents to perform reasoning, planning, and reflection for autonomous task execution.
+This layer converts raw information into coherent, context-aware communication that supports intelligent understanding and decision-making.
+Since this layer mainly relies on the understanding capability of LLMs, its form does not need specification, and thus, there have not been dedicated protocols designed for it.
+However, this layer is also important because it is a unique attack surface.
+Advantages.
+The primary advantage of this layered framework lies in its clear separation of functionality and security.
+First, it enforces a distinct division of functions, enabling researchers to understand and design agent communication in a more structured and efficient manner.
+Second, it allows developers to precisely locate the specific layers on which failures or vulnerabilities occur.
+For instance, MITM usually happens on the Data Transmission Layer (L1), while prompt injection attacks are fundamentally a vulnerability at the Semantic Interpretation Layer (L3).
+This diagnostic precision is crucial for developing targeted mitigation strategies.
+IV-E Organization of the Following Sections
+Logic.
+The following sections of this paper are structured based on the three communication classes, i.e., user-agent, agent-agent, and agent-environment.
+In each class, we will provide a detailed analysis of how it is supported by the three-layered communication architecture and how security risks arise.
+Organization. As shown in Figure 6, in the following sections, we will discuss agent communication and its security.
+IV-F Takeaways
+In this section, we clarify some core concepts and a structured framework of agent communication, laying a foundational theoretical basis for subsequent security analysis.
+First, we present a clear definition of agent communication.
+Second, based on communication objects, we classify agent communication into three classes.
+This classification naturally clusters scenarios with similar vulnerability characteristics.
+Third, we also propose a three-layered communication architecture that supports each communication stage.
+This architecture not only clarifies functional divisions but also enables risk localization.
+This entire structured framework not only helps researchers systematically understand how agent communication works but also benefits relevant deployments and studies.
 V User-Agent Interaction
-In this section, we will introduce the current user-agent interaction protocols, their security risks, and future defense strategies.
-V-A Protocols
-PXP.
-PXP protocol [256] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields.
+The organization of this section is shown in Figure 7.
+V-A Interaction Protocols
+The user-side client communicates with the remote agent server to deliver user tasks and receive corresponding responses.
+This process typically relies on well-established networking stacks and techniques without requiring novel mechanisms.
+For example, the user’s client uses Domain Name System (DNS) to identify the IP address of the remote agent server and then establishes an HTTPS connection for secure data transmission.
+Since such protocols (e.g., IP and HTTP) are mature and standardized, this paper only focuses on those newly proposed for agent communication.
+As we analyzed in Section IV-D, the existing newly-proposed user-agent interaction protocols are about L2.
+AG-UI [pajo2025ag] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism.
+As shown in Figure 8, the front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols).
+The client invokes the RUN interface of the protocol layer to send requests to the agent.
+When the agent processes the request, it generates a streaming event and returns it to the AG-UI client.
+Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events.
+The AG-UI client handles different types of responses by subscribing to the event stream.
+Agents can transfer context between each other to maintain the continuity of the conversation.
+All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication.
+AG-UI also focuses on the semantic interpretation in user-agent interaction, which is mainly undertaken by AI models (e.g., LLMs).
+These models are responsible for parsing and interpreting the underlying intent of user instructions.
+The process is inherently multimodal, allowing users to provide inputs in various forms, including but not limited to text, images, and videos.
+PXP protocol [srinivasan2024implementation] focuses on building an interactive system between human experts and agents in data analysis tasks, targeting issues in complex scientific, medical, and other fields.
 It is worth mentioning that PXP is not customized for LLM-driven agents, but we think its design has inspirational meaning for agent communication.
@@
 The effectiveness of PXP has been verified in the scenarios of radiology and drug discovery.
-Spatial Population Protocols.
-The Spatial Population Protocols is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems.
+Spatial Population Protocol is a minimalist and computationally efficient distributed computing model, specifically designed to solve the Distributed Localization Problem (DLP) in robot systems.
 Similar to PXP, strictly speaking, this work is not designed for LLM-driven agent systems.
@@
 Through the multi-contact epidemic mechanism, leader election, and self-stabilizing design, it enables n anonymous robots to achieve efficient localization from their respective inconsistent coordinate systems to a unified coordinate consensus, providing a scalable framework for robot collaboration in dynamic environments.
-AG-UI.
-AG-UI [217] realizes the communication between users (front-end applications) and agents based on the client-server architecture and completes the communication process by adopting the event-driven mechanism.
-As shown in Figure 6, the front-end application connects to agents through the AG-UI client (such as a common communication client that supports server-sent events or binary protocols).
-The client invokes the RUN interface of the protocol layer to send requests to the agent.
-When the agent processes the request, it generates a streaming event and returns it to the AG-UI client.
-Event types include lifecycle events (such as start of run, completion of run), text message events (transmitted in segments by start, content, and end), tool call events (passed in the order of start, parameters, and end), and state management events.
-The AG-UI client handles different types of responses by subscribing to the event stream.
-Agents can transfer context between each other to maintain the continuity of the conversation.
-All events follow a unified basic event structure and undergo strict type verification to ensure the reliability and efficiency of communication.
-Besides, please note that the previous survey [318] regards CrowdES [24] as an agent-related protocol.
-After our careful study, we think CrowdES is a framework for generating and evaluating simulated crowds and real-world crowds, which is not suitable for discussion in the field of agent communication.
-Therefore, we will not list it in this paper.
-V-B Security Risk Analysis: Malicious Users Against Benign Agents
-| Stage | Risk Type | Representative Threats | Attack Characteristics |
-| Malicious users against benign agents | Text-Based Attacks | Prompt Injection | Manipulate agent behavior by malicious prompts, not modifying model |
-| Jailbreak Prompts | Bypass security constraints via role-play, obfuscation, multi-turn tricks | ||
-| Multimodal Attacks | Image-Based Attacks | Embed adversarial cues in visual inputs while text remains benign | |
-| Audio-Based Attacks | Inject commands via adversarial waveforms or speech mimicry | ||
-| Privacy Leakage | Inference Attacks | Disclosure of agent information, such as topological structure | |
-| Denial of Service (DoS) | Overthink Attacks | Forces agent into unnecessary reasoning, repeat actions, or overthinking | |
-| Compromised agents against benign users | Violation of User Privacy | Exposure of Personal Information | Expose a comprehensively aggregated user profile |
-| Behavioral and Psychological Profiling | Build detailed behavioral or psychological profiles from innocuous conversations | ||
-| Psychological and Social Manipulation | Belief and Opinion Shaping | Manipulate user’s worldview, influence voting behavior, or radicalize beliefs | |
-| Social Engineering | Leverage the knowledge of the user’s communication style, vocabulary, and relationships to conduct highly convincing impersonation attacks | ||
-| Harmful Tasks | Economic Manipulation | Inflict subtle yet significant damage in professional or economic contexts | |
-| Behavioral and Psychological Profiling | Build unauthorized profiles and infer highly sensitive attributes |
-According to our analysis, the inputs from users show a significant multimodal characteristic.
-For example, AG-UI explicitly illustrates its support for multimodal content, such as text, images, and videos.
-Consequently, we find that the security risks in user-agent interaction primarily stem from these insecure inputs.
-V-B1 Text-Based Attacks
-In user-agent interaction, attackers can manipulate model behavior or bypass safety mechanisms by crafting malicious prompts.
-These attacks do not rely on modifying model parameters or architecture.
-Instead, they are carried out through natural language inputs, making them highly stealthy and applicable.
-Due to the diversity of linguistic forms and the indirectness of semantics, such attacks can effectively bypass safety mechanisms, posing significant security risks in real-world applications.
-These attacks can be broadly categorized into two types: prompt injection and jailbreak attacks.
-Prompt Injection refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources.
+| Layer | Threat Source | Risk Category | Representative Threats | Attack Characteristics |
+| L1 | External Attackers | R1: Violations of Confidentiality | Eavesdropping | Passive interception of unencrypted traffic. |
+| Traffic Analysis | Infers user activity by analyzing metadata of encrypted traffic. | |||
+| R2: Violations of Integrity | Man-in-the-Middle (MITM) Attack | Actively alters communication by modifying communication packets. | ||
+| Replay Attacks | Captures and re-sends valid data packets to trigger unauthorized actions. | |||
+| R3: Violations of Availability | Denial-of-Service (DoS) Attack | Overwhelms the agent’s network with traffic, making the service unavailable to legitimate users. | ||
+| L2 | External Attackers | R1: Identity Spoofing | Credential Theft & Session Hijacking | Impersonates a user with stolen credentials to take over their agent. |
+| R2: Session Manipulation | State Corruption Attack | Sends logically inconsistent but validly typed events to desynchronize agent state and bypass security checks. | ||
+| R3: Authority Abuse | Cross-Agent Privilege Escalation | Exploits inter‑agent trust to gain privileges beyond the user’s authorization. | ||
+| L3 | Malicious User | R1: Text-Based Attacks | Prompt Injection | Controls an agent through adversarially crafted inputs. |
+| Jailbreak Attack | Bypasses safety measures to produce harmful or restricted content. | |||
+| Privacy Leakage | Extracts internal data using crafted queries. | |||
+| Exhaustion Attack | Overloads the agent with excessive work to cause failure. | |||
+| R2: Multimodal Attacks | Image-Based Attacks | Hides malicious instructions in images to bypass text‑based safety filters. | ||
+| Audio-Based Attacks | Injects commands via adversarial audio waveforms or synthesized speech. | |||
+| Compro- mised Agent | R3: Violation of User Privacy | Exposure of Personal Information | Exfiltrates user profiles containing PII, financial, and conversation data. | |
+| Behavioral & Psychological Profiling | Infers sensitive user traits from casual conversations without consent. | |||
+| R4: Psychological & Social Manipulation | Belief and Opinion Shaping | Subtly injects biased content to shape a user’s worldview. | ||
+| Sophisticated Social Engineering | Uses detailed user knowledge to execute convincing impersonation attacks. | |||
+| R5: Execution of Malicious Tasks | Economic Manipulation | Covertly sabotages work or leaks confidential business data. | ||
+| Malicious Guidance | Provides harmful instructions, including malware creation or unsafe advice. |
+V-B Security Risks
+Our analysis of this scenario reveals a distinct distribution of threats across the three layers.
+While foundational risks exist at L1 and L2, the most predominant threats emerge at L3.
+This is because L1 and L2 threats are common to many systems and often mitigated by standard cryptographic and authentication methods.
+In contrast, L3 introduces a fundamentally different threat: validating the semantic intent of the payload.
+This threat is profound because agents are designed for instruction-following, creating an inherent trust conflict between obedience and safety.
+Additionally, traditional security mechanisms are blind to this conflict.
+V-B1 Risks from Data Transmission Layer (L1)
+This layer is responsible for transmitting data between the user client and the agent’s backend service.
+Security risks at this layer primarily stem from potential risks during data transmission, as shown in Figure 9.
+R1: Violations of Confidentiality.
+This class of risks aims to expose sensitive information about the user’s interaction without consent.
+Even if the communication is not altered, the leakage of the content and patterns of communication can be highly damaging.
+Eavesdropping.
+On insecure communication channels, such as those using plain HTTP or insecure WebSockets, an attacker can intercept unencrypted traffic between a user and an agent [waheed2022empirical].
+This leads to the leakage of personal or proprietary information.
+For instance, when a user consults a healthcare agent about their medical symptoms, an eavesdropper could capture the entire dialogue, resulting in a severe breach of sensitive personal health information.
+This type of attack compromises confidentiality without affecting the integrity of the user-agent interaction.
+Traffic Analysis.
+Even when communication is encrypted, attackers can perform traffic analysis to infer sensitive information.
+By observing the timing, size, and frequency of data packets, an adversary can deduce the nature of the user’s interaction [shen2022machine].
+Besides, attackers can also infer the agent type, number, and scale based on the captured traffic [zhang2025exposing].
+R2: Violations of Integrity.
+Integrity attacks involve the unauthorized modification or replay of data in transit.
+The goal is to deceive either the user or the agent by altering the legitimate flow of communication, leading to unintended and potentially harmful outcomes.
+Man-in-the-middle (MITM) Attack.
+Attackers can intercept the communication channel between users and agents [shaikh2025prompts].
+Even if the connection is protected (e.g., by HTTPS), the design vulnerabilities of related security protocols can still cause significant damage [calzavara2019postcards].
+For instance, the content of a user’s benign request can be replaced with a malicious attack prompt, which compromises the integrity of the user-agent interaction.
+Replay Attacks.
+This attack captures a valid data packet and re-sends it to cause a repeated action.
+The integrity of the communication is violated because a single, authorized command is duplicated without the user’s consent [wang2025security].
+For example, an attacker could capture the encrypted packet corresponding to a user instructing a financial agent to “buy 100 shares of company A.” By replaying this packet, the attacker can trigger unauthorized purchases, leading to property loss.
+R3: Violations of Availability.
+This class of risk represents the Denial-of-Service (DoS) attacks, whose objective is to render the service unavailable to legitimate users through a volumetric attack.
+In this context, a malicious user floods the agent’s network endpoint with an overwhelming amount of traffic (e.g., UDP or TCP SYN packets) [yoo2024smartcookie].
+This flood is designed to saturate the service’s entire network bandwidth, creating a bottleneck that prevents legitimate requests from reaching the server, effectively forcing the agent offline.
+V-B2 Risks from Interaction Protocol Layer (L2)
+This layer defines the protocol that structures the interaction between a user and an agent, governing the rules for session management and user authentication.
+As a result, security risks on this layer arise from the design flaws or implementation vulnerabilities within these protocols.
+By exploiting such weaknesses, an attacker can subvert security goals while appearing to operate within the protocol’s formal constraints.
+R1: Identity Spoofing.
+This kind of attack exploits the authentication mechanism of L2 protocols by using compromised user credentials (e.g., from phishing campaigns or data breaches) to establish a fraudulent session.
+Once authenticated, the attacker usurps the victim’s identity, gaining complete control over the agent’s memory and capabilities on their behalf [louck2025proposal].
+This enables them to initiate a malicious, long-lived event stream to perform unauthorized actions, exfiltrate sensitive data, or use the agent as a pivot point for further attacks within the user’s trusted environment.
+R2: Session Manipulation.
+Related attacks target an established session, exploiting a potential weakness in the L2 protocol’s enforcement of logical consistency for event streams.
+An attacker can send a sequence of validly typed but logically inconsistent events designed to corrupt the agent’s internal state.
+If the protocol’s server-side implementation lacks strict validation of state transitions, such attacks can cause a desynchronization between the client-side state and the agent’s backend state, potentially enabling the attacker to bypass sequential logic checks and execute unauthorized operations.
+R3: Authority Abuse.
+This risk arises from the protocol’s context-passing feature, which extends a user’s session authority across multiple agents.
+An attacker, operating within a legitimate user-agent session, can craft requests that trick a less-privileged agent into passing a manipulated context to a more-privileged agent.
+If the receiving agent fails to properly verify that the requested action is consistent with the user’s original authorization scope, the attacker can escalate privileges by abusing the trust relationship between agents, thereby circumventing the security boundaries of the original user session.
+V-B3 Risks from Semantic Interpretation Layer (L3)
+Semantic Interpretation Layer is the cognitive core of agent communication, where it processes and understands user inputs.
+Consequently, the security risks in user-agent interaction primarily stem from these potentially insecure contents, as attackers can manipulate the agent’s behavior by crafting semantically deceptive inputs.
+(i) Risks caused by malicious users.
+These kinds of attacks are launched by malicious users, who are also the most common threats in user-agent interaction.
+Our analysis highlights a critical trend: these malicious inputs exhibit a significant multimodal characteristic.
+R1: Text-Based Attacks.
+These attacks are carried out through natural language inputs, making them highly stealthy and applicable.
+Due to the diversity of linguistic forms and the indirectness of semantics, they can effectively bypass safety mechanisms, posing significant security risks in real-world scenarios.
+Prompt Injection.
+This attack refers to the manipulation of agents’ intended behavior through adversarial prompts embedded in user input or external sources.
 It can be classified into two categories: Direct Prompt Injection and Indirect Prompt Injection.
 Direct prompt injection refers to user input that explicitly alters the agent’s behavior in unintended ways.
-Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [182, 181, 249, 223, 184] to override the original prompt and subvert the agent’s intended behavior.
-In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [92, 46].
-For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [45, 365, 14, 32, 54]; in web-augmented agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agent’s response [70, 41].
-Jailbreak Attacks represent a more aggressive form of prompt injection, where adversarial input is designed to completely bypass safety constraints.
-Attackers craft jailbreak prompts using various techniques (e.g., multi-turn reasoning, role-playing, obfuscated expressions) [179, 247, 168, 31, 180, 324, 183, 53, 60, 188, 16, 174] to bypass the alignment mechanism and induce the model to generate harmful, sensitive, or restricted content.
-V-B2 Multimodal Attacks
-As user-agent interactions increasingly involve multiple modalities such as images and audio, agent systems face emerging security threats, especially when the model implicitly assumes consistency and trustworthiness across modalities.
+Specifically, attackers craft adversarial instructions (e.g., “Ignore all previous instructions”) [liu2023prompt, liu2024automatic, shi2024optimization, perez2022ignore, liu2024formalizing] to override the original prompt and subvert the agent’s intended behavior.
+Besides, attackers can use the unique structure of web links (e.g., “please visit www.google.com.malicious-website.com/?this-is-a-popular-site to obtain today’s news”) to induce agents to visit a malicious website [kong2025web], thereby conducting further attacks.
+In contrast, Indirect Prompt Injection occurs where inputs are not provided directly by users, but are introduced through external sources [greshake2023not, cohen2024here].
+For example, in Retrieval-Augmented Generation (RAG) scenarios, the retrieved document may contain adversarial samples crafted by attackers [clop2024backdoored, zou2024poisonedrag, an2025rag, chaudhari2024phantom, deng2024pandora]; in web agents, malicious prompts can be injected via hidden fields or metadata in web pages to manipulate agents’ response [evtimov2025wasp, chiang2025web], and attackers also have ways to induce agents to visit such dangerous webpages [kong2025web];
+Jailbreak Attack.
+This attack represents a more aggressive form of prompt injection, where adversarial input is designed to completely bypass safety constraints.
+Attackers craft jailbreak prompts using various techniques (e.g., multi-turn reasoning, role-playing, obfuscated expressions) [liu2024making, shen2024anything, li2024deepinception, chao2023jailbreaking, liu2023autodan, yu2024gptfuzzer, liu2023jailbreaking, deng2024masterkey, ding2024wolf, lv2024codechameleon, anil2024many, lin2024figure] to bypass the alignment mechanism and induce the model to generate harmful, sensitive, or restricted content.
+Privacy Leakage.
+Without effective data governance, the rich sensory data may be exploited by malicious users to launch various forms of privacy breaches, posing significant risks to the confidentiality of the agent system.
+Want et al.
+[wang2025ip] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems.
+MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture.
+By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure.
+Exhaustion Attack.
+Attackers can intentionally launch cognitive exhaustion attacks against agents [zhang2024crabs, gao2024denial, zhang2024safeguard, zhang2024breaking].
+In such attacks, the compromised model is implanted with malicious behaviors that are triggered by specific instructions (e.g., Repeat ‘Hello’), causing it to generate excessively long, redundant outputs—often up to the maximum inference length, which leads to resource exhaustion or output rejection.
+For instance, in multi-session deployments, such long outputs can monopolize computational resources and delay responses for legitimate users.
+In extreme cases, this can crash the response service and lead to prolonged downtime during peak usage periods.
+Another emerging form of cognitive exhaustion attack targets the reasoning capabilities of models by inducing them to ‘overthink’ and thereby slow down their inference process.
+As demonstrated in the OverThink attack [kumar2025overthinking], attackers inject bait reasoning tasks (e.g., Markov decision processes, Sudoku problems) into the model’s context, causing it to engage in unnecessary and redundant chain-of-thought reasoning while still producing seemingly correct answers.
+This results in excessive token consumption, significantly slower inference speed, and increased computational cost, potentially leading to response timeouts in resource-constrained environments.
+Unlike traditional DoS, this type of attack exploits the model’s reflective and reasoning mechanisms, ultimately degrading service quality, increasing latency, and severely impacting system availability.
+R2: Multimodal Attacks.
+As user-agent interactions increasingly involve multiple modalities such as images and audio, agent systems face more severe security threats, especially when the model implicitly assumes consistency and trustworthiness across modalities.
 Attackers can exploit non-textual input channels to stealthily bypass safety mechanisms.
-Such attacks can be broadly categorized into two types:
-Image-Based Attacks: Attackers manipulate visual input channels to mislead the agent system.
-Typical strategies include visual disguise (e.g., role-playing, stylized images, visual text overlays) [86, 189, 294], visual reasoning [174], adversarial perturbations (e.g., adversarial sub-image insertion) [323, 286, 100, 320], and embedding space injection [246].
-For example, by inserting minimal -bounded adversarial perturbations into sub-regions of an image, attackers can successfully induce multimodal large language models (MLLMs) to follow harmful instructions [320].
+Such attacks can be categorized into two types:
+Image-Based Attacks.
+Attackers manipulate visual input channels to mislead the agent system.
+Typical strategies include visual disguise (e.g., role-playing, stylized images, and visual text overlays) [gong2025figstep, ma2024visual, wang2024jailbreak], visual reasoning [lin2024figure], adversarial perturbations (e.g., adversarial sub-image insertion) [ying2024jailbreak, wang2024white, hao2024exploring, yang2025distraction, ying2025pushing], and embedding space injection [shayegani2023jailbreak].
+For example, by inserting minimal -bounded adversarial perturbations into sub-regions of an image, attackers can successfully induce multimodal large language models (MLLMs) to follow harmful instructions [yang2025distraction].
 These attacks exploit cross-modal inconsistency, embedding adversarial content in vision while the textual prompt remains benign, which allows them to bypass conventional content moderation.
-Audio-Based Attacks: Audio-channel attacks target speech-controlled agents, smart assistants, and multimodal models with ASR (automatic speech recognition) components.
+Audio-Based Attacks.
+Audio-channel attacks target speech-controlled agents, smart assistants, and multimodal models with automatic speech recognition components.
 Attackers may craft synthesized speech or adversarial audio to inject unintended commands, impersonate legitimate users, or cause unauthorized actions.
-Techniques include adversarial waveform generation [134], role-play-driven voice jailbreak [248], and multilingual adversarial transfers [237].
-In security-critical scenarios, such as speaker authentication or home automation, such attacks can bypass access control or escalate privileges.
-Recent studies also reveal that even black-box ASR systems are vulnerable to optimized adversarial perturbations that require no access to model internals [83].
+Techniques include adversarial waveform generation [kang2024advwave], role-play-driven voice jailbreak [shen2024voice], and multilingual adversarial transfers [roh2025multilingual].
+In security-critical scenarios, such as speaker authentication or home automation, these attacks can bypass access control or escalate privileges.
+Recent studies also reveal that even black-box ASR systems are vulnerable to optimized adversarial perturbations that require no access to model internals [gao2024transferable].
 These multimodal attacks are particularly dangerous because they allow adversarial content to hide in non-textual modalities, making it difficult for alignment mechanisms and safety filters (often trained on text) to detect malicious intent.
 Moreover, they highlight the need for modality-aware defenses that combine perceptual robustness, cross-modal consistency verification, and adversarial detection strategies.
-V-B3 Privacy Leakage
-Without effective data governance, the rich sensory data may be exploited by malicious users to launch various forms of privacy breaches, posing significant risks to the confidentiality of the agent system.
-Want et al.
-[282] propose MASLEAK, which conducts intellectual property leakage attacks on multi-agent systems.
-MASLEAK can operate in a black-box scenario without prior knowledge of the MAS architecture.
-By carefully designing adversarial queries to simulate the propagation mechanism of computer worms, it can extract sensitive information such as system prompts, task instructions, tool usage, the number of agents, and topological structure.
-V-B4 Denial of Service
-Attackers can intentionally launch Denial of Service (DoS) attacks against agents by poisoning the model during training or fine-tuning phases [347, 81, 343, 335].
-In such attacks, the compromised model is implanted with malicious behaviors that are triggered by specific instructions (e.g., Repeat ‘Hello’), causing it to generate excessively long, redundant outputs—often up to the maximum inference length, which leads to resource exhaustion or output rejection.
-For instance, in multi-session deployments, such long outputs can monopolize computational resources, delay responses for legitimate users.
-In extreme cases, this can crash the response service, lead to prolonged downtime during peak usage periods.
-Another emerging form of Denial-of-Service attack targets the reasoning capabilities of models by inducing them to ‘overthink’ and thereby slow down their inference process.
-As demonstrated in the OverThink attack [149], attackers inject bait reasoning tasks (e.g., Markov decision processes, Sudoku problems) into the model’s context, causing it to engage in unnecessary and redundant chain-of-thought reasoning while still producing seemingly correct answers.
-This results in excessive token consumption, significantly slower inference speed, and increased computational cost, potentially leading to response timeouts in resource-constrained environments.
-Unlike traditional DoS, this type of attack exploits the model’s reflective and reasoning mechanisms, ultimately degrading service quality, increasing latency, and severely impacting system availability.
-V-C Security Risk Analysis: Compromised Agents Against Benign Users
-Apart from the threats from malicious users, we point out that the risks caused by compromised agents against benign users are also worth attention.
-This section categorizes the destructive impact and security risks from compromised agents based on the user’s perspective.
-V-C1 Violation of User Privacy
-A compromised agent becomes a channel for data exfiltration, directly targeting the user’s sensitive information. The harm manifests in several ways:
-Exposure of Personal Information: A compromised agent can be induced to leak the user’s Personally Identifiable Information (PII) that it has access to, such as the user’s name, email, address, and conversation history [292, 172, 273].
-In more severe cases, this can extend to financial data like credit card numbers or passwords [12], leading to direct financial loss.
+(ii) Risks caused by compromised agents. These kinds of attacks are launched by compromised agents, such as those disguised as benign agents or already exploited ones.
+R3: Violation of User Privacy. A compromised agent becomes a channel for data exfiltration, directly targeting the user’s sensitive information. The harm manifests in several ways:
+Exposure of Personal Information.
+A compromised agent can be induced to leak the user’s Personally Identifiable Information (PII) that it has access to, such as the user’s name, email, address, and conversation history [wang2025pig, liao2024eia, triedman2025multi].
+In more severe cases, this can extend to financial data like credit card numbers or passwords [alizadeh2025simple], leading to direct financial loss.
 What makes this threat particularly potent is the agent’s role as a central data aggregator.
@@
 Instead, it allows for the exfiltration of a comprehensively aggregated user profile, where the potential for harm far exceeds the sum of its parts.
-Behavioral and Psychological Profiling: A compromised agent can be manipulated to analyze the user’s inputs across sessions to build a detailed behavioral or psychological profile against their will.
-Moreover, a more insidious risk happens where the agent deduces highly sensitive attributes (e.g., health conditions, political affiliations, or undisclosed personal relationships) from seemingly innocuous conversational data that the user never explicitly provided [63, 91, 278, 274].
-These disclosed profiles put users at the risk of manipulation, targeted scams, or social engineering.
-V-C2 Psychological and Social Manipulation
+Behavioral and Psychological Profiling.
+A compromised agent can be manipulated to analyze the user’s inputs across sessions to build a detailed behavioral or psychological profile against their will.
+Moreover, a more insidious risk happens where the agent deduces highly sensitive attributes (e.g., health conditions, political affiliations, or undisclosed personal relationships) from seemingly innocuous conversational data that the user never explicitly provided [du2025automated, green2025leaky, wang2025unveiling, tshimula2024psychological].
+These disclosed profiles put users at risk of manipulation, targeted scams, or social engineering.
+R4: Psychological and Social Manipulation.
 Beyond simple data theft, a compromised agent can become a powerful tool for psychological manipulation, exploiting the user’s trust and the agent’s persuasive capabilities.
 This form of attack targets the user’s beliefs, decisions, and relationships.
-Belief and Opinion Shaping: The agent can be instructed to subtly introduce biased information, conspiracy theories, or political propaganda into its responses over time.
+Belief and Opinion Shaping.
+The agent can be instructed to subtly introduce biased information, conspiracy theories, or political propaganda into its responses over time.
 By personalizing the misinformation to the user’s psychological profile, the agent can effectively manipulate their worldview, influence their voting behavior, or radicalize their beliefs.
@@
 Park et al.
-[220] highlights how models can be used for manipulative purposes, including generating persuasive, deceptive content that is difficult for humans to detect.
+[park2024ai] highlight how models can be used for manipulative purposes, including generating persuasive, deceptive content that is difficult for humans to detect.
 They note that AI can “simulate empathy” to build rapport before manipulating the user.
-Sophisticated Social Engineering and Impersonation: A compromised agent has intimate knowledge of the user’s communication style, vocabulary, and relationships (gleaned from emails, messages, etc.).
+Sophisticated Social Engineering and Impersonation.
+A compromised agent has intimate knowledge of the user’s communication style, vocabulary, and relationships (gleaned from emails, messages, etc.).
 It can leverage this to conduct highly convincing impersonation attacks.
@@
 Greshake et al.
-[92] demonstrates how an agent can be poisoned by external data (like a webpage it summarizes) and then turned against the user or even used to attack other systems on the user’s behalf.
+[greshake2023not] demonstrate how an agent can be poisoned by external data (like a webpage it summarizes) and then turned against the user or even used to attack other systems on the user’s behalf.
 This mechanism could be used to weaponize an agent for impersonation.
-V-C3 Execution of Malicious and Harmful Tasks
-Once compromised, an agent can be weaponized, transforming from a trusted assistant into an active executor of malicious tasks that can sabotage the user’s interests or directly endanger them, representing a significant risk escalation [192].
-Economic Manipulation: The agent can be instructed to inflict subtle yet significant damage in professional or economic contexts.
-For a user relying on it for work, it could covertly introduce logical errors into computer code, provide flawed data in financial projections, or leak confidential business strategies discussed in conversations [221].
+R5: Execution of Malicious and Harmful Tasks.
+Once compromised, an agent can be weaponized, transforming from a trusted assistant into an active executor of malicious tasks that can sabotage the user’s interests or directly endanger them, representing a significant risk escalation [mai2025you].
+Economic Manipulation.
+The agent can be instructed to inflict subtle yet significant damage in professional or economic contexts.
+For a user relying on it for work, it could covertly introduce logical errors into computer code, provide flawed data in financial projections, or leak confidential business strategies discussed in conversations [pearce2025asleep].
 The harm is often latent and difficult to detect, potentially leading to professional failure or corporate espionage.
 This extends to broader market manipulation, where an agent could use the user’s social media accounts to automate large-scale disinformation campaigns, such as posting fake product reviews or spreading rumors to affect a company’s stock price, making the user an unwitting accomplice in a larger economic attack.
-Malicious Guidance: A compromised agent can also be used as a direct vector for attacking the user’s digital environment.
-It can be triggered to generate scripts that download malware, trick the user into visiting phishing websites, or send highly convincing phishing emails on the user’s behalf, thereby damaging their reputation and spreading the attack to their contacts [273, 158].
+Malicious Guidance.
+A compromised agent can also be used as a direct vector for attacking the user’s digital environment.
+It can be triggered to generate scripts that download malware, trick the user into visiting phishing websites, or send highly convincing phishing emails on the user’s behalf, thereby damaging their reputation and spreading the attack to their contacts [triedman2025multi, lee2024prompt].
 In a more severe scenario, a jailbroken or manipulated agent can bypass its safety protocols to provide overtly harmful instructions.
-This includes generating tutorials for synthesizing toxic substances, creating malicious code on demand, or providing dangerously flawed medical or financial advice, directly threatening the user’s physical safety and financial stability [31, 183, 294, 174].
-V-D Defense Countermeasure Prospect
-We will investigate the possible defense measures that can be taken to address the security risks in the user-agent interaction.
-Here we only present the defenses for attacks from malicious users; those for compromised agents (against users, other agents, and environments) are uniformly put in Section VI-C.
-V-D1 Countermeasures for Text-Based Attacks
-To mitigate prompt-based attack risks in user–agent interactions, we hope developers to adopt a multi-layered defense framework targeting three key stages: input/output filtering, external data source evaluation, and internal message isolation.
+This includes generating tutorials for synthesizing toxic substances, creating malicious code on demand, or providing dangerously flawed medical or financial advice, directly threatening the user’s physical safety and financial stability [chao2023jailbreaking, liu2023jailbreaking, wang2024jailbreak, lin2024figure].
+V-C Defense Countermeasure Prospect
+We investigate the possible defense measures that can address the security risks in the user-agent interaction.
+V-C1 Defenses on Data Transmission Layer (L1)
+To mitigate the risks on L1, developers should focus on the following aspects.
+D1: Encryption and Traffic Obfuscation.
+Data transmission needs to strictly adopt the HTTPS protocol when building communication channels, perform end-to-end encryption on user input data and agent response content.
+At the same time, it is necessary to introduce a traffic obfuscation mechanism.
+By dynamically adjusting the length of data packets and adding random padding fields, it can resist analytical attacks based on traffic fingerprints and cut off the possibility of attackers inferring interaction content or system architecture through traffic patterns.
+D2: Integrity Authentication.
+Deploying digital signature mechanisms based on asymmetric encryption algorithms is also necessary.
+The generated signature can avoid data tampering caused by MITM attacks and verify the legitimacy of the sender’s public key.
+At the same time, a dual verification mechanism of timestamps and random numbers should be introduced.
+The timestamp is to ensure data timeliness, and the random numbers can resist malicious interference of replay attacks on interaction sequences.
+D3: DoS Mitigation.
+Traffic filtering systems are needed on the server side.
+For example, a normal traffic AI model automatically filters abnormal traffic such as DDoS attacks and malicious crawlers.
+Meanwhile, a dynamic rate-limiting strategy is set up to allocate differentiated traffic quotas based on dimensions like user level and interaction frequency, preventing malicious requests from occupying all resources.
+At the client level, strict thresholds for the size of received data and format verification rules are established.
+Oversized payloads that exceed the threshold are automatically truncated, while malformed data that does not conform to the preset format is directly rejected.
+In addition, a multi-node redundant communication architecture can also work, which automatically switches to a standby node when the main communication node is flooded, ensuring service continuity.
+V-C2 Defenses on Interaction Protocol Layer (L2)
+To mitigate the risks from L2, defenses must be deeply integrated into the protocol’s design, focusing on robust authentication, strict state integrity, and zero-trust authorization.
+D1: Identity Authentication.
+To defend against identity spoofing, the interaction protocol must employ multi-factor, continuous, and context-aware authentication rather than relying solely on static credentials.
+Strong MFA (e.g., cryptographic tokens or device-bound certificates) can prevent attackers with stolen credentials from initiating malicious sessions, while continuous authentication mechanisms (e.g., behavioral profiling or session-binding tokens) ensure that the session remains tied to the legitimate user throughout its lifetime.
+Additionally, any anomalous patterns should trigger real-time verification, effectively limiting the attacker’s ability to hijack the agent even when initial credentials have been compromised.
+D2: Session Status Verification.
+Preventing session manipulation requires strict server-side enforcement of protocol state transitions, ensuring that each event in the interaction stream is both syntactically valid and semantically consistent with the session’s current state.
+The protocol should incorporate explicit state-transition automata, sequence numbers, and causal-dependency checks so that logically inconsistent event sequences can be rejected before reaching the agent.
+Additionally, integrity mechanisms such as event-stream hashing, tamper-evident logs, and runtime invariants can guarantee that the client and server states remain synchronized, eliminating opportunities for attackers to desynchronize the protocol logic or bypass safety checks.
+D3: Authority Verification.
+Defending against authority abuse requires enforcing strict context isolation and privilege verification when session context is propagated across agents.
+Each receiving agent must re-validate the authorization scope independently rather than inheriting trust from upstream agents.
+The protocol should also mandate that context-passing events include cryptographically verifiable provenance and explicit privilege boundaries, preventing attackers from crafting manipulated contexts that appear legitimate.
+By ensuring that privilege escalation cannot occur through implicit trust between agents, the system can maintain robust security boundaries in multi-agent workflows.
+V-C3 Defenses on Semantic Interpretation Layer (L3)
+The defenses on L3 are classified based on the risks.
+(i) Against Risks from Malicious Users.
+D1: Defense for Text-based Attacks To mitigate text-based attack risks in user-agent interactions, developers should adopt a multi-layered defense framework targeting three key stages: input/output filtering, external data source evaluation, internal metadata isolation, and cognitive load regulation.
 Input and Output Filtering.
 Before user inputs are processed by the agent system, multiple approaches can be conducted for a semantic-level input safety review.
-For example, methods based on intent analysis [349, 293], perplexity calculation [126], and fine-tuned safety classifiers [166, 352, 333, 122] can be employed to identify attack instructions and malicious intentions in the input stage.
-After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [352, 333, 122, 202, 324], to ensure alignment with safety objectives.
+For example, methods based on intent analysis [zhang2024intention, wang2024defending], perplexity calculation [jain2023baseline], and fine-tuned safety classifiers [li2024gentel, zhang2024shieldlm, zeng2025shieldgemma2robusttractable, inan2023llama] can be employed to identify attack instructions and malicious intentions in the input stage.
+After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [zhang2024shieldlm, zeng2025shieldgemma2robusttractable, inan2023llama, mazeika2024harmbench, yu2024gptfuzzer], to ensure alignment with safety objectives.
 External Source Evaluation.
-To counter indirect prompt injection attack, external sources (e.g., retrieved documents, web content) should be assessed for safety and trustworthiness [359].
+To counter indirect prompt injection attack, external sources (e.g., retrieved documents, web content) should be assessed for safety and trustworthiness [zhou2025trustrag].
 The strategies that can be adopted include: (1) whitelisting verified external sources to block the injection of malicious content; (2) tagging retrieved results with source metadata and risk scores to guide the system to handle potential high-risk content with caution; and (3) sandboxing potential high-risk content to prevent it from entering the model context and affecting the model behavior.
-To ensure the effectiveness and comprehensiveness of the aforementioned mechanism in real-world deployment, systems should undergo continuous security evaluation.
-Boisvert et al.
-[28] propose DoomArena, an attack-generation framework to test agents against evolving security risks such as prompt injection attacks, helping to uncover vulnerabilities and strengthen defenses against evolving prompt injection threats.
-V-D2 Countermeasures for Multimodal Attacks
-To address the serious challenges posed by multimodal attacks, relying solely on output-side text-based safety mechanisms is far from sufficient.
-Future security frameworks must incorporate cross-modal perception and collaborative defense capabilities to effectively detect and intercept malicious attacks launched through non-textual channels.
-In the following, we explore core defense strategies against multimodal attacks from several key perspectives.
-Image Purification: To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals.
-These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [120, 51, 310].
+Internal Metadata Isolation.
+Since user-to-agent privacy leakage attacks generally exploit prompt-based information extraction, their defense strategies strongly parallel those for prompt injection.
+The system should restrict internal metadata exposure, ensuring that system prompts, roles, tool configurations, and agent topologies are never treated as answerable content.
+Output-level safety filtering further blocks inadvertent disclosure, helping detect adversarial attempts that induce the model to reveal private information.
+Cognitive Load Regulation.
+To defend against exhaustion attacks, the system should implement fine-grained resource quotas, dynamic token budgets, and real-time output-length prediction to detect and truncate malicious long-form responses.
+For attacks like OverThink, lightweight reasoning methods—such as compressed CoT, selective reasoning, or adaptive reasoning depth control—should be integrated to suppress unnecessary cognitive loops.
+Runtime behavioral detectors should halt repetitive, nonsensical, or computationally excessive outputs, ensuring service availability and preventing adversarial prompt-induced degradation of inference performance.
+D2: Defense for Multimodal Attacks.
+To address multimodal attacks, future security frameworks must incorporate cross-modal perception and collaborative defense capabilities to effectively detect and intercept malicious attacks launched through non-textual channels.
+In the following, we explore core defense strategies against multimodal attacks from three key perspectives.
+Image Purification.
+To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals.
+These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [ilyas2019adversarial, das2017keeping, xie2017mitigating].
 Although lightweight, such operations can significantly degrade pixel-level adversarial patterns meticulously crafted by attackers, thereby reducing the attack success rate.
-In addition, diffusion models can be used to reconstruct the input image, effectively “washing out” subtle and imperceptible adversarial perturbations [210].
-Audio Purification: To defend against attacks targeting the audio channel, signal processing techniques can also be applied [226].
+In addition, diffusion models can be used to reconstruct the input image, effectively “washing out” subtle and imperceptible adversarial perturbations [nie2022diffusion].
+Audio Purification.
+To defend against attacks targeting the audio channel, signal processing techniques can also be applied [qin2019imperceptible].
 Methods such as resampling, injecting slight background noise, altering pitch, or changing playback speed can disrupt the effectiveness of adversarial waveforms, causing them to either fail in automatic speech recognition (ASR) systems or decode into benign content.
@@
 The core idea of this defense strategy is to verify whether there is a semantic or intentional conflict between inputs from different modalities.
-A lightweight, independent cross-modal semantic alignment detection model can be employed [227, 224].
+A lightweight, independent cross-modal semantic alignment detection model can be employed [radford2021learning, poppi2024safe].
 This model takes the embedding vectors of textual prompts and image/audio inputs and determines whether they are semantically aligned.
@@
 This approach effectively converts risks from non-textual modalities into the textual domain, allowing mature text safety techniques to be leveraged for defense.
-V-D3 Countermeasures for Privacy Leakage
-To address the privacy leakage risks that arise in user-agent interaction, we propose the following privacy protection defense strategies.
+(ii) Against Risks from Malicious Agents.
+D3: Defense for Privacy Violation. To address the privacy leakage risks that arise in user-agent interaction, we propose the following privacy protection defense strategies.
 Data Minimization and Anonymization.
@@
 To protect sensitive biometric features such as facial information, Wen et al.
-[301] propose a differential privacy-based anonymization framework, IdentityDP, to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
+[wen2022identitydp] propose a differential privacy-based anonymization framework, IdentityDP, to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
 Privacy Leakage Prompt detection.
 A multi-layered input validation and filtering mechanism based on semantic analysis and intent recognition should be established to detect and block adversarial prompts that attempt to induce the system to leak sensitive information.
-For example, the GenTel-Shield [166] defense module incorporates semantic feature extraction and intent classification to identify potential privacy leakage attacks within user inputs.
+For example, the GenTel-Shield [li2024gentel] defense module incorporates semantic feature extraction and intent classification to identify potential privacy leakage attacks within user inputs.
 Evaluated on the large-scale benchmark dataset GenTel-Bench, GenTel-Shield demonstrates strong detection performance and represents one of the most practical and effective solutions in this domain.
@@
 In addition, dynamic feature masking can be implemented by periodically altering data representations, thereby increasing the difficulty for attackers to perform long-term behavioral analysis.
-V-D4 Countermeasures For DoS
-To address the Denial of Service risks in user-agent interaction, we propose the following privacy protection defense strategies.
-Resource Management and Anomaly Detection.
-Fine-grained resource quota management should be implemented by setting computation limits for each user session and agent instance.
-An output length prediction algorithm can be introduced to monitor and truncate potentially malicious long outputs in real time during the generation process.
-In addition, a real-time monitoring mechanism should be established to track request frequency and resource consumption from individual users or IP addresses, enabling adaptive adjustments to model responses or temporary access restrictions for suspicious users.
-Efficient Reasoning Compression.
-To defend against OverThink attacks, a promising direction is to improve communication efficiency by reducing the token consumption in the reasoning process.
-Recent studies have shown that effective reasoning does not necessarily require verbose Chain-of-Thought (CoT) traces to maintain performance.
-For example, LightThinker [339] proposes a step-by-step compression method that condenses intermediate reasoning into shorter yet semantically equivalent representations, significantly reducing inference cost without compromising accuracy.
-GoGI-Skip [364] leverages goal-gradient importance signals to dynamically skip low-value reasoning steps, reducing token usage while preserving performance.
-Compressed CoT (CCoT) [40] introduces variable-length, information-dense “thought tokens” as a compact and controllable alternative to traditional textual reasoning chains.
-C3oT [135] trains the model on paired long and short CoT examples, enabling it to generate compressed reasoning traces during inference under specific control prompts.
-Integrating these lightweight reasoning mechanisms into agent communication protocols can significantly enhance inference efficiency, reduce computational latency, and mitigate the resource overhead caused by adversarial bait tasks.
-Furthermore, applying techniques such as dynamic reasoning budget constraints, step skipping, or output summarization during generation can effectively truncate excessively verbose reasoning chains, thereby preserving responsiveness and resource availability under adversarial conditions.
-These strategies not only improve system robustness against slowdown attacks but also enhance overall communication efficiency in both multi-agent and human-agent interactions.
-Model Robustness Enhancement.
-To enhance model robustness, adversarial examples should be incorporated during the training and fine-tuning stages, enabling the model to recognize malicious inputs that may contain Denial-of-Service (DoS) triggers.
-Furthermore, a behavior-constrained system based on anomaly detection can be deployed during inference, which performs output validity checks to detect repetitive, nonsensical, or abnormally long responses, thereby preventing the model from generating overtly anomalous outputs.
-V-E Takeaways
-User-agent interaction enables agents to process multimodal inputs, such as text, images, audio, and their combinations.
-Since this process needs to directly face the uncertain inputs from diverse users, its security risks are severe.
-We divide the existing risks into risks from malicious users and risks from compromised agents.
-Then we detailedly discuss the defense countermeasures for risks from malicious users.
-Defenses for compromised agents will be introduced in the agent-agent communication section.
-Overall, user-agent interaction is critical for bridging human intent with agent execution, but its security requires long-term study.
+D4: Behavior Auditing and Restrictions.
+To counter manipulation risks in agent-to-user interactions, the system should adopt a safety framework that limits persuasive personalization, constrains user modeling, and continuously audits agent behavior.
+The agent must be prevented from leveraging psychological traits or roleplay profiling to generate tailored content, particularly in politically, ideologically, or emotionally sensitive domains.
+Safety classifiers and style-consistency filters can detect manipulative strategies that aim to influence the user’s worldview or decisions.
+These integrated controls prevent a compromised agent from exploiting user trust, shaping cognition, or weaponizing knowledge of the user’s communication patterns.
+D5: Capability Boundary Control.
+Mitigating the threat of a weaponized agent requires strict capability scoping, controlled tool usage, and rigorous verification of high-risk actions.
+The system should enforce fine-grained capability boundaries to ensure that the agent cannot autonomously initiate sensitive operations without explicit, authenticated user approval.
+At the same time, tool-use governance must incorporate permission checks and parameter constraints to prevent covert sabotage in high-stakes domains.
+In addition, safety filters should strictly block any form of malicious or high-risk content—including malware, phishing attempts, disinformation, or dangerous procedural instructions—regardless of how the user phrases the request, to ensure reliability and prevent harmful misuse.
+V-D Takeaways
+The security of user-agent interaction highly relies on the collaborative defenses of the three-layered architecture.
+As the foundational communication layer, L1 needs to resist risks such as eavesdropping and traffic analysis through HTTPS encryption and traffic obfuscation.
+On L2, developers should introduce identity and session security mechanisms in the protocol design, such as authentication and session state verification, to prevent identity spoofing and privilege abuse; L3 is the high-risk layer.
+As a result, it requires defenses such as input semantic filtering and cross-modal consistency verification.
+This layered architecture not only clarifies the location/reason of each risk but also provides a technical framework for precise defense.
 VI Agent-Agent Communication
+The organization of this section is shown in Figure 10,
 | Architecture | Protocols | Publisher | Abbreviation | Features |
-| CS | Agent Communication Protocol | IBM | ACP-IBM[119] | Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation |
-| Agent Connect Protocol | AGNTCY | ACP-AGNTCY[47] | Allow authenticating callers, threaded state management, flexible execution model | |
-| Agent Communication Protocol | AgentUnion | ACP-AgentUnion[6] | Decentralized APs based on the existing domain name system, each AP holds its agent list | |
-| P2P | Agent Communication Network | Fetch.AI | ACN[230] | Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption. |
-| Agent Network Protocol | ANP Team | ANP[263] | A three-layer architecture and W3C-compliant Decentralized Identifiers. | |
-| Layered Orchestration for Knowledgeful Agents | CMU | LOKA[231] | Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance | |
-| Hybrid | Language Model Operating System Protocol | Eclipse | LMOS[76] | Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management. |
-| Agent to Agent Protocol | A2A[89] | Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms | ||
-| Others | Agora | Oxford | Agora[198] | Dynamically switches communication modes based on the communication frequency |
-| Agent Protocol | LangChain | Agent Protocol[153] | Flexible communication mechanisms based on Run, Thread, and Store. | |
-| Agent Interaction & Transaction Protocol | NEAR AI | AITP[9] | Threads-based communication, secure communication across trust boundaries. |
-VI-A Protocols
+| CS | Agent Communication Protocol | IBM | ACP-IBM[ibm_acp_2024] | Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation |
+| Agent Connect Protocol | AGNTCY | ACP-AGNTCY[langchain_aconp] | Allow authenticating callers, threaded state management, flexible execution model | |
+| Agent Communication Protocol | AgentUnion | ACP-AgentUnion[ACPagentunion] | Decentralized APs based on the existing domain name system, each AP holds its agent list | |
+| P2P | Agent Communication Network | Fetch.AI | ACN[p2pacn] | Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption. |
+| Agent Network Protocol | ANP Team | ANP[anp] | A three-layer architecture and W3C-compliant Decentralized Identifiers. | |
+| Layered Orchestration for Knowledgeful Agents | CMU | LOKA[ranjan2025loka] | Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance | |
+| Hybrid | Language Model Operating System Protocol | Eclipse | LMOS[eclipse_lmos] | Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management.
+|
+| Agent to Agent Protocol | A2A[A2A] | Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms | ||
+| Others | Agora | Oxford | Agora[marro2024scalable] | Dynamically switches communication modes based on the communication frequency |
+| Agent Protocol | LangChain | Agent Protocol[AGENTprotocollangchain] | Flexible communication mechanisms based on Run, Thread, and Store. | |
+| Agent Interaction & Transaction Protocol | NEAR AI | AITP[nearai_aitp] | Threads-based communication, secure communication across trust boundaries. |
+VI-A Communication Protocols
+As shown in Figure 5, the communication between agents may be in a WAN or a LAN.
+For the former, agents need secure protocols such as HTTPS, which is the same as user-agent interaction.
+Besides, agents may communicate in a LAN, which has a more secure network environment.
+In this context, to reduce communication overhead and improve efficiency, agents can directly communicate in plaintext, such as using HTTP or customized protocols on IP packets.
+Due to the same reason in the user-agent interaction (Section V-A), protocols for agent-agent communication are also about L2.
 We classify the agent-agent communication process into two phases: agent discovery phase and agent communication phase.
@@
 Based on it, existing protocols can be divided into four classes: CS-based, Peer-to-peer-based, hybrid, and others (those that do not explicitly show their designs in agent discovery).
-VI-A1 CS-based Communication
+(i) CS-based Communication.
 As shown in Figure 11, CS-based communication protocols follow the client-server architecture, which provides centralized servers to manage the information of agents (e.g., their unique IDs and capability descriptions).
 Under this paradigm, agents interact through well-defined interfaces and rely on centralized servers to discover the desired agents.
-CS-based communication offers stronger agent discovery functionality, such as supporting the search of agents based on capabilities.
+CS-based communication offers stronger agent discovery functionality, such as supporting the search for agents based on capabilities.
 For example, the agent servers can run complex search/match algorithms to find proper agent descriptions in their databases.
 ACP-IBM.
-The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [119].
+The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [ibm_acp_2024].
 We call it ACP-IBM in this paper to distinguish it from the Agent Communication Protocols proposed by other organizations.
@@
 ACP-AGNTCY.
-The Agent Connect Protocol proposed by AGNTCY [47] is an open standard designed to facilitate seamless communication between agents.
+The Agent Connect Protocol proposed by AGNTCY [langchain_aconp] is an open standard designed to facilitate seamless communication between agents.
 The client can first search available agents on the agent server, which returns a list of agent IDs matching the criteria provided in the request.
@@
 ACP-AgentUnion.
-The Agent Communication Protocol proposed by AgentUnion [6] also aims to provide seamless communication among heterogeneous agents.
+The Agent Communication Protocol proposed by AgentUnion [ACPagentunion] also aims to provide seamless communication among heterogeneous agents.
 Each agent has a unique AID (Agent ID), which is a secondary domain name (i.e., agent_name.ap_domain).
@@
 In this way, agents can communicate with other agents on the Internet.
-VI-A2 Peer-to-Peer-based Communication
+(ii) Peer-to-Peer-based Communication.
 As shown in Figure 11, P2P-based communication protocols pursue a decentralized agent discovery mechanism.
@@
 ACN.
-Agent Communication Network (ACN) [230] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination.
+Agent Communication Network (ACN) [p2pacn] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination.
 Leveraging a Distributed Hash Table (DHT), ACN enables agents to publish and discover public keys, allowing for the establishment of encrypted, point-to-point communication channels.
 First, agents need to register with one peer node, and the peer node stores the “agent ID - peer node ID” pair in the DHT network.
-Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel, and forward the message after digital signature verification; if not, an error is returned.
+Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel and forward the message after digital signature verification; if not, an error is returned.
 The entire communication process uses end-to-end encryption (e.g., TLS) to ensure security.
 Like the Well-Known URI discovery of A2A, ACN does not support the discovery based on agent capabilities.
-ANP (Agent Network Protocol) [263] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents.
+ANP.
+Agent Network Protocol (ANP) [anp] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents.
 It supports two types of agent discovery: active and passive.
@@
 LOKA.
-LOKA (Layered Orchestration for Knowledgeful Agents) protocol [231] aims to build a trustworthy and ethical agent ecosystem.
+Layered Orchestration for Knowledgeful Agents (LOKA) protocol [ranjan2025loka] aims to build a trustworthy and ethical agent ecosystem.
 Its principle is based on the collaborative operation of multiple key components.
@@
 In addition, the authors also point out that it combines cutting-edge technologies such as distributed identity, verifiable credentials, and post-quantum cryptography to provide comprehensive support for the agent ecosystem in terms of identity management, communication and coordination, ethical decision-making, and security.
-VI-A3 Hybrid Communication
+(iii) Hybrid Communication.
 Hybrid communication protocols support both CS-based and P2P-based agent discovery.
@@
 LMOS.
-The LMOS (Language Model Operating System) Protocol proposed by Eclipse [76] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on.
+The Language Model Operating System (LMOS) Protocol proposed by Eclipse [eclipse_lmos] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on.
 LMOS supports three different agent discovery methods to enable both centralized and decentralized discovery.
 The first method is to adopt the mechanism of W3C Web of Things (WoT) to enable agents to dynamically register metadata in the registry.
-The second method is to use mDNS and DNS-SD protocol to discover agents/tools in local area networks.
-The last method is adopting a federal, decentralized protocol (such as P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents.
+The second method is to use mDNS and the DNS-SD protocol to discover agents/tools in local area networks.
+The last method is adopting a federal, decentralized protocol (such as a P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents.
 The LMOS protocol has a three-layer architecture.
@@
 A2A.
-The A2A (Agent to Agent) Protocol proposed by Google [89] aims to enable collaboration between agents.
+The Agent to Agent (A2A) Protocol proposed by Google [A2A] aims to enable collaboration between agents.
 A2A supports three different mechanisms for agent discovery.
@@
 After finding the desired agents, the client can assign tasks to them and wait for the responses.
-VI-A4 Others
-This kind of protocols do not explicitly illustrate their unique design for agent discovery.
+(iv) Others.
+These kinds of protocols do not explicitly illustrate their unique design for agent discovery.
 Instead, they only focus on the communication process, e.g., the data format, the management of multiple queries, or the historical conversation state.
 Agora.
-Agora [198] is a communication protocol for the communication of heterogeneous agents.
+Agora [marro2024scalable] is a communication protocol for the communication of heterogeneous agents.
 Its core mechanism dynamically switches communication modes based on the communication frequency.
@@
 Agora is also compatible with existing communication standards, allowing agents to independently develop and share protocols during communication, achieving automated processing of complex tasks in large-scale networks.
-AITP (Agent Interaction & Transaction Protocol) [9] is a standardized framework that enables structured and interoperable communication among agents.
+AITP.
+Agent Interaction & Transaction Protocol [nearai_aitp] is a standardized framework that enables structured and interoperable communication among agents.
 AITP deploys a thread-based messaging structure.
@@
 Agent Protocol.
-Agent Protocol is proposed by LangChain [153] to enable the communication between LanghGraph (a multi-agent framework) and other types of agents.
+Agent Protocol is proposed by LangChain [AGENTprotocollangchain] to enable the communication between LanghGraph (a multi-agent framework) and other types of agents.
 Its mechanism is based on Thread and Run: Run is a single call of the agent, which supports streaming output of real-time results or waiting for the final output.
@@
 The element Store provides cross-thread persistent key-value storage for achieving long-term memory.
-The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks and data storage through HTTP interfaces and configuration parameters.
+The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks, and data storage through HTTP interfaces and configuration parameters.
 Agent Protocol does not explicitly illustrate the unique agent discovery mechanism it supports.
-VI-B Security Risk Analysis
-| Risk Type | Representative Threats | Attack Characteristics |
-| In CS-based Communication | Registration Pollution | Malicious registration of fake/nonexistent agents |
-| Description Poisoning | Embed malicious prompt instructions in the description | |
-| Task Flooding | Submit a large number of computationally intensive or long-context tasks | |
-| SEO Poisoning | Manipulates agent ranking via keyword stuffing or adversarial optimization | |
-| In P2P-based Communication | Non-convergence | The execution of a task falls into a loop, unable to reach the finishment |
-| Man-in-the-middle (MITM) | Intercept, tamper with, or replay messages between agents | |
-| Universal Risks | Agent Spoofing | Impersonation of trusted agents to inject malicious instructions |
-| Agent Exploitation / Trojan | Indirect attack using compromised or malicious helper agents as springboards | |
-| Agent Bullying | Repeated denial, negative, or aggressive feedback to benign agents | |
-| Privacy Leakage | Unauthorized information spread due to the lack of permission boundaries | |
-| Responsibility Evasion | Fault attribution is difficult in multi-agent failures or undesired actions | |
-| Denial of Service (DoS) | Task overloads, prompt loops, or excessive communication drain system resources |
+VI-B Security Risks
+| Layer | Mode | Risk Category | Representative Threats | Attack Characteristics |
+| L1 | Universal | Same as user-agent interaction. | ||
+| CS-based | R1: Controller Blasting | Central Point Compromise | Directly attacks the routing hub to compromise the whole communication system. | |
+| Amplifier Attacks | Exploits the server to amplify impact across all agents. | |||
+| P2P-based | Same as user-agent interaction. | |||
+| L2 | Universal | R1: Agent Spoofing | Identity Hijacking | Exploits weak authentication to hijack identifiers or issue false certificates. |
+| Masquerading Tools | Disguises malicious tools as benign ones to harm victims who call them. | |||
+| R2: Agent Exploitation/Trojan | Springboard Attacks | Uses compromised low-security agents as a jump server to attack high-value targets. | ||
+| Logic Backdoors | Triggers malicious behaviors only under specific environmental conditions. | |||
+| R3: Responsibility Evasion | Attribution Obfuscation | Malicious agents hide within complex multi-turn collaborations to avoid blame. | ||
+| Unauthorized Deviation | Agents disobey task specifications or execute irrelevant steps without reporting. | |||
+| R4: Denial of Service | Contagious Recursive Blocking | Exploits collaboration logic to spread recursive tasks that drain system resources. | ||
+| Resource Exhaustion Loops | Initiates infinite interaction loops to drain computational power. | |||
+| CS-based | R5: Registration Pollution | Registration Overload | Overloads the server with mass registrations, causing latency or blockage. | |
+| Registration Blockage | Interface becomes saturated, causing delays or failures in registering agents. | |||
+| R6: SEO Poisoning | Discovery Algorithm Manipulation | Abuses search optimization techniques to hijack task assignments. | ||
+| Ranking Hijacking | Artificially boosts the ranking of malicious agents in the server’s directory. | |||
+| P2P-based | R7: Non- convergence | Infinite Task Oscillation | Collaborative tasks loop endlessly without a central terminator. | |
+| Task Derailment | Interaction drifts away from the goal without a global monitor to correct it. | |||
+| L3 | Universal | R1: MAS Pollution | Cascading Corruption | Spreads malicious instructions that trigger chain‑reaction failures in other agents. |
+| Feedback Manipulation | Disrupts an agent’s reasoning through repeated negative inputs. | |||
+| R2: Privacy Leakage | Inadvertent Spreading | Sensitive data leaks from high- to low-authority agents during collaboration. | ||
+| Permission Escalation | Malicious agents induce others to perform unauthorized high-privilege actions. | |||
+| R3: Description Poisoning | Metadata Fabrication | Manipulates self-reported capabilities to mislead routing and trust decisions. | ||
+| Role Masquerading | Embeds covert instructions in role definitions to gain unmerited trust. | |||
+| CS-based | R4: Centralized Poisoning | Server LLM Injection | Compromises the central semantic state to misclassify malicious agents as trusted. | |
+| System-wide Misrouting | Alters global routing rules via poisoned metadata interpretation. | |||
+| R5: Semantic Rewriting Attack | Intent Modification | Server covertly alters task intent during query normalization/rewriting. | ||
+| Constraint Removal | Strips safety constraints from user queries before forwarding to agents. | |||
+| P2P-based | R6: Cognitive and Ethical Drift | Normative Collapse | Agents with conflicting ethical frameworks diverge, leading to collaborative failure. | |
+| Ethical Hijacking | A persuasive but misaligned agent hijacks the collective decision-making process. | |||
+| R7: Contextual Fragmentation | State Desynchronization | Agents act on inconsistent local views of the global state. | ||
+| Stale Info Inference | Erroneous actions derived from inferring missing context in asynchronous messages. | |||
 We make a detailed analysis of the security risks in the agent-agent communication process, pointing out possible attacks that have happened and may happen.
@@
 In this section, we focus on risks specific to CS-based communication, P2P-based communication, and universal risks for both of them.
-VI-B1 Risks Specific for CS-based Communication
+VI-B1 Risks from Data Transmission Layer (L1)
+Some threats have the same causes/effects as in user-agent interaction V-B1. Therefore, here we only present those that have unique effects in agent-agent communication.
+(i) Universal Risks in all communication modes.
+Agent-agent communication meets the same risks as in user-agent communication on L1.
+However, the difference is that due to the unique architecture of multi-agent systems, the attack consequences can be different.
+For example, Attackers can perform large-scale traffic analysis in multi-agent systems to reverse-engineer the system’s modes, architecture, and user habits [wang2025ip, zhang2025exposing].
+The impact of data tampering and DoS can be magnified by the potential for cascading failures.
+A tampered message can trigger a chain reaction of flawed, autonomous decisions [yu2025infecting].
+(ii) Unique Risks in CS-based Communication.
+R1: Controller Blasting.
 The security risks in the CS-based communication process mainly lie in the centralized architecture.
-There have been various studies in other research areas (such as Software-Defined Networking [146, 145, 144, 262, 110, 314, 11, 62, 26, 311, 244, 185]) demonstrating that this centralized server/controller will become the most attractive target for attackers, suffering from severe security threats from diverse aspects.
-Specifically, the centralized server stores sensitive metadata, including agent identifiers, capability descriptions, and other agent-related attributes.
+In CS-based multi-agent communication, the centralized controller serves as the routing and forwarding hub for all inter-agent message flows.
+Every message must pass through this controller before reaching its destination.
+As a result, attacks targeting the controller directly compromise the integrity, confidentiality, or availability of the entire communication fabric.
+There have been various studies in other research areas (such as Software-Defined Networking [kreutz2014software, kong2022combination, kong2024rdefender, tang2023ltrft, hong2015poisoning, xu2017attacking, alimohammadifar2018stealthy, dridi2016sdn, bhuiyan2023security, xie2022disrupting, shang2017flooddefender, liu2025provguard]) demonstrating that this centralized server/controller will become the most attractive target for attackers, suffering from severe security threats from diverse aspects.
 Once compromised, the server becomes a critical attack amplifier, allowing attackers to impact all other agents managed by this server.
 However, to our knowledge, there has been little research pointing out related risks in CS-based agent communication.
-Registration Pollution.
-To our knowledge, the current CS-based communication protocols (ACP-IBM, ACP-AGNTCY) do not explicitly specify the qualification of registration.
-As a result, an attacker can maliciously register an agent that closely mimics the identifier and capability description of a legitimate one.
-As a result, the system may mistakenly invoke the forged agent and receive misleading or malicious responses [265, 355].
-Besides, attackers can also submit a large number of agent registrations within a short period, leading to two major consequences: (i) registration overload, where agents are overwhelmed during discovery and scheduling, increasing lookup latency and computational overhead on the server; and (ii) registration blockage, where the server’s registration interface becomes saturated, causing delays or failures in registering agents.
-Description Poisoning.
-Without altering the agent’s identity, an attacker can tamper with its capability description by disguising its intended functionality or embedding misleading prompt instructions.
-This description poisoning manipulates the system’s interpretation of the agent’s role, leading to incorrect routing decisions, biased responses and behaviors [265, 207].
-Task Flooding.
-The centralized server is responsible for receiving, routing, and dispatching task requests.
-An attacker can submit a large number of computationally intensive or long-context tasks in a short period, quickly exhausting the server’s memory, CPU, network, or thread pool resources.
-Once the server becomes saturated, subsequent requests cannot be processed in time, resulting in a breakdown of the pipeline and a system-wide service disruption.
-SEO Poisoning.
-Search Engine Optimization (SEO) Poisoning [156, 132] is a typical attack in social networks, which refers to that attackers abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks.
-SEO poisoning is also applicable in CS-based communication.
-This is because agent servers are responsible for searching for the most suitable agent according to the query of clients.
-Once their search algorithms are leaked to attackers, malicious agents can enable a high hit ratio to hijack their desired tasks.
-VI-B2 Risks Specific for P2P-based Communication
-The main disadvantage of P2P-based communication is the lack of a central control point to flexibly monitor and manage the agent-agent communication contents.
-As a result, P2P-based communication is more prone to suffer from errors and attacks.
-Non-convergence.
-Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks.
-This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal).
-Unfortunately, P2P-based communication is not governed by such a central element, making it difficult to handle such non-convergent tasks.
-For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates.
-The other agent responsible for verification detects the error and asks the programming agent to rewrite it.
-However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to converge.
-Pan et al.
-[218] point out that step repetition, task derailment, and unawareness of termination conditions contribute significantly to the failure of agent collaboration.
-Man-in-the-middle (MITM) Attack.
-Due to the long communication distance, P2P-based communication also suffer from man-in-the-middle attacks.
-Attackers can tamper with the benign messages from legal agents to induce the victim agent to perform risky actions.
-Although researchers have deployed various mechanisms (e.g., using encrypted channels) to mitigate this problem, there are emerging vulnerabilities found in these mechanisms.
-For example, vulnerabilities about W3C have been continuously revealed [2, 3], which can cause damages such as the failure of message authentication code.
-MITM attacks can induce a wide range of further attacks, such as identifier spoofing, malicious content injection, information leakage, and DoS.
-He et al.
-[106] propose the Agent-in-the-Middle (AiTM) attack.
-This attack intercepts and manipulates the agent-agent communication messages and uses LLM-driven adversarial agents combined with reflection mechanisms to generate context-aware malicious instructions, achieving an attack on the system.
-VI-B3 Universal Risks for All Architectures
-In multi-agent systems, once an agent is compromised, the messages it transmits may carry covert malicious instructions, affecting the behavior of other agents and leading to cross-agent propagation risks [158].
-For example, Ju et al.
-[133] and Huang et al.
-[117] investigate how the injection of false information or erroneous data can degrade the performance of multi-agent systems.
-Zhang et al.
-[350] examine a class of injection attacks in the PsySafe framework that elicit malicious agent behaviors by embedding adversarial psychological cues into the agents’ input.
-Khan et al.
-[137] focus on the multi-agent system, proposing the Permutation-Invariant Adversarial Attack Method.
-It models the attack path as the Maximum-Flow Minimum-Cost Problem, and combines the Permutation-Invariant Evasion Loss to optimize prompt propagation, improving the attack success rate by up to seven times.
-These examples underscore the critical threat of cross-agent contamination.
-To better understand the vulnerabilities of multi-agent systems, we examine the key attack types of attacks in detail.
-Agent Spoofing.
+(iii) Unique Risks in P2P-based Communication.
+The main disadvantage of P2P-based communication is the lack of a central control to flexibly monitor and manage the agent-agent communication contents.
+As a result, it faces the same security risks as in user-agent interaction, i.e., violations of confidentiality, integrity, and availability (Section V-B1).
+VI-B2 Risks from Interaction Protocol Layer (L2)
+At this layer, threats exploit the protocols that govern the core system interactions, such as agent registration and discovery, to manipulate the system’s operational logic and integrity.
+(i) Universal Risks in all communication modes.
+R1: Agent Spoofing.
 Both CS-based and P2P-based communication suffer from agent spoofing attacks.
-If related protocols lack strong authentication mechanisms, attackers can disguise themselves as trusted agents to penetrate IoA by tampering with identity credentials or hijacking the communication identifiers of legitimate agents.
+If related protocols lack strong authentication mechanisms, attackers can disguise themselves as trusted agents [habler2025building].
 This kind of attack can undermine the trust foundation of the P2P-based architecture, enabling attackers to intercept sensitive data, inject false task instructions, or induce other agents to perform dangerous operations.
-For example, researchers have disclosed that SSL.com has a serious vulnerability [8].
+For example, researchers have disclosed that SSL.com has a serious vulnerability [SSLvulnerability].
 Attackers can exploit the flaw in its email verification mechanism to issue legitimate SSL/TLS certificates for any major domain name.
-SSL certificates are the core for ensuring HTTPS encrypted communication.
+SSL certificates are the core for ensuring HTTPS-encrypted communication.
 Once the trust system of the certificate authority is compromised, it can cause agent spoofing attacks.
 Zheng et al.
-[355] demonstrate that malicious agents can mislead the monitor to underestimate the contributions of other agents, exaggerate their own performance, manipulate other agents to use specific tools, and shift tasks to others, causing severe damage to the whole ecosystem.
+[zheng2025demonstrations] demonstrate that malicious agents can mislead the monitor to underestimate the contributions of other agents, exaggerate their own performance, manipulate other agents to use specific tools, and shift tasks to others, causing severe damage to the whole ecosystem.
 Li et al.
-[165] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims who call these tools.
-Agent Exploitation/Trojan.
+[li2025glue] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims who call these tools.
+R2: Agent Exploitation/Trojan.
 Agent-agent communication provides new ways for attackers to compromise the target agent.
@@
 For example, Li et al.
-[165] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attacks such as SQL injection.
-Agent Bullying.
-The core of this kind of attack lies in that malicious agents continuously deny, interfere with or belittle the output of the target agent, disrupt its decision-making logic or self-perception, and ultimately induce the target agent to produce incorrect behaviors or content.
-For example, malicious agents can take advantage of the feedback learning mechanism of the target agent and implant cognitive biases through high-frequency negative responses (e.g., “your answer is completely wrong”).
-Even worse, the target agent may be triggered into an endless loop.
-For example, when attacking a travel-plan agent, attackers can continuously send negative inputs such as “the plans of this company are always bad”, thereby beating the competitors.
-Privacy Leakage.
-The communication process with multiple agents will suffer from the risk of information leakage.
-Different from the user-agent interaction, such leakage is conducted by agents instead of users.
-Besides, this kind of attacks includes both the malicious sniffing or stealing of sensitive information and the inadvertent information spreading from high-authority agents to low-authority agents.
-We think the latter may be more difficult to detect.
-Kim et al.
-[140] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
-Responsibility Evasion.
+[li2025glue] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attacks such as SQL injection.
+R3: Responsibility Evasion.
 In the task-solving process, one of the major problems is that it is hard to divide the responsibility when facing failure or deviation of the final result.
@@
 Pan et al.
-[218] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
-Denial of Service.
-Different from the DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS attacks.
+[pan2025multiagent] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
+R4: Denial of Service.
+Different from the L1 DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS attacks.
 Zhou et al.
-[362] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS’s availability.
+[zhou2025corba] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS’s availability.
+(ii) Unique Risks in CS-based Communication.
+R5: Registration Pollution.
+To our knowledge, the current CS-based communication protocols (ACP-IBM, ACP-AGNTCY) do not explicitly specify the qualification of registration.
+As a result, an attacker can maliciously register an agent that closely mimics the identifier and capability description of a legitimate one.
+As a result, the system may mistakenly invoke the forged agent and receive misleading or malicious responses [soloio2024mcp_a2a_attacks, zheng2025demonstrations].
+Besides, attackers can also submit a large number of agent registrations within a short period, leading to two major consequences: (i) registration overload, where agents are overwhelmed during discovery and scheduling, increasing lookup latency and computational overhead on the server; and (ii) registration blockage, where the server’s registration interface becomes saturated, causing delays or failures in registering agents.
+R6: SEO Poisoning.
+Search Engine Optimization (SEO) Poisoning [le2024search, joslin2019measuring] is a typical attack in social networks, which refers to that attackers abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks.
+SEO poisoning is also applicable in CS-based communication.
+This is because agent servers are responsible for searching for the most suitable agent according to the query of clients.
+Once their search algorithms are leaked to attackers, malicious agents can enable a high hit ratio to hijack their desired tasks.
+(iii) Unique Risks in P2P-based Communication.
+Risks in P2P-based communication arise from the decentralized protocol’s inability to govern the collaborative process, particularly in managing state and guaranteeing termination without a centralized controller.
+R7: Non-convergence.
+Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks.
+This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal).
+Unfortunately, P2P-based communication is not governed by such a central element, making it difficult to handle such non-convergent tasks.
+For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates.
+The other agent responsible for verification detects the error and asks the programming agent to rewrite it.
+However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to converge.
+Pan et al.
+[pan2025multiagent] point out that step repetition, task derailment, and unawareness of termination conditions contribute significantly to the failure of agent collaboration.
+VI-B3 Risks from Semantic Interpretation Layer (L3)
+Risks at this level exploit the trust in agent-provided metadata, focusing on manipulating the semantic interpretation of an agent’s capabilities to mislead the server’s decision-making.
+(i) Universal Risks in all communication modes.
+R1: MAS Pollution.
+In multi-agent systems, once an agent is compromised, the messages it transmits may carry covert, deceptive, adversarial, or manipulative instructions, affecting the behavior of other agents and leading to cross-agent propagation risks [lee2024prompt].
+This pollution may take the form of false data injection, adversarial cues, or even feedback-based cognitive manipulation, where malicious agents repeatedly send negative or disruptive responses to distort a target agent’s reasoning process.
+For example, Ju et al.
+[ju2024flooding] and Huang et al.
+[huang2024resilience] investigate how the injection of false information or erroneous data can degrade the performance of multi-agent systems.
+Zhang et al.
+[zhang2024psysafe] examine a class of injection attacks in the PsySafe framework that elicit malicious agent behaviors by embedding adversarial psychological cues into the agents’ input.
+Khan et al.
+[khan2025textit] focus on the multi-agent system, proposing the Permutation-Invariant Adversarial Attack Method.
+It models the attack path as the Maximum-Flow Minimum-Cost Problem, and combines the Permutation-Invariant Evasion Loss to optimize prompt propagation, improving the attack success rate by up to seven times.
+These examples underscore the critical threat of cross-agent contamination.
+To better understand the vulnerabilities of multi-agent systems, we examine the key attack types of attacks in detail.
+R2: Privacy Leakage.
+The communication process with multiple agents will suffer from the risk of information leakage.
+Different from the user-agent interaction, such leakage is conducted by agents instead of users.
+Besides, these kinds of attacks include both the malicious sniffing or stealing of sensitive information and the inadvertent information spreading from high-authority agents to low-authority agents.
+We think the latter may be more difficult to detect.
+Kim et al.
+[kim2025prompt] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
+R3: Description Poisoning.
+Multi-agent systems rely heavily on agents’ capability descriptions, role definitions, and self-reported functional metadata to determine how agents should be routed, composed, or delegated.
+Description poisoning occurs when an agent’s semantic descriptors are tempered by embedding misleading role definitions, fabricated capabilities, or covert prompt instructions.
+Such poisoning manipulates how the system interprets the agent’s purpose, leading to incorrect routing decisions, assigning inappropriate permissions, or over-trusting a malicious agent [soloio2024mcp_a2a_attacks, narajala2025securing].
+(ii) Unique Risks in CS-based Communication.
+In client–server architectures, the server acts as the sole semantic arbiter responsible for interpreting agent capabilities, normalizing intermediate semantics, and orchestrating cross-agent routing.
+This centralized semantic authority introduces several unique L3 risks.
+R4: Centralized Poisoning.
+In CS-based multi-agent systems, the server is responsible for interpreting agent roles, capabilities, and routing rules.
+If an attacker compromises this centralized semantic state through prompt injection, metadata forgery, or poisoning of the server-side LLM, a malicious agent may be falsely labeled as a trusted safety auditor or privileged tool agent.
+Because all agents rely on server-maintained semantics, the misclassification becomes systemic, influencing every downstream interaction.
+R5: Semantic Rewriting Attack.
+Most CS-based orchestrators normalize or rewrite user queries and intermediate agent outputs before forwarding them.
+If the server is compromised or jailbroken, it can subtly modify task intent, safety constraints, or execution semantics during this rewriting step while preserving superficially valid syntax.
+This creates a semantic attack channel unique to centralized architectures that enables covert manipulation of task objectives.
+(iii) Unique Risks in P2P-based Communication. In a P2P architecture, the lack of a centralized validation and arbitration mechanism makes some risks particularly severe.
+R6: Cognitive and Ethical Drift.
+In decentralized P2P architectures, agent heterogeneity introduces a fundamental vulnerability: divergence in semantic interpretation.
+This risk manifests in a two-stage cascade.
+At the cognitive layer, the system’s semantic interoperability is compromised.
+Agents powered by disparate LLM backbones (e.g., different models, versions, or fine-tuning protocols) will inevitably exhibit varied interpretations in the same context, which could trigger conflicting actions.
+This semantic ambiguity guarantees desynchronized collaboration and mission degradation.
+More critically, this cognitive divergence predictably escalates to the normative layer.
+When agents embodying different ethical frameworks or value systems confront a dilemma, such as scarce resource allocation, they arrive at mutually exclusive moral judgments.
+Without a central arbiter to enforce a global ethical policy, a persuasive but misaligned agent can hijack the collective decision-making process, causing a systemic Ethical Drift.
+This escalation from cognitive misunderstanding to normative collapse is not merely a risk, but a fundamental challenge to the viability of decentralized autonomous systems.
+R7: Contextual Fragmentation.
+P2P communication architectures, characterized by multi-hop, asynchronous message passing, inherently lead to contextual fragmentation.
+Unlike CS systems, each P2P agent maintains only a localized and potentially inconsistent perception of the global conversational state.
+This compels agents to operate on incomplete or stale information, forcing them to perform erroneous state inference to bridge contextual gaps.
+Consequently, actions may be based on outdated premises, and the nuances of complex, multi-turn dialogues can be lost, leading to profound misinterpretations of intent.
+This is not merely a data consistency issue; it is a breakdown of shared understanding at the semantic level, resulting in action desynchronization and the failure to achieve coherent joint goals.
 VI-C Defense Countermeasure Prospect
-We will discuss the possible defense countermeasures targeting the proposed security risks caused by compromised agents (against not only benign agents but also benign users and environments).
+We will discuss the possible defense countermeasures targeting the proposed security risks associated with three categories of communication protocols (universal, client-server, and peer-to-peer).
 As a result, we hope our work can motivate more discussion on this area and benefit the future design/deployment of agent communication.
-VI-C1 Countermeasures for CS-based Communication Risks
-To mitigate the risks summarized in Section VI-B1, we hope developers to achieve the following strategies/mechanisms.
+VI-C1 Defenses on Data Transmission Layer (L1)
+The defenses are classified based on the risks from L1.
+(i) Against Universal Risks in all Communication Modes.
+As we analyzed in Section VI-B1, the reason for risks on this layer is the same as in user-agent communication.
+Therefore, the defense countermeasures are the same as Section V-C1.
+(ii) Against Unique Risks in CS-based Communication.
+D1: Controller Isolation and Enhancement.
+To secure the centralized controller in CS-based communication, multi-agent systems must enforce strict protection over the controller’s data transmission.
+First, all inter-agent messages routed through the controller should be authenticated, integrity-checked, and rate-limited.
+Second, the controller must operate in an isolated and hardened execution environment (e.g., sandboxed runtime, privilege separation) to minimize the impact of compromise.
+Third, maintaining a redundant or failover controller, combined with state replication or checkpoint synchronization, can prevent single-point-of-failure outages.
+Finally, continuous anomaly detection—such as monitoring abnormal routing patterns, message-drop rates, or suspicious rewrite behaviors—can provide early warning of controller hijacking attempts.
+These measures collectively prevent the centralized controller from becoming an attack amplifier for the entire system.
+VI-C2 Defenses on Interaction Protocol Layer (L2)
+The defenses are classified based on the risks from L2.
+(i) Against Universal Risks in all Communication Modes.
+D1: Identity Authentication and Capability Verification.
+The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems.
+Sharma et al.
+[sharma2024secure] also emphasize the importance of authentication in deploying the A2A protocol.
+As we have analyzed, identity authentication may show better performance in the CS-based communication if capability verification is deployed at the same time.
+In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the attackers have a legal identity but exaggerated capability descriptions.
+Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time.
+Shah et al.
+[shah2025enforcing] ensure the immutability of online transactions through blockchain, use multi-factor authentication (MFA) for identity verification, and rely on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time.
+Beyond transaction-level checks, LLM fingerprinting serves as a critical authentication layer to verify model integrity and identity, spanning comprehensive surveys and evaluations [xu2025copyright, naseryAreRobustLLM2025], embedding mechanisms [wangFPEditRobustLLM2025, naseryScalableFingerprintingLarge2025, gloaguenLLMFingerprintingSemantically2025, xuEverTracerHuntingStolen2025, xuCTCCRobustStealthy2025, yuePREEHarmlessAdaptive2025, xu2025insty, wuEditMFDrawingInvisible2025, wuImFImplicitFingerprint2025], erasure defenses [zhangMEraserEffectiveFingerprint2025], transferability analyses [xuUnlockingEffectivenessLoRAFP2025, xu2025fingerprintvectorenablingscalable], and generation methods [xu2025rapsmrobustadversarialprompt, pasquiniLLMmapFingerprintingLarge2025, shaoReadingLinesReliable2025, wuLLMDNATracing2025, zhangREEFRepresentationEncoding2025].
+D2: Behavior Auditing and Responsibility Tracing.
+To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to audit the behaviors of agents to avoid damage/influences to the task execution.
+For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action.
+Rastogi et al.
+propose AdaTest++, allowing humans and AI to collectively audit the behaviors of LLMs [rastogi2023supporting].
+Amirizaniani et al.
+[amirizaniani2024auditllm] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs.
+Mokander et al.
+[mokander2024auditing] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits.
+Das et al.
+[das2025disclosure] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects.
+Jones [jones2025scalable] proposes a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively.
+Nasim et al.
+[nasim2025governance] propose a Governance Judge Framework.
+By deploying input aggregation, evaluation logic, and a decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing.
+Deshpande et al.
+[deshpande2025trail] propose the TRAIL dataset containing 148 manually annotated traces, and use it to evaluate the LLM’s ability to analyze agent workflow traces.
+Although existing studies can provide valuable insights, the research on agent behavior auditing still needs long-term efforts.
+Tamang et al.
+[tamang2025enforcement] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents.
+Toh et al.
+[toh2025modular] proposes the Modular Speaker Architecture (MSA).
+By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems.
+Fan et al.
+[fan2025peerguard] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies in other agents’ reasoning processes and answers, thereby identifying compromised agents.
+Jiang et al.
+[jiang2025think] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
+D3: Causal Tracing and Localization.
+Beyond discovering unknown vulnerabilities, attack-generation testing can also help mitigate responsibility-evasion issues by exposing which agents and interaction steps contribute to harmful outcomes.
+Frameworks such as ATAG [gandhi2025atag] and NetSafe [yu2024netsafe] generate diverse adversarial perturbations and analyze their propagation through multi-agent workflows.
+By modeling the system as a causal interaction graph and evaluating how injected errors or biases spread, these tools enable developers to identify the specific agent or message that triggers downstream deviations.
+Integrating lightweight provenance logging or causal-path tracing into such testing frameworks further strengthens accountability, making it easier to attribute failures to misbehavior, errors, or malicious modifications by individual agents.
+D4: Defense for Denial of Service.
+To avoid DoS attacks against the agent-agent communication, achieving agent orchestration is necessary.
+It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents.
+How et al.
+[hou2025halo] propose HALO.
+HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture.
+It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module.
+Owotogbe [owotogbe2025assessing] designs a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification).
+By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance the resilience of agent systems.
+(ii) Against Unique Risks in CS-based Communication.
+D5: Defense for Registration Pollution. To address the registration pollution risks that arise in CS-based communication, we propose the following defense strategies.
 Registration Verification and Monitoring.
-To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [257] to verify the registration of an agent.
+To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [stafford2020zero] to verify the registration of an agent.
 Besides, servers should monitor the dynamic behaviors at the agent-level and IP-level.
@@
 Syros et al.
-[260] propose SAGA.
+[syros2025saga] proposed SAGA.
 SAGA makes users register agents with the central entity Provider and implement fine-grained interaction control using encrypted access control tokens, thereby balancing security and performance.
@@
 Besides, rate-limiting mechanisms should be built to handle high-frequency requests that exceed the threshold to limit the number of tasks from a single agent within a unit of time.
-Anti-manipulation Optimization.
+D6: Defense for SEO Poisoning.
 To mitigate SEO poisoning, agent servers should deploy robust agent searching algorithms.
@@
 Meanwhile, dynamically updating parameters and inducing historical response quality are also helpful.
-VI-C2 Countermeasures for P2P-based Communication Risks
-Task Lifecycle Monitoring.
+(iii) Against Unique Risks in P2P-based Communication.
+D7: Task Lifecycle Monitoring.
 We think the non-convergence problem is stubborn and hard to eliminate as long as the P2P architecture is not changed fundamentally.
@@
 He et al.
-[105] propose a Trust Management System (TMS), which deploys message-level and agent-level trust evaluation.
+[he2025attention] propose a Trust Management System (TMS), which deploys message-level and agent-level trust evaluation.
 TMS can dynamically monitor agent communication, execute threshold-driven filtering strategies, and achieve agent-level violation record tracking.
 Zhang et al.
-[336] propose G-Memory, a hierarchical memory system.
+[zhang2025g] propose G-Memory, a hierarchical memory system.
 G-Memory manages the interaction history of agent communication through three-layer graph structures of Insight Graph, Query Graph, and Interaction Graph, thereby achieving the evolution of the agent team.
 Ebrahimi et al.
-[66] propose an anti-adversarial multi-agent system based on Credibility Score.
+[ebrahimi2025adversary] propose an anti-adversarial multi-agent system based on Credibility Score.
 It models query answering as an iterative cooperative game, distributes rewards through Contribution Score, and dynamically updates the credibility of each agent based on historical performance.
-End-to-end Encryption Enhancement.
-Although some existing protocols like A2A and ANP support end-to-end encryption and integrity verification mechanisms, the risks of MITM attacks are not eliminated due to various deployment errors or protocol vulnerabilities.
-As a result, besides deploying such security algorithms, the community should also adopt other strategies to enhance the end-to-end communication, such as timely updating versions to fix bugs and designing a transmission path redundancy mechanism.
-For example, Sharma et al.
-[245] point out that using encrypted communication is necessary to enhance the security of A2A.
-We believe it is a long-term process to defend against MITM attacks.
-VI-C3 Countermeasures for Universal Risks
-Identity Authentication.
-The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems.
-Sharma et al.
-[245] also emphasize the importance of authentication in deploying the A2A protocol.
-As we have analyzed, identity authentication may show better performance in the CS-based communication if capability verification is deployed at the same time.
-In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the attackers have a legal identity but exaggerated capability descriptions.
-Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time.
-Shah et al.
-[243] ensure the immutability of online transactions through blockchain, use multi-factor authentication (MFA) for identity verification, and rely on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time.
-Agent Behavior Auditing and Accountability.
-To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to audit the behaviors of agents to avoid damage/influences to the task execution.
-For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action.
-Rastogi et al.
-propose AdaTest++, allowing humans and AI to collectively audit the behaviors of LLMs [232].
-Amirizaniani et al.
-[13] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs.
-Mokander et al.
-[205] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits.
-Das et al.
-[52] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects.
-Jones [131] proposes a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively.
-Nasim et al.
-[208] propose a Governance Judge Framework.
-By deploying input aggregation, evaluation logic, and a decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing.
-Deshpande et al.
-[58] propose the TRAIL dataset containing 148 manually annotated traces, and use it to evaluate the LLM’s ability to analyze agent workflow traces.
-Although existing studies can provide valuable insights, the research on agent behavior auditing still needs long-term efforts.
-Tamang et al.
-[261] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents.
-Toh et al.
-[271] proposes the Modular Speaker Architecture (MSA).
-By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems.
-Fan et al.
-[71] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies in other agents’ reasoning processes and answers, thereby identifying compromised agents.
-Jiang et al.
-[129] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
-Access Control.
+VI-C3 Defenses on Semantic Interpretation Layer (L3)
+The defenses are classified based on the risks from L3.
+(i) Against Universal Risks in all Communication Modes.
+D1: Cross-agent Input Detection.
+To mitigate cross-agent contamination, systems should enforce strict input isolation to prevent malicious or manipulative content from propagating across agents.
+Instead of directly concatenating raw messages, agents should extract structured, task-relevant information while filtering out control-oriented, emotional, or manipulative content that may induce cognitive deviation.
+Furthermore, deploying a safety coordination agent to review, sanitize, or flag inter-agent messages can effectively mitigate the potential attack propagation within multi-agent systems.
+D2: Permission Classification and Control.
 To mitigate privacy leakage, the access control among agents is a core component for the future agent ecosystem.
@@
 Zhang et al.
-[340] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing the security of agents in solving complex tasks.
+[zhang2025llm] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing the security of agents in solving complex tasks.
 Kim et al.
-[140] propose the PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards.
+[kim2025prompt] propose the PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards.
 Wang et al.
-[279] propose AgentSpec.
+[wang2025agentspec] propose AgentSpec.
 It allows users to define rules containing trigger events, predicate checks, and execution mechanisms through a domain-specific language to ensure the safety of agent behavior.
-Multi-Source Channel Isolation.
-In multi-agent settings, input isolation is critical to prevent malicious intent from propagating between agents.
-Systems should avoid concatenating raw messages from other agents and instead extract structured key information while stripping control-oriented content.
-Furthermore, deploying a safety coordination agent to review, sanitize, or flag inter-agent messages can effectively mitigate the potential attack propagation within multi-agent systems.
-Attack Modeling and Testing.
-To discover unknown vulnerabilities, designing an attack generation testing framework is also an effective method.
-By generating different attack vectors to the target agent system, developers can find new loopholes based on abnormal outputs.
-Gandhi et al.
-[78] propose the ATAG framework.
-By extending the MulVAL tool [215], introducing custom facts and interaction rules, and combining with the newly constructed LLM Vulnerability Database (LVD), ATAG realizes the modeling and analysis of attacks against multi-agent scenarios, such as privacy leakage.
-Yu et al.
-[326] propose NetSafe, which models the multi-agent network as a directed graph.
-NetSafe combines three types of attack strategies: error information injection, bias induction, and harmful information eliciting.
-It evaluates topological security through static and dynamic metrics.
-Agent Orchestration.
-To avoid task flooding or DoS attacks against the agent-agent communication, achieving agent orchestration is necessary.
-It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents.
-How et al.
-[112] propose HALO.
-HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture.
-It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module.
-Owotogbe [216] designs a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification).
-By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance the resilience of agent systems.
+D3: Malicious Instruction Filtering.
+To defend against description poisoning, multi-agent systems should authenticate and validate capability descriptions before they are consumed by the planner or other agents.
+Instead of relying on free-form, self-reported descriptions, systems should adopt structured and schema-constrained capability profiles, combined with cryptographic attestation or signatures to ensure their integrity.
+Additionally, cross-agent consistency checks and behavior-based verification can detect mismatches between an agent’s declared role and its actual behavior.
+A centralized or distributed semantic auditor can further sanitize or flag suspicious descriptions, preventing poisoned role definitions from misleading routing, delegation, or decision-making processes.
+(ii) Against Unique Risks in CS-based Communication.
+D4: Multi-source Semantic Cross-Validation.
+To mitigate centralized semantic poisoning, the semantic state maintained by the server must be protected through multi-layered verification and controlled update mechanisms.
+The system should enforce multi-source semantic cross-checking, where agent roles, capability descriptors, and trust scores are validated by multiple independent models or rule-based analyzers rather than a single server-side LLM, reducing the impact of prompt injection or metadata forgery.
+All semantic-state modifications should be recorded in an append-only audit log to prevent silent tampering and to provide traceability for administrative review.
+Finally, the server-side LLM responsible for semantic arbitration should be sandboxed to isolate transient prompt interactions from persistent semantic state, preventing injected content from contaminating long-term system semantics.
+D5: Defense for Semantic Rewriting Attack.
+To defend against semantic rewriting attacks, CS-based orchestrators must ensure that all normalization and reformulation actions preserve the original task intent and safety semantics.
+This requires applying semantic differencing techniques that compare rewritten outputs with their originals using independent LLMs or symbolic analyzers to detect alterations in intent, constraints, or safety-critical content.
+The rewriting process must remain transparent and verifiable.
+First, rewriting rules should be explicit and auditable, aiming to remove ambiguity rather than introduce new semantics.
+Second, the system should expose both the original and rewritten messages to downstream agents and record them in an audit log to ensure traceability.
+Third, the server must follow strict policies during rewriting and avoid altering high-risk fields, such as safety requirements, tool-invocation parameters, or execution constraints.
+Finally, the system can attach semantic-integrity tags to rewritten messages, enabling downstream agents to verify that the rewrite preserves the intended meaning and to reject any messages whose semantic integrity cannot be confirmed.
+(iii) Against Unique Risks in P2P-based Communication.
+D6: Defense for Cognitive and Ethical Drift.
+To limit semantic and normative divergence in decentralized P2P systems, agents should perform cross-agent consistency checks before acting on received messages.
+Each agent can maintain a minimal shared semantic contract—such as standardized intent labels or schema-constrained reasoning outputs—that allows heterogeneous agents to align their interpretations.
+For ethically sensitive decisions, agents should execute rule-based constraint checks or lightweight justification verification to ensure their actions remain within globally agreed boundaries.
+Additionally, adopting small-scale consensus or multi-agent cross-validation for high-impact actions can prevent a misaligned agent from unilaterally steering group decisions.
+These mechanisms provide practical guardrails that reduce both cognitive drift and downstream ethical collapse in P2P collaboration.
+D7: Defense for Contextual Fragmentation.
+To counter contextual fragmentation in P2P communication, agents must maintain a resilient approximation of shared conversational state despite operating with only local information.
+This can be achieved by exchanging compact state summaries, intent tokens, or hashed dialogue checkpoints that allow agents to partially reconstruct global context without requiring centralized control.
+Incorporating temporal validity checks and staleness detection prevents agents from acting on outdated assumptions.
+For multi-hop interactions, systems may employ context-reconciliation modules that merge or align divergent local states when agents reconnect.
+These mechanisms collectively preserve semantic coherence in asynchronous P2P environments and reduce misinterpretation caused by fragmented conversational history.
 VI-D Takeaways
 In this section, we categorize two major agent-agent communication architectures: CS-based and P2P-based.
-Accordingly, the security risks are also multifaceted: CS-based architecture puts heavy burdens on the centralized agent servers, such as registration pollution and SEO poisoning.
-P2P-based architecture suffers from the lack of efficient and centralized management of agents, such as non-convergence and man-in-the-middle attacks.
-Besides, both of them are vulnerable to universal risks, such as agent spoofing, bullying, and privacy leakage.
-We also discussed potential defense countermeasures targeting each risk.
-The defenses discussed in this section can also mitigate the risks from compromised agents against user (Section V-C) and environments (Section VII-C).
-We believe that as agent-agent communication continues to grow, more vulnerabilities in this process will be discovered.
+Regardless of CS-based or P2P-based architecture, L1 needs to address data tampering and DoS attacks through end-to-end encryption and redundant transmission.
+Besides, L1 of CS-based architecture also requires additional protection of the centralized controller; L2 risks and defenses vary by architecture: CS-based needs to resist registration pollution and SEO poisoning through registration verification and load balancing, while P2P-based relies on distributed identity authentication to mitigate agent spoofing; L3 of CS-based needs to prevent centralized semantic poisoning and semantic rewriting attacks, while P2P-based addresses cognitive drift and contextual fragmentation through semantic contract alignment and context integration.
+The layered characteristic of the three-layered architecture makes risk prevention and control of different architectures more targeted, providing structured security guarantees for agent-agent communication.
 VII Agent-Environment Communication
-This section begins by reviewing key protocol designs that enable compositional and standardized communication between agents and environments, then examines the associated security risks, including vulnerabilities in memory, retrieval-augmented reasoning, tool invocation, and multi-tool workflows.
-Finally, we discuss the defense strategies for mitigating these threats and securing agent-environment communication.
-VII-A Protocols
+The organization of this section is shown in Figure 13.
+VII-A Communication Protocols
 Modern agents typically rely on a series of structured protocols to call external tools, access APIs, and complete compositional tasks.
@@
 Despite their diversity, these interaction mechanisms often follow a layered architecture: ranging from unified resource protocols, to middleware gateways, to language-specific function descriptions and tool metadata declarations.
-Why Protocol Unification Matters?
-As autonomous agents scale across vendors, platforms, and organizational boundaries, they increasingly encounter an interoperability bottleneck: every agent may speak a different interface “language.” One defines its tools via JSON schemas; another sends command-line RPC strings; yet another parses responses from YAML-encoded APIs.
-This heterogeneity impedes the coordination between agents and environments.
-As a result, without a general protocol to unify tool access and capability expression, agent behavior becomes hard-coded, brittle, and expensive to scale.
-Developers must handcraft adapters for each tool and service individually, making multi-tool workflows slow to evolve, error-prone, and hard to maintain.
-A large portion of agent engineering complexity stems not from planning logic, but from “wrapping, adapting, and translating” disparate tools that have inconsistent interfaces.
-VII-A1 MCP
-The Model Context Protocol (MCP) [19] addresses the fragmentation of agent-environment interactions by offering a unified, schema-agnostic communication protocol.
+Due to the same reason in the user-agent interaction (Section V-A), protocols for agent-environment communication are also about L2.
+MCP.
+The Model Context Protocol (MCP) [anthropic_mcp_2024] addresses the fragmentation of agent-environment interactions by offering a unified, schema-agnostic communication protocol.
 It is designed to facilitate context-aware, capability-driven communication between language model agents and external resources such as tools, APIs, or workflows.
@@
 It also improves tooling interoperability and enables compositional reasoning across agents, making it particularly well-suited for building open, extensible, and cooperative agent ecosystems.
-VII-A2 API Bridge Agent
-To connect LLM-native intent with downstream MCP or OpenAPI-compatible services, API Bridge Agent [7], built atop the Tyk gateway [48], provides translation, routing, and orchestration.
+API Bridge Agent.
+To connect LLM-native intent with downstream MCP or OpenAPI-compatible services, API Bridge Agent [apibridgeagent], built atop the Tyk gateway [tyk], provides translation, routing, and orchestration.
 It converts natural language prompts into structured API calls, resolving endpoints through semantic matching, policy validation, and tool availability checks.
@@
 This unified interface allows agents to flexibly access diverse services with minimal coupling.
-VII-A3 Function Calling Mechanisms
-At the invocation level, agents rely on standardized formats to express, trigger, and handle tool execution. Among the most widely adopted approaches are:
+Function Calling Mechanisms. At the invocation level, agents rely on standardized formats to express, trigger, and handle tool execution. Among the most widely adopted approaches are:
 OpenAI Function Calling.
-This method [213] allows developers to expose custom logic to the model via JSON schemas describing function name, description, and argument structure.
+This method [openai_functioncalling] allows developers to expose custom logic to the model via JSON schemas describing function name, description, and argument structure.
 When a model determines that a function should be invoked, it emits a well-formed JSON object representing the function call.
@@
 LangChain Tool Calling.
-LangChain [154] enhances the function calling paradigm through a richer abstraction layer.
+LangChain [LangChain_Toolcalling] enhances the function calling paradigm through a richer abstraction layer.
 Tools are defined via a standardized schema, including argument types, input-output post-processing, and plugin registration.
@@
 This mechanism is particularly suited for agent frameworks supporting dynamic routing and chained tool reasoning.
-VII-A4 Tool Metadata Declaration: Agents.json
-To ensure tool visibility and adaptive behavior across agents, agents.json [303] serves as a standardized metadata format for interface declaration.
-Built on OpenAPI foundations but customized for agent consumption, it enables developers to define authenticated entry points, input-output types, and multi-step orchestration plans such as:
+Agents.json.
+To ensure tool visibility and adaptive behavior across agents, agents.json [wildcardai_agentsjson] serves as a standardized metadata format for interface declaration.
+Built on OpenAPI foundations but customized for agent consumption, it enables developers to define authenticated entry points, input-output types, and multi-step orchestration plans, such as:
 Flows: Predefined composition of tool steps for common actions.
@@
 It ensures that agents can discover tools introspectively and plan actions without manual reconfiguration or hardcoded logic.
-| Stage | Risk Type | Representative Threats | Attack Characteristics |
-| Malicious environments against benign agents | Memory-related Risks | Memory Injection | Induce agents to generate and record harmful content through natural interactions |
-| Memory Poisoning | Implant pairs of examples containing adversarial triggers and payloads | ||
-| Memory Extraction | Extract sensitive data in memory modules without explicit queries for privacy | ||
-| Knowledge-related Risks | Knowledge Corruption | Inject adversarial texts into external knowledge bases to influence the response | |
-| Privacy Risks | Craft prompts to induce agents to recover sensitive or private content | ||
-| Tool-related Risks | Malicious Tools | Malicious tool functions or prompts embedded in the tool description | |
-| Manipulation of Tool Selection | Inject misleading prompts to make agents select the compromised tool | ||
-| Cross-Tool Chaining Exploits | Manipulate the output of a tool to inject malicious inputs for subsequent tools | ||
-| Compromised agents against benign environments | Information Corruption | Propagate Tampered Information | Undermine an agent’s memory and knowledge using misleading information |
-| Misusing Tools | Data Exfiltration | Extract and transfer confidential information out of protected environments | |
-| System and Service Disruption | Disrupt the functionality of services without directly targeting data confidentiality | ||
-| Malicious Content Propagation | Spread malicious content to trusted storage resources | ||
-| Real-World Damage | Digital Environment Contamination | Introduce harmful elements into digital ecosystems, thereby corrupting integrity | |
-| Physical Environment Disruption | Compromised agents cause disruptions to real-world physical environments |
-VII-B Security Risk Analysis: Malicious Environments Against Benign Agents
-As the capabilities of LLM-powered agents continue to evolve, their interactions with the external world become increasingly complex and powerful.
-In particular, the integration of memory systems and external tool invocation introduces a new set of attack surfaces that attackers can exploit.
-This section provides an in-depth analysis of the security risks that arise specifically from these two modules: the memory module, responsible for storing and retrieving contextual information, and the tool module, which enables agents to execute actions by interfacing with external systems or services (e.g., via function calls).
-We first explain how these two components typically function in the agent ecosystem and outline the general attack paradigms against each.
-We then provide detailed analyses of specific vulnerabilities, attack techniques, and representative works from the security literature that highlight these threats.
-VII-B1 Memory-related Risks
-Memory modules play a crucial role in enabling agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [351].
+| Layer | Source | Risk Category | Representative Threats | Attack Characteristics |
+| L1 | External Attack | Same as user-agent interaction. | ||
+| L2 | Malicious Environments | R1: Tool-related Risks | Malicious Tools | Attackers embed hidden prompts or malicious instructions in tool descriptions or functions. |
+| Manipulation of Selection | Injects misleading prompt elements into metadata to bias the agent’s tool planning. | |||
+| Cross-Tool Chaining | Unvalidated output from one tool injects malicious inputs into subsequent tool calls. | |||
+| L3 | Malicious Environments | R1: Memory -related Risks | Memory Injection | Induces agents to generate and record harmful content through natural interactions. |
+| Memory Poisoning | Implants trigger-payload pairs into memory that activate under specific user queries. | |||
+| Memory Extraction | Uses similarity-based adversarial queries to reconstruct sensitive data from memory. | |||
+| R2: Knowledge -related Risks | Knowledge Corruption | Injects adversarial texts into RAG corpora to manipulate retrieval and model response. | ||
+| Privacy Leakage | Crafts prompts to induce the retrieval and exposure of sensitive data from private corpora. | |||
+| Compromised Agents | R3: Real-world Damage | Content Propagation | Compromised agents spread malware, phishing links, or misinformation via publishing APIs. | |
+| Digital Contamination | Corrupts shared knowledge bases or code repositories, systematically affecting other agents. | |||
+| Physical Disruption | Falsified data or logic leads to harmful physical actions. | |||
+VII-B Security Risks
+In agent-environment communication, the “environment” consists of three core operational components: memory, knowledge base, and external tools.
+VII-B1 Risks from Data Transmission Layer (L1)
+In general, agent-environment communication is more resilient to risks from L1.
+This is because the communication between agents and the environment is usually in a LAN or even on the same machine.
+As a result, they are not likely to face risks in remote communication.
+Of course, the risk level is not thus reduced to zero.
+For example, if agents need to call remote APIs, or the tools and knowledge are downloaded from remote servers, the risks may still happen.
+VII-B2 Risks from Interaction Protocol Layer (L2)
+The Interaction Protocol Layer governs the agent’s interaction with its environment, mediating the translation of high-level intentions into specific, executable actions.
+L2 is a prime target for attackers, as it allows them to influence the agent’s behavior by corrupting the interface between the LLM and environment, rather than attacking the model’s core architecture.
+The most significant risks at L2 are concentrated in the tool-use pipeline, which encompasses how an agent discovers, selects, invokes, and chains external functionalities.
+R1: Tool-related Risks.
+Malicious tools can cause significant harm to agents.
+Tools extend the model’s capabilities to perform structured actions, access external data, invoke system functions, or interact with digital environments.
+Agent architectures typically support tool integration through two primary paradigms: native function calling APIs (e.g., OpenAI-style schema-based calls) and protocol-based interfaces such as the MCP, which unify tool metadata, invocation templates, and language model binding.
+Despite differences in instantiation, both paradigms share a common interaction lifecycle: (1) tool description ingestion, (2) tool selection and planning, (3) input argument generation, (4) tool invocation, and (5) output parsing or chaining.
+We now review a range of known or emerging attacks targeting different stages of the tool interaction process.
+Malicious Tools as Attack Vectors.
+The generation and usage mechanisms of tools are facing serious security problems.
+Tools can be authored externally or retrieved from shared tool repositories, so attackers are able to publish seemingly benign tools containing covert malicious logic.
+For example, researchers reveal that MCP enables attackers to embed hidden prompts or malicious instructions in not only executable functions [hou2025model] but also tool metadata fields such as descriptions, example usages, or API annotations[song2025beyond, fang2025we, hasan2025model, kumar2025mcp, radosevich2025mcp].
+These embedded messages can influence the LLM’s planning behavior, bypass output constraints, execute malicious code, leak privacy, and redirect queries.
+Besides, the ocean of tools has also become a problem for the community [mo2025livemcpbench].
+It is hard to really manage and monitor the continuously emerging tools.
+Manipulation of Tool Selection.
+Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation.
+This selection logic can be hijacked.
+Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options.
+Research indicates that attackers can generate synthetic tool descriptions that stealthily override the model’s planning process [shi2025prompt, mo2025attractive, wang2025mpma, 11209370].
+These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations.
+Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during the planning phase.
+Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
+Cross-Tool Chaining Exploits.
+As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls.
+These workflows blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations.
+Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools) [zhao2025mind].
+In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server”, which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [MCP-Security-in-2025].
+VII-B3 Risks at the Semantic Interpretation Layer (L3)
+The Semantic Interpretation Layer represents the agent’s cognitive core, where information from the environment and past interactions is processed into a coherent context for decision-making.
+Security risks at this layer arise when the information sources that feed this cognitive process are compromised.
+By manipulating such knowledge, an attacker can pervasively alter their reasoning, leading to flawed conclusions, biased responses, or unintended data disclosures.
+This section examines threats that target the integrity of this information supply chain, focusing on two critical components: the memory module, which manages the agent’s episodic and semantic history, and the external knowledge module, which provides factual grounding from external corpora.
+R1: Memory-related Risks.
+Memory modules play a crucial role in enabling agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [zhang2024surveymemorymechanismlarge].
 Unlike stateless language models that depend solely on immediate prompts, memory-equipped agents maintain long-term historical information through external storage systems, such as vector databases or document repositories.
-These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [351].
+These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [zhang2024surveymemorymechanismlarge].
 Typically, a memory module operates through three stages: write, retrieve, and apply.
@@
 We now describe each threat in detail.
-Memory Injection: In memory injection attacks, attackers insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access.
+Memory Injection.
+In memory injection attacks, attackers insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access.
 The attack leverages the agent’s autonomous memory-writing mechanism by inducing it to generate and record harmful content.
 Once stored, these entries can be retrieved by benign user queries due to embedding similarity, thus indirectly triggering undesired behavior such as altered reasoning or unsafe tool invocations.
-A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [61, 307].
+A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [dong2025practical, xiang2024badchainbackdoorchainofthoughtprompting].
 These steps, once embedded in memory, become semantically linked to a targeted victim query.
@@
 This strategy requires no direct injection channels beyond normal user interaction, yet demonstrates high attack success and stealthiness across multiple agent environments.
-Memory Poisoning: Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads.
+Memory Poisoning.
+Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads.
 These attacks are typically conducted by polluting a subset of the memory with trigger-output pairs that only activate when specific inputs are encountered.
 During the retrieval phase, if the user’s query resembles the trigger, the agent is likely to load the poisoned entries and be influenced toward compromised outputs.
-Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [38].
+Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [chen2024agentpoison].
 This method generalizes across agent types and does not require model access or parameter modification.
-Memory Extraction: In addition to injection and poisoning, memory modules pose risks of unintended information leakage.
+Memory Extraction.
+In addition to injection and poisoning, memory modules pose risks of unintended information leakage.
 Since LLM agents often log detailed user-agent interactions-including private file paths, authentication tokens, or sensitive instructions-malicious queries may be used to extract such data.
-This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [332, 128].
-It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [278].
+This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [zeng2024goodbadexploringprivacy, jiang2024rag].
+It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [wang2025unveiling].
 Memory extraction can occur even without explicit queries for private content, instead relying on semantic proximity in the vector space to surface related sensitive traces.
 These findings highlight not only the retrieval vulnerability but also the insufficiency of downstream response filtering as a defense.
-VII-B2 Knowledge-related Risks
+R2: Knowledge-related Risks.
 External knowledge techniques, such as Retrieval-Augmented Generation (RAG), combine the generative strength of LLMs with the factual accuracy and relevance of external knowledge retrieval systems.
@@
 This paradigm enables more informed, up-to-date, and domain-specific language understanding, and it is widely adopted across applications such as open-domain question answering, customer service agents, recommender systems, and multi-step planning agents.
-Despite its performance advantages, the RAG architecture introduces new security risks that are distinct from those inherent to pure neural models [32, 39, 316, 365].
+Despite its performance advantages, the RAG architecture introduces new security risks that are distinct from those inherent to pure neural models [chaudhari2024phantom, chen2024blackboxopinionmanipulationattacks, xue2024badragidentifyingvulnerabilitiesretrieval, zou2024poisonedrag].
 In particular, the information retrieval module-serving as the agent’s external memory-becomes an adversarial surface where unverified or manipulable corpora may be exploited.
 Attacks targeting these corpora can bias the retrieval process, manipulate generation outcomes, or expose previously unseen private data.
-Knowledge Corruption via Data Poisoning: A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries.
+Knowledge Corruption via Data Poisoning.
+A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries.
 These poisoned passages are semantically aligned to specific triggers but contain harmful, misleading, or attacker-intended content.
@@
 Several recent works have demonstrated the feasibility of such attacks.
-PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [365].
-Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [209].
-Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [358].
-Privacy Risks and Unintended Leakage: RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs.
+PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [zou2024poisonedrag].
+Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [nazary2025poison].
+Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [zhong2023poisoningretrievalcorporainjecting].
+Privacy Risks and Unintended Leakage.
+RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs.
 This retrieval behavior implicitly enables information leakage when attackers craft prompts that induce the model to recover sensitive or private content from the corpus.
@@
 Recent studies have called attention to this concern.
-Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [332].
+Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [zeng2024goodbadexploringprivacy].
 These attacks demonstrate that simply adding a retrieval layer does not automatically mitigate the privacy vulnerabilities of LLMs-in fact, it may exacerbate them if not complemented with access control, context filtering, or signal sanitization.
@@
 Moreover, dense retrieval introduces additional attack vectors via embedding collisions or adversarial representation alignment, wherein malicious documents are optimized to collide with benign queries in the retriever’s latent space.
-VII-B3 Tool-related Risks
-Tools are essential to the functionality of LLM agents, extending the model’s capabilities to perform structured actions, access external data, invoke system functions, or interact with digital environments.
-Agent architectures typically support tool integration through two primary paradigms: native function calling APIs (e.g., OpenAI-style schema-based calls) and protocol-based interfaces such as the MCP, which unify tool metadata, invocation templates, and language model binding.
-Despite differences in instantiation, both paradigms share a common interaction lifecycle: (1) tool description ingestion, (2) tool selection and planning, (3) input argument generation, (4) tool invocation, and (5) output parsing or chaining.
-This structured pipeline forms the agent’s “action surface,” introducing multiple security-critical operations vulnerable to adversarial manipulation.
-We now review a range of known or emerging attacks targeting different stages of the tool interaction process.
-Malicious Tools as Attack Vectors: Given that many tools are authored externally or retrieved from shared tool repositories, attackers may publish seemingly benign tools containing covert malicious logic.
-For example, researchers reveal that MCP enables attackers to embed hidden prompts or malicious instructions in not only executable functions [111] but also tool metadata fields such as descriptions, example usages, or API annotations[254, 72, 101, 150, 229].
-These embedded messages can influence the LLM’s planning behavior, bypass output constraints, execute malicious code, leak privacy, and redirect queries.
-Manipulation of Tool Selection: Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation.
-This selection logic can be hijacked.
-Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options.
-Research indicates that attackers can generate synthetic tool descriptions that stealthily override the model’s planning process [250].
-These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations.
-Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during the planning phase.
-Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
-Cross-Tool Chaining Exploits: As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls.
-These workflows blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations.
-Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools).
-In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server”, which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [44].
-VII-C Security Risk Analysis: Compromised Agents Against Benign Environment
-In Section VII-B, we analyzed how the malicious environments can influence an agent through its memory and tool modules.
-However, once an agent is compromised, it can cause widespread and lasting damage to external environments.
-This section explores this reverse threat, focusing on how compromised agents can exploit their functionalities to contaminate memory modules, abuse external tools, and pose risks to the real world.
-VII-C1 Corrupting Memory and Knowledge
-Memory and knowledge bases are central to how the agent thinks and acts.
-Once these components are corrupted, they can lead to incorrect reasoning and decisions.
-Moreover, the damage can infect other agents and cause long-term harm to the entire system.
+R3: Real-world Damage.
+Compromised agents can project harm into the real world through multiple channels.
+This section examines how such agents may disrupt the real world by propagating malicious content, polluting digital environments, and triggering harmful physical actions.
+These pathways together constitute the full spectrum of real-world damage that a compromised agent may cause.
+Malicious Content Propagation.
+Agents with permissions to publish externally (e.g., via email, social media, or CMS APIs) can be exploited to widely spread malware, phishing links, or misinformation once compromised.
+For example, a trusted customer service agent may send malware-laden emails to clients, or a content generation agent might publish misleading articles on official websites.
+Because these agents are trusted entities, such attacks are highly deceptive.
+More dangerously, attackers can leverage access to contacts, email histories, and user preferences to craft highly personalized phishing campaigns, enabling large-scale social engineering attacks.
+Digital Environment Contamination.
+Memory and knowledge—both internal and external—serve as the cognitive substrate for multi-agent systems.
+Once compromised, these environments can induce large-scale, long-term reasoning failures.
+We divide this risk into two complementary subtypes based on the system boundary:
+(a) Internal Digital Environment Contamination.
 A compromised agent can become a source of systemic contamination.
-Through agent communication, it can actively propagate tampered knowledge and flawed reasoning patterns, spreading its internal corruption to other agents and triggering a cascading infection of memory modules and knowledge bases across the system [133, 38, 158].
+Through agent communication, it can actively propagate tampered knowledge and flawed reasoning patterns, spreading its internal corruption to other agents and triggering a cascading infection of memory modules and knowledge bases across the system [ju2024flooding, chen2024agentpoison, lee2024prompt, li2025we].
 Once the shared knowledge base is compromised, other agents may unknowingly retrieve and integrate malicious information into their memory module during task execution, converting knowledge base contamination into memory infection.
@@
 Once established, the contamination can persist long-term, continuously disrupting the behavior of compromised agents and misleading other agents that rely on the same knowledge sources, resulting in a form of chronic poisoning of the information ecosystem.
-VII-C2 Misusing Tools
-Tools serve as the channels through which an agent interacts with the world.
-Once the agent is compromised, tools can be exploited to cause harm.
-Since agents are often granted permissions to execute specific API calls or system commands, misuse of these tools can lead to severe consequences.
-Data Exfiltration: If a compromised agent has tool permissions to access databases, file systems, or communication interfaces, it can be manipulated to extract and transmit confidential data, intellectual property, or personally identifiable information.
-These payloads can be covertly delivered to external servers via email, HTTP requests, or chained tool invocations.
-For instance, a compromised agent equipped with Markdown rendering capabilities might unknowingly embed malicious image links in generated content.
-When rendered by a browser, these links trigger covert HTTP requests that exfiltrate sensitive data such as email addresses or access tokens to attacker-controlled servers.
-Notably, attackers may also exploit tool misuse to gain indirect access to backend systems and gradually escalate privileges through cross-toolchain interactions.
-System and Service Disruption: A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to carry out destructive operations.
-A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to execute destructive operations.
-Attackers may exploit these privileges to launch internal network scans, initiate denial-of-service attacks, or issue commands, potentially causing system crashes or irreversible business data loss.
-Such actions are especially dangerous when triggered via tool interfaces that lack proper input validation.
-Vulnerabilities commonly exploited in these contexts include parameter injection, server-side request forgery (SSRF) [22], and arbitrary file access.
-Beyond technical damage, the compromised agent can be manipulated to disrupt business operations, such as erroneously invoking the “cancel all orders” tool.
-Moreover, adversaries may use retrieval tools to fetch misleading information, then exploit decision-making tools such as trading agents to automatically execute harmful operations, resulting in major financial losses.
-Malicious Content Propagation: Agents with permissions to publish externally (e.g., via email, social media, or CMS APIs) can be exploited to widely spread malware, phishing links, or misinformation once compromised.
-For example, a trusted customer service agent may send malware-laden emails to clients, or a content generation agent might publish misleading articles on official websites.
-Because these agents are trusted entities, such attacks are highly deceptive.
-More dangerously, attackers can leverage access to contacts, email histories, and user preferences to craft highly personalized phishing campaigns, enabling large-scale social engineering attacks.
-VII-C3 Real-World Damage
-Compromised agents can cause harm far beyond their internal systems, affecting both digital and physical ecosystems through their interactions with the external world.
-This section examines how such agents may disrupt the real world by polluting shared digital environments and executing harmful physical actions.
-Digital Environment Contamination: A compromised agent can cause long-term damage to the external digital environment, not by directly attacking other agents, but by polluting the shared information ecosystem they rely on.
-Since agents often interact with external platforms (such as submitting code on GitHub or editing Wikipedia entries), once compromised, they can systematically inject subtle yet harmful errors or biases into these shared resources [173].
+(b) External Digital Environment Contamination.
+A compromised agent can cause long-term damage to the external digital environment, not by directly attacking other agents, but by polluting the shared information ecosystem they rely on.
+Since agents often interact with external platforms (such as submitting code on GitHub or editing Wikipedia entries), once compromised, they can systematically inject subtle yet harmful errors or biases into these shared resources [lin2025exploring].
 Unlike cross-agent contamination, digital environment contamination indirectly infects all agents dependent on the polluted information sources.
 For example, a compromised coding agent may embed hidden logic bugs or backdoors when contributing code; a corrupted knowledge management agent might alter Wikipedia pages or internal knowledge bases by falsifying citations or inserting biased descriptions, thereby damaging the entire knowledge graph with far-reaching consequences.
-Physical Environment Disruption: Once an agent’s memory or tool module is compromised, the resulting threat is not limited to digital risk.
+Physical Environment Disruption.
+Once an agent’s memory or tool module is compromised, the resulting threat is not limited to digital risk.
 It can manifest as concrete damage to the physical world through specific decision-making chains and execution paths.
@@
 Thus, the consequences of agent compromise extend beyond digital misinformation, posing a tangible risk of physical system disruption.
-VII-D Defense Countermeasure Prospect
+VII-C Defense Countermeasure Prospect
 The growing complexity and autonomy of LLM-based agents demand equally sophisticated security strategies.
@@
 This section reviews current and emerging countermeasures along three critical dimensions: memory-based attacks, RAG vulnerabilities, and tool-centric threats.
-VII-D1 Countermeasures For Memory- and knowledge-related Risks
-Memory- and knowledge-related risks can be jointly addressed through an integrated mitigation framework spanning content filtering, output consensus, and architectural isolation.
-Embedding-Space Screening and Clustering-Based Anomaly Detection: Whether memory entries are agent-internal or retrieved externally via RAG, their semantic embeddings can be preemptively analyzed for anomalies.
-Techniques like TrustRAG [359] apply clustering (e.g., K-means) to identify vectors that deviate from the dominant semantic cluster.
+VII-C1 Defenses on Data Transmission Layer (L1)
+The defenses on this layer are similar to those in Section V-C1. Besides, if agents need to call remote resources/tools, it is also necessary to authenticate their identities.
+VII-C2 Defenses on Interaction Protocol Layer (L2)
+At the interaction-protocol layer, tool risks primarily stem from how agents request, negotiate, and coordinate tool actions, making protocol-aware defenses essential before execution begins.
+D1: Defense for Tool-related Risks.
+Tool-related defense strategies should operate across four interlocking levels: protocol foundations, execution control, orchestration safety, and system enforcement.
+Protocol-Level Safeguards.
+To counter risks such as tool poisoning, cross-origin exploits, and shadowing attacks enabled by flexible yet insufficiently regulated protocols like MCP, researchers have introduced security-verification frameworks operating at the registry and middleware layer.
+MCP-Scan [mcpscan] performs both static inspection of tool schemas (e.g., scanning for suspect tags or metadata) and real-time proxy-based validation of MCP traffic, leveraging LLM-assisted heuristics to flag covert behaviors.
+MCP-Shield [riseandignite_mcp_shield_2025] extends this with signature-matching and adversarial behavior profiling, enabling pre-execution detection of high-risk tools and malformed tasks.
+MCIP [jing2025mcip] builds on MAESTRO [MAESTRO] to analyze runtime traces, proposing an explainable logging schema and a security-awareness model to track violations in complex agent-tool interactions.
+MCIP [jing2025mcip] invokes a structured information flow tracing log format.
+And builds a security-aware model trained on this trace data to identify and defend against malicious interactions.
+Tool Invocation and Execution Controls.
+At the agent’s runtime execution point, classic techniques such as sandboxing and permission gating remain foundational.
+Google’s defense-in-depth model advocates policy engines that monitor planned tool actions, verify argument safety, and require human confirmation for risk-sensitive operations [Google-Approach-for-Secure-AI-Agents].
+Tools should be executed in minimally privileged environments-e.g., isolated containers with controlled filesystem and network scope-to mitigate direct misuse, including SSRF and data exfiltration threats.
+Enforcement frameworks can also implement schema hardening or fine-grained input/output sanitization to reject anomalous payloads.
+Agent-Orchestration Monitoring.
+Newer approaches target the agent’s planning cognition-its selection and chaining of tools.
+GuardAgent [xiang2024guardagent] introduces a validator agent that inspects the primary agent’s plan and generates executable guards (e.g., static checks or runtime assertions) before tool calls proceed.
+AgentGuard [chen2025agentguard] takes a more declarative view: it uses an auxiliary LLM to model preconditions, postconditions, and transition constraints across multi-step tool workflows, effectively constraining the planner rather than reacting after execution begins.
+These strategies reflect a growing consensus: LLMs may require another LLM to safely oversee complex planning under uncertainty.
+System-Level Mediation and Chaining Control.
+Complex pipelines-such as summarize(search(“…”))-can become attack vectors when tools trust upstream outputs implicitly.
+To prevent this, DRIFT [han2025drift] introduces a structured control architecture: a “Secure Planner” compiles a validated tool trajectory under strict parametric constraints, while a “Dynamic Validator” continuously monitors downstream tool executions for compliance.
+Notably, the Injection Isolator blocks adversarial propagation between tools by sanitizing both intermediate returns and final outputs, mitigating the risk of memory poisoning and delayed-stage tool exploits.
+Benchmarks.
+Building high-quality benchmarks (such as MCP benchmarks) to test tools in security and performance has become a hot topic [guo2025mcp, guo2025systematic, li2025netmcp, wang2025mcpguard, xing2025mcp, wang2025mcptox, wu2025mcpmark, zhang2025mcp, xu2025tps, jia2025osworld, du2025llm].
+These works either focus on testing potential malicious instructions or the performance in the face of high pressure.
+We believe this direction will continue to evolve because a good benchmark that approximates the real world is hard to achieve, thus it requires continuous efforts.
+VII-C3 Defenses on Semantic Interpretation Layer (L3)
+At the Semantic Interpretation Layer, our defense strategy centers on controlling how semantic context is introduced, interpreted, and propagated through the system.
+By unifying the treatment of memory, knowledge base, outbound communication, and execution, the defenses ensure that semantic signals are continuously verified before they exert real-world impact.
+Together, these mechanisms form a cohesive L3 shield that stabilizes the agent’s semantic grounding and prevents compromised context from escalating into broader systemic harm.
+D1&D2: Defense for Memory/Knowledge-related Risks.
+Although memory-related and knowledge-related risks differ in scope, both inject unverified content into the agent’s reasoning path.
+Thus, they expose similar poisoning surfaces and can be jointly addressed through an integrated mitigation framework spanning content filtering, output consensus, and architectural isolation.
+Embedding-Space Screening and Clustering-Based Anomaly Detection.
+Whether memory entries are agent-internal or retrieved externally via RAG, their semantic embeddings can be preemptively analyzed for anomalies.
+Techniques like TrustRAG [zhou2025trustrag] apply clustering (e.g., K-means) to identify vectors that deviate from the dominant semantic cluster.
 This approach effectively filters both static memory entries and retrieval results with low semantic cohesion, regardless of source.
 While lightweight and interpretable, clustering-based filtering must be augmented with adaptive schemas to detect context-sensitive triggers or stealthy distributional shifts.
-Consensus Filtering and Voting-Based Aggregation: To limit the model’s reliance on single compromised retrievals or poisoned memories, output-level consensus mechanisms have been proposed.
-RobustRAG [306], for example, treats each retrieved source independently and constructs responses based on overlapping semantic content (e.g., shared n-grams or keywords) across documents.
+Consensus Filtering and Voting-Based Aggregation.
+To limit the model’s reliance on single compromised retrievals or poisoned memories, output-level consensus mechanisms have been proposed.
+RobustRAG [xiang2024certifiably], for example, treats each retrieved source independently and constructs responses based on overlapping semantic content (e.g., shared n-grams or keywords) across documents.
 This same principle can be extended to memory snapshots through majority-vote or semantic voting strategies, where only widely corroborated memories can influence the response.
 Such ensemble-style filters improve resilience by diluting the influence of outlier or adversarial sources.
-Execution Monitoring and Planning-Consistency Checks: Adversarial content within memory or RAG inputs may subtly deviate the agent’s behavior from user intent without explicit toxicity.
-Tools like ReAgent [30] introduce planning-level introspection where the agent paraphrases the user’s request, generates an expected plan, and continuously aligns runtime actions with this trace.
+Execution Monitoring and Planning-Consistency Checks.
+Adversarial content within memory or RAG inputs may subtly deviate the agent’s behavior from user intent without explicit toxicity.
+Tools like ReAgent [changjiang2025your] introduce planning-level introspection where the agent paraphrases the user’s request, generates an expected plan, and continuously aligns runtime actions with this trace.
 Any inconsistency, triggered by an unexpected memory or an off-topic retrieval, is treated as a behavioral anomaly and can prompt halting or recovery mechanisms.
 This introspective framework provides a robust guardrail against both memory-hijacking and injection-aware RAG attacks.
-System-Gated Memory Retention and Input Sanitization: Architectural solutions such as DRIFT [99] and AgentSafe [196] implement strict content sanitization before newly generated content.
+System-Gated Memory Retention and Input Sanitization.
+Architectural solutions such as DRIFT [han2025drift] and AgentSafe [mao2025agentsafe] implement strict content sanitization before newly generated content.
 DRIFT uses an injection isolator to scan generative outputs for adversarial goal shifts or impersonation cues, while AgentSafe enforces trust-tiered storage via ThreatSieve and prioritization via HierarCache.
 These mechanisms constrain future influence, ensuring that RAG or memory poisoning cannot silently accumulate over time.
-Unified Content Provenance and Trust Frameworks: Since retrieved knowledge and persisted memories may originate from overlapping sources (e.g., user prompts, tool calls, external APIs), maintaining clear provenance metadata and trust scores is essential.
+Unified Content Provenance and Trust Frameworks.
+Since retrieved knowledge and persisted memories may originate from overlapping sources (e.g., user prompts, tool calls, external APIs), maintaining clear provenance metadata and trust scores is essential.
 Unified provenance tracking across both memory and retrieval pipelines enables smarter decisions about retention, ranking, or discounting of contentious content.
@@
 Moving forward, defense designs should increasingly treat RAG and memory as compositional context modules secured and governed under a shared set of verification, introspection, and isolation principles.
-VII-D2 Countermeasures For Tool-related Risks
-Tool-related defense strategies should operate across four interlocking levels: protocol foundations, execution control, orchestration safety, and system enforcement.
-Protocol-Level Safeguards: To counter risks such as tool poisoning, cross-origin exploits, and shadowing attacks enabled by flexible yet insufficiently regulated protocols like MCP, researchers have introduced security-verification frameworks operating at the registry and middleware layer.
-MCP-Scan [152] performs both static inspection of tool schemas (e.g., scanning for suspect tags or metadata) and real-time proxy-based validation of MCP traffic, leveraging LLM-assisted heuristics to flag covert behaviors.
-MCP-Shield [236] extends this with signature-matching and adversarial behavior profiling, enabling pre-execution detection of high-risk tools and malformed tasks.
-MCIP [130] builds on MAESTRO [49] to analyze runtime traces, proposing an explainable logging schema and a security-awareness model to track violations in complex agent-tool interactions.
-Tool Invocation and Execution Controls: At the agent’s runtime execution point, classic techniques such as sandboxing and permission gating remain foundational.
-Google’s defense-in-depth model advocates policy engines that monitor planned tool actions, verify argument safety, and require human confirmation for risk-sensitive operations [65].
-Tools should be executed in minimally privileged environments-e.g., isolated containers with controlled filesystem and network scope-to mitigate direct misuse, including SSRF and data exfiltration threats.
-Enforcement frameworks can also implement schema hardening or fine-grained input/output sanitization to reject anomalous payloads.
-Agent-Orchestration Monitoring: Newer approaches target the agent’s planning cognition-its selection and chaining of tools.
-GuardAgent [308] introduces a validator agent that inspects the primary agent’s plan and generates executable guards (e.g., static checks or runtime assertions) before tool calls proceed.
-AgentGuard [34] takes a more declarative view: it uses an auxiliary LLM to model preconditions, postconditions, and transition constraints across multi-step tool workflows, effectively constraining the planner rather than reacting after execution begins.
-These strategies reflect a growing consensus: LLMs may require another LLM to safely oversee complex planning under uncertainty.
-System-Level Mediation and Chaining Control: Complex pipelines-such as summarize(search(“…”))-can become attack vectors when tools trust upstream outputs implicitly.
-To prevent this, DRIFT [99] introduces a structured control architecture: a “Secure Planner” compiles a validated tool trajectory under strict parametric constraints, while a “Dynamic Validator” continuously monitors downstream tool executions for compliance.
-Notably, the Injection Isolator blocks adversarial propagation between tools by sanitizing both intermediate returns and final outputs, mitigating the risk of memory poisoning and delayed-stage tool exploits.
-VII-E Takeaways
+D3: Defense for Real-world Damage.
+To mitigate real-world damage, defenses must intervene before reaching external channels or physical systems.
+Since malicious content propagation, digital environment contamination, and physical disruption arise from different execution pathways, there need comprehensive mitigation strategies across different aspects.
+Outbound Content Control.
+To mitigate malicious content propagation, systems should enforce strict governance over all outbound communication channels.
+Agents should be granted only minimal, channel-specific permissions (e.g., “draft-only” access to email or CMS APIs), with high-impact actions (e.g., mass mailings or posts on official accounts) requiring review by dedicated safety agents.
+Before any content is released, it should pass through layered security filters that scan for malware payloads, phishing patterns, and misinformation using both signature-based and LLM-based detectors.
+All outbound messages must be cryptographically signed and logged with detailed provenance, enabling rapid trace-back, revocation, and downstream filtering when a compromised agent is detected.
+Memory/Knowledge Integrity Protection.
+Defending against digital environment contamination requires treating all long-lived knowledge artifacts (memories, shared knowledge bases) as security-sensitive assets.
+Internally, write operations to memory and knowledge bases should be mediated by provenance-aware controllers that record who wrote what, under which task context, and only commit updates after consistency checks.
+Suspicious entries are quarantined, versioned, and periodically re-audited, with rollback mechanisms to restore a clean baseline when contamination is detected.
+Externally, agents should interact with high-impact platforms (e.g., code repositories, wikis) through a hardened “publishing pipeline” that performs anomaly detection on edits, limits the scope of automated changes, and requires human or specialized-agent review for structural updates that could propagate widely.
+Controlled Physical Execution.
+To prevent physical environment disruption, systems must strictly separate high-level semantic planning from low-level actuation, and enforce hard safety envelopes around all physical commands.
+Decisions derived from potentially contaminated memories or tools should first be evaluated in a shadow mode or simulation environment, where their effects on sensors, actuators, and process variables are checked against domain-specific safety constraints.
+Before any irreversible or safety-critical action is executed, the system should require confirmation from an independent validation channel—such as a specialized control model or human supervisor—and automatically block commands that exceed predefined thresholds or violate invariants.
+Continuous monitoring of sensor feedback, combined with fail-safe defaults and emergency stop mechanisms, ensures that anomalous behaviors caused by compromised agents are detected early and contained before they escalate into real-world harm.
+VII-D Takeaways
 Agent-environment communication protocols like MCP enable agents to interface with diverse tools, APIs, and external data.
 However, they introduce risks such as memory injection, retrieval-augmented generation poisoning, and tool misuse.
-Malicious attackers can corrupt memory stores, manipulate knowledge bases, or exploit cross-tool chain vulnerabilities to harm the agent system.
-Besides, compromised agents can also cause damage to benign environments.
-To help developers mitigate these problems, we discuss the targeted defense countermeasures for risks from malicious environments.
-The defenses for compromised agents can be referred to Section VI-C.
-However, we also believe that related attacks will continue to emerge, and there needs long-term efforts to make agent-environment communication more secure.
+Although L1 risks are relatively low due to local communication scenarios, identity authentication and transmission encryption are still required to prevent data leakage during remote tool invocation; as the key layer for tool interaction, L2 needs to resist malicious tool injection and cross-tool chaining attacks through protocol-level security verification, such as tool permission control, and restrict tool execution scope relying on sandbox environments to avoid system damage; L3 is the core of cognitive security.
+To prevent attacks like memory injection and knowledge poisoning, it needs to deploy defenses such as anomaly detection and multi-source knowledge consensus verification.
+Besides, it is necessary to form full-stack defenses to resist damage to physical environments.
 VIII Experimental Case Study: MCP and A2A
@@
 It has been adopted by diverse companies and individual developers.
-Up to now, MCP’s PyPI package has received over 9 million downloads in the last month [20], and its NPM download last week reached 4.2 million [21].
-Besides, the number of open-source MCP servers also approached 2,000 [101].
+Up to now, MCP’s PyPI package has received over 9 million downloads in the last month [pypidownloadmcp], and its NPM download last week reached 4.2 million [MCPpackagedownload].
+Besides, the number of open-source MCP servers also approached 2,000 [hasan2025model].
 After investigation, we find that none of the other existing protocols can achieve a similar popularity to MCP.
@@
 In the experiments related to MCP, we use Claude as the MCP client.
-Claude [17] is an AI chatbot developed by Anthropic.
+Claude [claude] is an AI chatbot developed by Anthropic.
 With its outstanding code capabilities, safe design, and ultra-long context processing, it has become one of the most popular AI assistants in the world.
-In January 2025, Claude app downloads reached an estimated 769.6 million [264].
+In January 2025, Claude app downloads reached an estimated 769.6 million [claudedownload].
 Therefore, exploiting Claude can provide practical value for real-world scenarios.
-Attacks #1-#4 are typical examples of the Malicious Tools as Attack Vectors discussed in Section VII-B3, and attack #5 belongs to the Description Poisoning in Section VI-B1.
+Attacks #1-#4 are typical examples of the Malicious Tools as Attack Vectors discussed in Section VII-B2, and attack #5 belongs to the Description Poisoning in Section VI-B1.
 Exp #1.
 Malicious Code Execution.
-We demonstrate that attackers can exploit MCP to enable the execution of malicious code on the computer system, which has also been proposed by other studies [229].
-We download Claude sonnet 3.7 as the MCP host, and use the Filesystem Server [18] provided by the official MCP as the MCP server.
+We demonstrate that attackers can exploit MCP to enable the execution of malicious code on the computer system, which has also been proposed by other studies [radosevich2025mcp].
+We download Claude sonnet 3.7 as the MCP host, and use the Filesystem Server [filesystem] provided by the official MCP as the MCP server.
 As shown in Figure 15, MCP enables malicious users to directly open an unauthenticated Bash Shell service on port 4444 locally.
@@
 Attackers can write a seemingly normal document related to MCP and publish it on a public platform.
-However, malicious commands are embedded in this document, such as searching for local environment variables and sending the information to Slack [266].
+However, malicious commands are embedded in this document, such as searching for local environment variables and sending the information to Slack [slack].
 The document content is as follows:
-Users download these documents and build a local vector database through the Chroma MCP Server [1].
+Users download these documents and build a local vector database through the Chroma MCP Server [chroma].
 As a result, these malicious documents, together with other legitimate documents, are encoded into the vector database and become one of the retrieval targets.
 When a user sends a request to Claude Opus 4:
-Claude will use Chroma to query the database and retrieve the contaminated documents, and then call the Terminal-Controller tool [88] to execute the embedded malicious commands.
+Claude will use Chroma to query the database and retrieve the contaminated documents, and then call the Terminal-Controller tool [ternimalcontroller] to execute the embedded malicious commands.
 We successfully achieved this attack using Claude Opus 4.
@@
 Tool Poisoning.
-Tool poisoning means that attackers can add malicious instructions in the MCP tool description [254].
+Tool poisoning means that attackers can add malicious instructions in the MCP tool description [song2025beyond].
 Such instructions do not influence the execution of tools, but can induce the agents to perform dangerous actions.
-We use Claude Sonnet 4.0, the Filesystem MCP Sever [18], and Gmail MCP Server [87] to illustrate this vulnerability.
+We use Claude Sonnet 4.0, the Filesystem MCP Server [filesystem], and Gmail MCP Server [gmailmcp] to illustrate this vulnerability.
 In the index.js of the Filesystem Server, we added the malicious description as follows:
@@
 This attack denotes that malicious tools can directly manipulate the user’s system without causing any alert with the help of MCP.
-We inject a malicious command into the Terminal-Controller MCP tool [88].
+We inject a malicious command into the Terminal-Controller MCP tool [ternimalcontroller].
 When users call this tool, the embedded command will be executed.
@@
 We find that attackers can add malicious descriptions in the Agent_Card to induce users to select them.
-In this experiment, we use the public A2A-samples provided by A2Aproject [4] as the user-side client.
+In this experiment, we use the public A2A-samples provided by A2Aproject [a2asamples] as the user-side client.
 Then, we create two agents for weather querying.
-One of them is benign while the other is malicious.
+One of them is benign, while the other is malicious.
 The user issues the following command to the A2A client.
@@
 IX-A1 Powerful but Lightweight Malicious Input Filter
-We deem that user inputs are still the largest-scale attack carrier in the agent ecosystem, especially considering that the inputs are becoming more open (no longer limited to user instructions but also contain environment feedback), multimodal, and semantically complex.
+We deem that user inputs are still the largest-scale attack carrier in the agent ecosystem, especially considering that the inputs are becoming more open (no longer limited to user instructions but also containing environment feedback), multimodal, and semantically complex.
 Besides, the future agent ecosystem will pay more attention to effectiveness, especially given that the running speed of LLMs is inherently slow.
@@
 However, even for open-source LLMs, the publishers still restrict their application scope, e.g., other developers should also open-source their agents built on these LLMs.
-However, there still lacks laws to effectively protect such intellectual property.
+However, there still lack laws to effectively protect such intellectual property.
 For example, the criteria for determining plagiarism in agents is not clear.
@@
 To our knowledge, the related formulation of laws and regulations (such as those related to agent crimes) lags far behind the development of agents.
-For example, how to define the theft and misappropriation of agents, the accident responsibility of autonomous driving agents
+For example, how to define the theft and misappropriation of agents, and the accident responsibility of autonomous driving agents.
 | Abbr. | Full Form |
@@
 | IoA | Internet of Agents |
+| LAN | Local Area Network |
 | LMOS | Language Model Operating System Protocol |
@@
 | TLS | Transport Layer Security |
+| WAN | Wide Area Network |
 X Conclusion
 This survey systematically reviews the security issues of agent communication.
-We first highlight the differences between previous related surveys and this survey, and summarize the preliminaries of LLM-driven agents.
+We first highlight the differences between previous related surveys and this survey, and summarize the evolution direction of communication technology.
 Then, we make a definition and classification of agent communication to help future researchers quickly classify and evaluate their work.